### v1

# New Cold Subdwarf Discoveries from Backyard Worlds and a Metallicity Classification System for T Subdwarfs

Abstract
We report the results of a spectroscopic survey of candidate T subdwarfs identified by the Backyard Worlds: Planet 9 program. Near-infrared spectra of 31 sources with red colors and large -band reduced proper motions show varying signatures of subsolar metallicity, including strong collision-induced H2 absorption, obscured methane and water features, and weak K I absorption. These metallicity signatures are supported by spectral model fits and 3D velocities, indicating thick disk and halo population membership for several sources. We identify three new metal-poor T subdwarfs ([M/H] 0.5), CWISE J062316.19+071505.6, WISEA J152443.14262001.8, and CWISE J211250.11-052925.2; and 19 new “mild” subdwarfs with modest metal deficiency ([M/H] 0.25). We also identify three metal-rich brown dwarfs with thick disk kinematics. We provide kinematic evidence that the extreme L subdwarf 2MASS J053253.46+824646.5 and the mild T subdwarf CWISE J113010.07+313944.7 may be part of the Thamnos population, while the T subdwarf CWISE J155349.96+693355.2 may be part of the Helmi stream. We define a metallicity classification system for T dwarfs that adds mild subdwarfs (d/sdT), subdwarfs (sdT), and extreme subdwarfs (esdT) to the existing dwarf sequence. We also define a metallicity spectral index that correlates with metallicities inferred from spectral model fits and iron abundances from stellar primaries of benchmark T dwarf companions. This expansion of the T dwarf classification system supports investigations of ancient, metal-poor brown dwarfs now being uncovered in deep imaging and spectroscopic surveys.
1 Introduction
Over the past 25 years, wide-field red optical and near-infrared surveys have dramatically expanded our census of the lowest-mass stars, brown dwarfs, and free-floating exoplanets, uncovering thousands of low temperature ultracool dwarfs (UCDs) that span the late-M, L, T, and Y dwarf spectral classes (Kirkpatrick et al., 1999; Burgasser et al., 2006b; Cushing et al., 2011). This sequence is well-sampled among solar-metallicity objects down to effective temperatures 300 K (Luhman, 2014), providing a rich sample for studying the formation, evolution, atmospheres, structures, and spatial distribution of objects that bridge the gap between the lowest mass stars and gas giant planets.
Metal-poor ultracool subdwarfs, including members of the thick disk and halo populations of the Milky Way, remain a very small fraction (0.1%) of the currently known population. The rarity of these sources can be directly attributed to the sparsity of metal-poor halo stars in the vicinity of the Sun (0.5%; Jurić et al. 2008), the long-term cooling of brown dwarfs that makes old metal-poor objects cold and faint, and the relatively small distances (30 pc) to which low-temperature brown dwarfs can be detected. Nevertheless, metallicity variety among the known brown dwarf population is crucial for understanding how composition influences atmospheric chemistry, thermal evolution, interior structure, and the hydrogen burning mass threshold (Saumon et al., 1994; Baraffe et al., 1997; Burrows et al., 2011; Gerasimov et al., 2020). In addition, since UCDs have either long hydrogen fusion lifetimes (1012 yr; Laughlin et al. 1997) or never fuse, these objects record the conditions and chemical abundances of the earliest generations of star formation in the Milky Way (Burgasser, 2009; Lodieu et al., 2015; Caiazzo et al., 2017).
While metal-poor M-type subdwarfs have been known for over 75 years (Joy, 1947; Mould & McElroy, 1978; Gizis, 1997; Lépine et al., 2007), significantly metal-poor brown dwarfs are a recent discovery. The first serendipitous discoveries of metal-poor L-type subdwarfs (Burgasser et al., 2003; Scholz et al., 2004; Sivarani et al., 2009; Lodieu et al., 2010) were made with data from the Two Micron All-Sky Survey (hereafter 2MASS, Skrutskie et al. 2006), the Sloan Digital Sky Survey (hereafter SDSS, York et al. 2000), and the UKIRT Infrared Deep Sky Survey (hereafter UKIDSS, Lawrence et al. 2007). These sources have distinct near-infrared colors and spectra compared to solar-metallicity L dwarfs, shaped by enhanced collision-induced H2 absorption (Linsky, 1969; Burgasser et al., 2003) and metallicity-induced variations in optical molecular absorption and condensate formation (Gizis & Harvin, 2006; Burgasser et al., 2007a; Gonzales et al., 2021). However, most L subdwarfs are low-mass stars, since brown dwarfs at the age of the Galactic thick disk and halo (8 Gyr; Kilic et al. 2017) have temperatures 1500 K, within the late-L and T dwarf regime (Burrows et al., 2001; Baraffe et al., 2003). While several modestly metal-poor T dwarfs ([M/H] 0.5) were found early on by wide-field surveys (Burgasser et al., 2002b; Burningham et al., 2014; Pinfield et al., 2014; Kellogg et al., 2018) and as companions to metal-poor stars (Murray et al., 2011; Pinfield et al., 2012; Scholz, 2010a; Burningham et al., 2013; Mace et al., 2013a), significantly metal-poor substellar T subdwarfs with [M/H] 1 were initially elusive.
The first examples of true T subdwarfs were uncovered through the deep infrared photometry, multi-epoch astrometry, and search efforts of citizen scientists deployed by the Backyward Worlds: Planet 9 program (BYW; Kuchner et al. 2017). This project aims to build the census of low-temperature brown dwarfs in the Solar Neighborhood by providing multi-epoch unWISE images (Lang, 2014; Meisner et al., 2018; Schlafly et al., 2019) to citizen scientists who visually identify faint moving sources. This approach is particularly well-suited for finding low-temperature, nearby thick disk and halo brown dwarfs, whose large velocities translate into large proper motions. Schneider et al. (2020) reported the discovery of WISEA J041451.67585456.7 and WISEA J181006.18101000.5 (hereafter J04145854 and J18101010), whose near-infrared spectra are indicative of both low temperatures ( 1300 K) and significantly subsolar metallicities ([M/H] 1). Meisner et al. (2020, 2021) subsequently reported three additional T subdwarf candidates in BYW data, one (WISE J155349.96+693355.2; hereafter J1553+6933) that exhibits spectroscopic evidence of modest subsolar metallicity ([M/H] 0.5). Brooks et al. (2022) reported the BYW T subdwarf candidate, CWISE J052306.42015355.4 (hereafter J05230153) whose distinct and colors and large proper motion suggest it is a metal-poor member of the thick disk or halo. Finally, WISEA J153429.75104303.3 (hereafter WISE J15341043; aka, “The Accident”) may be the first known metal-poor Y subdwarf (Kirkpatrick et al., 2021a). Candidate thick disk and halo T subdwarfs have also been identified in deep multi-band imaging and spectroscopic surveys conducted with JWST (Nonino et al., 2023; Glazebrook et al., 2023; Wang et al., 2023; Burgasser et al., 2024; Hainline et al., 2024; Holwerda et al., 2024).
Low-temperature subdwarf candidates are exceptionally faint, and most lack the spectroscopic validation or astrometric follow-up necessary to assess physical properties. Indeed, Lodieu et al. (2022) used improved spectroscopy, photometry, and a parallax measurement for J18101010 to conclude that it is both colder ( = 800100 K, equivalent to a T7 dwarf) and more metal-poor ([Fe/H] = 1.50.5 dex) than originally estimated. This study also found evidence that J18101010 lacks methane absorption even in the strong 3.3 band, a defining trait for T dwarfs, suggesting significantly modified atmospheric chemistry. Separately, Burgasser et al. (2024) have claimed evidence for modified PH3/CO2 chemistry in the distant T subdwarf UNCOVER-BD-3 identified with JWST, drawing an analogy to variations in CaH/TiO chemistry among metal-poor M subdwarfs, although this claim has yet to be verified (Beiler et al., 2024).
In this article, we report spectroscopic follow-up of a sample of late-L and T subdwarf candidates uncovered by Backyward Worlds: Planet 9. In Section 2 we describe the selection of candidate T subdwarfs and summarize their associated photometric and astrometric properties. We also estimate distances for metal-poor candidates using new empirical color-magnitude relations based on late-L and T subdwarfs with parallax distance measurements. In Section 3 we describe new near-infrared spectroscopic observations of 24 candidates and three metal-poor L and T dwarf benchmarks. We also define comparison samples of solar metallicity and metal-poor brown dwarf spectra from the literature, including sources that match our selection criteria. In Section 4 we analyze these data, reporting classifications from dwarf standards and spectral indices, K I equivalent widths, spectral model fits, and radial velocities (RVs) from forward modeling. We also use estimated distances, proper motions, and RVs to compute 3D velocities and assess Galactic population membership. In Section 5 we assess the status of our candidates, identifying three new T subdwarfs, 19 new slightly metal-poor T dwarfs, two brown dwarfs with exceptional Galactic orbits, and three apparently metal-rich brown dwarfs with thick disk kinematics. In Section 6 we use our full spectral sample to define anchors for a two-dimensional classification scheme for T dwarfs, extending the existing dwarf sequence to include mild subdwarfs (d/sdT), subdwarfs (sdT), and extreme subdwarfs (esdT). We also construct a calibrated metallicity index that correlates with model-fit metallicities and the iron abundances of the stellar primaries of benchmark T dwarf companions. In Section 7 we examine metallicity dependence in the temperature scale of T dwarfs, continued challenges in modeling T subdwarf spectra, and prospects for expanding the known cold subdwarf population. In Section 8 we summarize our results.
2 BYW T Subdwarf Candidates
2.1 Color and Reduced Proper Motion Selection
The Backyard Worlds: Planet 9 program uses multi-epoch imaging data from the WISE mission to find faint, infrared moving sources, which are identified by a community of over 80,000 citizen scientists. Objects identified as moving point sources are matched to optical and near-infrared photometry and astrometry from 2MASS, UKIDSS, CatWISE2020 (Marocco et al., 2021), Pan-STARRS DR1 (Chambers et al., 2016), Gaia DR3 (Gaia Collaboration et al., 2021), the UKIRT Hemisphere Survey (UHS; Dye et al. 2018), and the Vista Hemisphere Survey (VHS/VIKING; McMahon et al. 2013). There are currently over 4,000 motion-confirmed cool brown dwarf candidates in the BYW sample, most without existing spectroscopic data. Of these, roughly 3,400 sources have well-measured -band (2MASS, UHS/UKIRT, VHS/VIKING) and -band (WISE) photometry, as well as proper motions from CatWISE2020 or Gaia.
Guided by the photometric and astrometric properties of previously identified L and T subdwarfs compiled in Zhang et al. (2018a), Meisner et al. (2021), and Brooks et al. (2022), and “normal” field ultracool dwarfs compiled in Best et al. (2018), we identified a series of selection criteria to isolate high-probability T subdwarf candidates. We imposed a color limit of mag to select sufficiently low temperature objects, and applied a reduced proper motion (RPM) constraint
to select sources with large tangential velocities, indicative of thick disk and halo objects (Figure 1). Note that unlike Meisner et al. (2021) and Brooks et al. (2022) we do not impose a color constraint, using to segregate potential T subdwarfs from contaminant field late-L and T dwarfs. We further required , mag and declination to facilitate follow-up from Keck Observatory. We identified 95 candidate T subdwarfs, including 19 “high-priority” sources with measured mag and/or estimated 150 km/s, the latter based on color/absolute magnitude relations from Kirkpatrick et al. (2021a) or measured parallaxes. These candidates include the previously identified T subdwarfs J18101010 and the sdT5 CWISE J113019.19-115811.3 (hereafter J1130-1158; Kirkpatrick et al. 2021a). Here, we investigate a subset of 31 candidates with infrared spectroscopy.
2.2 Photometric Distance Estimation
The majority of our sample lack trigonometric parallaxes, and assessment of physical properties (temperature, radius) and population membership (kinematics) requires a distance estimate. Gonzales et al. (2018) derived linear relationships between spectral type and absolute 2MASS and WISE magnitudes for late-M and L subdwarfs, but these do not extend into the T dwarf regime. Zhang et al. (2019b) derived separate linear relationships between spectral type and absolute MKO and WISE magnitudes for L and for mid/late T subdwarfs (the latter encompassing relatively modest subsolar metallicities), but were unable to connect these due to the lack of L/T transition subdwarfs known at the time of that study. With several more cool subdwarfs now known, we revisited these analyses with a sample of 18 metal-poor late-L and T dwarfs with measured parallactic distances, either of the source or its primary companion111We exclude three sources from this sample: the esdL6 J0616-6407 and the sdT1 J03012319, whose measured parallaxes from Faherty et al. (2012) and Best et al. (2020) were deemed too low quality for this analysis; and the subdwarf candidate J0055+5947, a co-moving companion to the white dwarf LSPM J0055+5948 (Meisner et al., 2020), which does not appear to be significantly metal-poor. We also excluded previously reported subdwarfs J00042604 and J1055+5443 as our analysis and that of Robbins et al. (2023) indicate they are not significantly metal-poor. (Table 2.2). Figure 2 compares absolute , , and magnitudes to and colors for this sample, as well as empirical relationships for subdwarfs from Gonzales et al. (2018) and Zhang et al. (2019b), and for dwarfs from Dupuy & Liu (2012). There are reasonably linear trends between color and absolute magnitude for these combinations, and good overlap with prior empirical relations for most of the metal-poor T dwarfs. However, the metal-poor late L dwarfs show considerable scatter, particularly in absolute and magnitudes where empirical dwarf relations become nearly vertical (a large range of absolute magnitudes over nearly constant color). In addition, both the esdT J18101010 and the putative Y subdwarf J15341043 are significantly offset from the bulk of the sample in versus , while J15341043 and the T subdwarf companion Wolf 1130C are offset in all of the absolute-magnitude versus color samples. We therefore chose to estimate distances from and versus color-magnitude trends, fitting 16 sources over 1 mag 8.5 mag to second order polynomials by linear least squares regression:
Both relations have a scatter of 0.44 mag, or 22% error in distance. For our analysis, we used the average of both relations to estimate distances, and include in our uncertainty estimates a 0.3 mag scatter in addition to the photometric uncertainties. We emphasize that these relations are based on a limited sample and require validation with a larger parallax sample of verified low-temperature subdwarfs.
Name | SpTaaSpectral type from this paper unless otherwise noted; see Table 10. | Ref. | |||||||
(mag) | (mag) | (mag) | (mas) | (mag) | (mag) | (mag) | |||
2MASS J085039.11-022154.3 | d/sdL6.5 | 15.440.04 | 2.340.05 | 0.2820.016 | 28.21.3 | 12.700.11 | 10.640.10 | 10.360.10 | [1];[2] |
2MASS J11582077+0435014 | d/sdL8 | 15.610.06 | 2.100.06 | 0.2860.023 | 373 | 13.470.16 | 11.650.16 | 11.370.16 | [1];[2] |
2MASS J053253.46+824646.5 | esdL8: | 15.180.06 | 1.500.06 | 0.5960.016 | 40.710.15 | 13.230.06 | 12.3280.013 | 11.7320.014 | [1];[2] |
2MASS J064531536646120 | d/sdT0 | 15.5250.009 | 2.2270.017 | 0.4410.023 | 543 | 14.180.12 | 12.390.12 | 11.950.12 | [3];[4] |
WISEA 181006.18101000.5bbUsing WISE photometry as reported in Lodieu et al. (2022). | esdT3: | 17.290.04 | 4.710.06 | 1.340.05 | 1128 | 17.540.16 | 14.170.16 | 12.830.16 | [5];[5] |
HD 65486B | d/sdT4.5 | 15.780.02 | 1.920.03 | 1.910.04 | 54.1570.018ccParallax from primary. | 14.450.02 | 14.440.04 | 12.5310.017 | [6];[2] |
WISE J044853.28193548.6 | sdT5ddBased on Zhang et al. (2019b). | 16.6080.011 | 2.3780.019 | 2.070.03 | 583 | 15.410.11 | 15.100.12 | 13.030.11 | [3];[4] |
GJ 576B | d/sdT5.5 | 16.590.02 | 2.2840.023 | 1.900.03 | 52.510.02ccParallax from primary. | 15.1920.016 | 14.810.03 | 12.9080.017 | [7];[2] |
2MASS 09373487+2931409 | sdT6 | 14.650.04 | 2.460.04 | 2.4320.017 | 169.41.8 | 15.790.04 | 15.760.03 | 13.330.02 | [1];[8] |
WISE J030919.67501614.3 | d/sdT7.5 | 17.100.02 | 3.4370.021 | 3.060.05 | 674 | 16.220.13 | 15.850.13 | 12.790.13 | [3];[9] |
ULAS J141623.94+134836.3 | sdT7 | 17.260.02 | 4.4280.020 | 3.210.03 | 107.70.2ccParallax from primary. | 17.4210.018 | 16.210.03 | 12.9930.012 | [7];[2] |
LHS 6176B | d/sdT7.5 | 18.050.04 | 3.550.04 | 3.330.10 | 51.000.03ccParallax from primary. | 16.580.04 | 16.360.10 | 13.030.02 | [7];[2] |
HD 126053B | sdT7.5 | 18.710.05 | 3.930.06 | 3.110.12 | 57.270.04ccParallax from primary. | 17.500.05 | 16.670.12 | 13.570.03 | [10];[2] |
WISE J083641.10185947.0 | sdT8ddBased on Zhang et al. (2019b). | 19.100.12 | 3.970.09 | 2.500.09 | 44.22.2 | 17.330.16 | 15.850.14 | 13.360.11 | [3];[4] |
2MASS J093935482448279 | d/sdT8 | 15.7000.007 | 4.0820.011 | 3.310.02 | 19610 | 17.160.12 | 16.390.12 | 13.080.12 | [1],[7];[8] |
Wolf 1130C | sdT6: | 19.640.09 | 4.500.09 | 2.010.04 | 60.300.03ccParallax from primary. | 18.540.09 | 16.060.04 | 14.040.02 | [11];[2] |
WISE J083337.82+005214.1 | d/sdT8 | 20.280.10 | 5.370.10 | 835 | 19.860.15 | 14.490.12 | [12];[13] | ||
WISEA J153429.75104303.3 | (e)sdY:eeBased on Kirkpatrick et al. (2021a). | 24.50.3 | 8.40.3 | 2.040.21 | 615 | 23.40.3 | 17.10.3 | 15.090.19 | [14],[4] |
Note. — photometry is from CATWISE2020 (Marocco et al., 2021) as provided by the Vizier catalog service (Ochsenbein et al., 2000), with exceptions noted.
References. — The first reference is for -band photometry, the second reference is for parallax; [1] 2MASS (Cutri et al., 2003); [2] Gaia EDR3 (Gaia Collaboration et al., 2021); [3] VISTA VHS DR5 (McMahon et al., 2013); [4] Kirkpatrick et al. (2021b); [5] Lodieu et al. (2022); [6] Best et al. (2021); [7] UKIDSS DR9 (Lawrence et al., 2013); [8] Faherty et al. (2012); [9] Kirkpatrick et al. (2019); [10] Pinfield et al. (2012); [11] Mace et al. (2013a); [12] Pinfield et al. (2014); [13] Kirkpatrick et al. (2019); [14] Meisner et al. (2023)
3 Spectroscopic Observations
3.1 Keck/NIRES
We obtained moderate-resolution near-infrared spectroscopy of 20 of our T subdwarf candidates with the Near-Infrared Echellette Spectrometer (NIRES; Wilson et al. 2004) on the Keck II 10m telescope. NIRES is a cross-dispersed spectrograph providing 2700 spectra spanning 0.9–2.45 over four spectral orders, with a fixed 05518 slit. Data were obtained between 2019 October and 2024 January in various conditions (Table 2). All sources were observed with individual exposure times of 300 s in multiple ABBA nodding sets separated by 10 along the slit for background subtraction. One candidate, J0623+0715, resides in a crowded field, so we obtained six exposures at one slit position, and offset to measure the sky background at the start, middle, and end of the sequence. For all sources, we observed a nearby A0 V calibrator star at equivalent airmass to measure telluric and instrumental throughput. We also obtained calibration dome flat lamp exposures at the start of both nights. We observed three additional benchmarks for comparative analysis: the late L subdwarf 2MASS J05325346+8246465 (hereafter J0532+8246; Burgasser et al. 2003) and the metal-poor T dwarf benchmark companions GJ 576B (T5p, [Fe/H] = 0.370.09; Scholz 2010a; Arentsen et al. 2019) and LHS 6176B (T8p, [Fe/H] = 0.300.10; Burningham et al. 2013).
Name | Designation | Instrument | UT Date | Exposure | Airmass | Seeing | Calibrator | SNR | |
(J2000) | (mag) | (s) | () | ||||||
Metallicity Benchmarks and Comparison Sources | |||||||||
J0532+8246 | J053310.53+824620.5 | 15.180.06 | Keck/NIRES | 2022 Jan 19 | 2400 | 2.67 | 07 | HD 48049 | 277 |
GJ 576B | J150457.52+053759.8 | 16.590.02 | Keck/NIRES | 2022 Jun 11 | 2400 | 1.03 | 06 | HD 123233 | 224 |
LHS 6176B | J095047.39+011731.6 | 18.050.04 | Keck/NIRES | 2022 Jan 20 | 3600 | 1.06 | 04bbSeeing significantly less than the NIRES 055 slit width, hence systematic radial velocity offsets are possible (see Section 4.5). | HD 74721 | 91 |
Subdwarf Candidates | |||||||||
J0045+7958 | J004556.21+795847.8 | 16.970.21 | APO/TSpec | 2018 Jan 21 | 3600 | 1.53 | 15 | HD 9612 | 5 |
J0140+0150 | J014016.91+015054.7aaAstrometry from UKIDSS DR9 (Lawrence et al., 2007) epoch 2010.716. | 17.200.02 | IRTF/SpeX | 2020 Dec 24 | 1800 | 1.25 | 05 | HD 1154 | 23 |
Keck/NIRES | 2021 Jan 02 | 2400 | 1.09 | 08 | HD 18571 | 64 | |||
J0411+4714 | J041102.42+471422.6 | 17.670.02 | Keck/NIRES | 2023 Dec 27 | 1200 | 1.34 | 05 | BD+45 981 | 90 |
J0429+3201 | J042941.12+320125.4 | 18.280.08 | Keck/NIRES | 2023 Nov 03 | 2400 | 1.02 | 04bbSeeing significantly less than the NIRES 055 slit width, hence systematic radial velocity offsets are possible (see Section 4.5). | BD+43 1021 | 37 |
J0623+0715 | J062316.20+071506.3 | 17.750.03 | Keck/NIRES | 2022 Jan 19 | 1800 | 1.46 | 05 | HD 37887 | 46 |
J0659+1615 | J065919.43+161552.0 | 17.970.06 | Keck/NIRES | 2024 Jan 30 | 2400 | 1.24 | 05 | HD 47540 | 69 |
J0843+2904 | J084325.37+290435.8 | 17.210.02 | IRTF/SpeX | 2022 Feb 22 | 1800 | 1.18 | 08 | HD 71906 | 15 |
J11101747 | J111055.13174738.8 | 17.680.02 | Keck/NIRES | 2022 Jan 19 | 2400 | 1.27 | 05 | HD 98884 | 76 |
J1130+3139 | J113010.21+313947.0 | 17.790.05 | Keck/NIRES | 2022 Jan 20 | 2400 | 1.05 | 05 | HD 98989 | 90 |
J1138+7212 | J113833.47+721207.8 | 17.080.23 | Gemini/GNIRS | 2013 Dec 03 | 2160 | 1.71 | 07 | HIP 57348 | 13 |
J12042359 | J120444.33235927.4 | 17.910.03 | Keck/NIRES | 2022 Jan 19 | 2400 | 1.40 | 05 | HD 110649 | 44 |
J1304+2819 | J130440.45+281927.7 | 18.100.04 | Keck/NIRES | 2022 Jul 11 | 2400 | 1.30 | 06 | HD 109055 | 40 |
J13080321 | J130841.33032157.8 | 18.070.02 | Keck/NIRES | 2022 Jan 20 | 1800 | 1.09 | 05 | HD 114381 | 72 |
J1401+4325 | J140118.31+432553.6 | 18.400.08 | Keck/NIRES | 2022 Jan 23 | 1800 | 1.09 | 06 | HD 128039 | 36 |
J1458+1734 | J145837.85+173450.2 | 18.390.07 | Keck/NIRES | 2022 Jul 11 | 3600 | 1.15 | 05 | HD 124773 | 58 |
J15152157 | J151521.28215737.3 | 17.480.03 | Magellan/FIRE | 2024 May 29 | 400 | 1.03 | 06 | HD 139273 | 33 |
J15242620 | J152443.15262001.8 | 18.420.10 | Keck/NIRES | 2022 Jun 11 | 2400 | 1.45 | 05 | HD 147384 | 25 |
J1710+4537 | J171003.01+453756.4 | 18.230.06 | Keck/NIRES | 2022 Jun 11 | 3600 | 1.17 | 07 | HD 123299 | 60 |
J1801+4717 | J180122.59+471741.4 | 18.490.09 | Keck/NIRES | 2021 Aug 23 | 2400 | 1.15 | 05 | HD 164899 | 33 |
J20130326 | J201342.30032643.4 | 18.240.12 | Keck/NIRES | 2021 Aug 23 | 2400 | 1.10 | 05 | HD 192230 | 49 |
J2021+1524 | J202130.11+152418.3 | 17.180.02 | Keck/NIRES | 2023 Jul 03 | 2400 | 1.01 | 09 | HD 333145 | 70 |
J21120529 | J211250.14052924.7 | 17.500.03 | Keck/NIRES | 2022 Jun 11 | 2400 | 1.11 | 05 | HD 203769 | 62 |
J2112+3030 | J211255.59+303037.6 | 18.180.06 | Keck/NIRES | 2023 Jul 03 | 3600 | 1.02 | 09 | HD 333145 | 42 |
J22510740 | J225109.56074037.4 | 18.470.08 | Keck/NIRES | 2023 Jul 10 | 2400 | 1.16 | 05 | HD 203769 | 41 |
Note. — All designations are from the CatWISE2020 catalog (Marocco et al., 2021) and are proper motion corrected to epoch 2015 May 28 except for noted sources.
All NIRES data were reduced using a modified version of the Spextool package (Cushing et al., 2004) with standard settings. In brief, we used the dome flat lamps to rectify the spectral orders and correct for pixel-to-pixel response variations, computed the wavelength calibration from OH emission lines in deep exposures, optimally extracted spectra from the rectified 2D traces, averaged multiple spectra for each source order-by-order with a sigma-clipped weighted mean (individual cosmic ray hits were removed by hand), and corrected the spectra for telluric absorption and instrumental response using the A0 V star spectra following the methodology of Vacca et al. (2003). We chose to stitch the individual spectral orders into a single spectrum manually to correct for slight inter-order flux scaling variations not accounted for in the Spextool package. The final data have median signal-to-noise ratios between 23 and 275 at the 1.27 peak at native resolution.
3.2 IRTF/SpeX
Two candidates, J0140+0150 and J0843+2904, were observed with the SpeX spectrograph on the 3m NASA Infrared Telescope Facility (Rayner et al., 2003). J0140+0150 was observed on 2020 Dec 24 (UT) in clear conditions at an average airmass of 1.25; J0843+2904 was observed on 2022 Feb 22 (UT) in partly cloudy conditions at an average airmass of 1.18. Both sources were observed in prism mode with the 0815 slit, providing a resolution of 75 spanning 0.75–2.5 m in a single order, with ten exposures of 180 s each obtained in an ABBA nodding pattern along the slit for background subtraction. A0 V stars were observed immediately after each source, as well as internal flat field and arc lamp exposures. Data were reduced using Spextool version 4.1 with standard settings, following the procedures described above.
3.3 APO/TripleSpec
One candidate, J0045+7958, was observed on 2018 January 21 (UT) with the TripleSpec spectrograph (TSpec; Wilson et al. 2004) on the 3.5m Astrophysical Research Consortium Telescope at Apache Point Observatory (APO). Like NIRES, TSpec is a cross-dispersed spectrograph spanning 0.95–2.46 m at a resolution of 3500 for a 11-wide slit. Conditions were clear with 15 seeing. We acquired twelve 300 s exposures in an ABBA nod pattern at an average airmass of 1.53, followed by observations of the B8 IV standard HD 9612 ( = 6.57) at an airmass of 1.42. We also obtained quartz lamp flat field and NeAr arc lamp exposures illuminated on the dome at the start of the night. Data were reduced using a modified version of Spextool that incorporates instrument parameters for APO/TSpec (cf. Schmidt et al. 2012) following the procedures described above.
3.4 Magellan/FIRE
One candidate, J15152157, was observed on 2024 May 29 (UT) with the Folded-port InfraRed Echellette (FIRE; Simcoe et al. 2013) on the 6.5m Magellan Baade Telescope at Las Campanas Observatory. FIRE was operated in its prism-dispersed mode, providing 0.82–2.51 m spectra at an average resolution of 300 for the deployed 08-wide slit. We acquired eight 50 s exposures in an ABBA nod pattern at an average airmass of 1.03, followed by observations of the A0 V standard HD 139273 ( = 10.13) at the same airmass. We obtained quartz lamp flat field exposures reflected off the Baade flat field screen at two settings to obtain uniform illumination for pixel calibation at the start of the night, and NeAr arc lamp exposures after the science observation for wavelength calibration. Data were reduced using a custom version of the FIREHOSE package (Gagné et al., 2015), which is based on the MASE pipeline (Bochanski et al., 2009) and Spextool, following procedures described in Bickle et al. (2024).
3.5 Gemini-North/GNIRS
One candidate, J1138+7212, was observed on 2013 December 3 (UT) with the Gemini Near-Infrared Spectrograph (GNIRS; Elias et al. 2006) on the 8m Gemini-North telescope. Data were obtained as part of program GN-2013B-Q-91 in queue service mode, and were obtained through light clouds and seeing of 07. The cross-dispersed mode of GNIRS was used with the 0675 slit, short 015 pix-1 camera, and 31.7 lines/mm grating, yielding a 0.9-–2.4 m spectrum at an average resolution of 800. A series of 12 exposures of 180 s each were obtained in an ABBA nod pattern at an average airmass of 1.71, preceded by observations the A5 V star HIP 57348 (V = 10.67) at similar airmass. Data were reduced using the pipeline described in Delorme et al. (2008) and Albert et al. (2011), following similar procedures as noted above, with wavelength calibration achieved using a combination of Ar arc lamps and bright OH telluric emission lines.
3.6 Additional Spectra
In addition to these newly-observed candidates, we included in our analysis spectra of late-L and T dwarfs and subdwarfs previously published in the literature. A sample of 199 L7–T8 spectra were obtained from the SpeX Prism Library Analysis Toolkit (SPLAT; Burgasser & Splat Development Team 2017), which excluded known resolved and candidate binaries and young brown dwarfs (Table A in Appendix A). We also included previously published spectra of T dwarf companions to primary stars with independent iron abundance measurements ranging over 0.64 [Fe/H] 0.25, and previously identified metal-poor late-L and T dwarfs (Table 3.6). For the late L subdwarf J0532+8246, we stitched our Keck/NIRES spectrum with the Keck/LRIS (Oke et al., 1995) optical spectrum reported in Burgasser et al. (2003) by matching in the overlapping region of 0.98–1.0 m. For the T subdwarf J18101010, we stitched together the Palomar/TripleSpec (Wilson et al., 2004) near-infrared spectrum reported by Schneider et al. (2020) with the GTC/OSIRIS (Cepa et al., 2000) optical and GTC/EMIR (Garzón et al., 2007) spectra reported by Lodieu et al. (2022) by matching in the overlapping region of 0.98–1.32 m.
For those sources observed with cross-dispersed spectrographs (Keck/NIRES, APO/TSpec, Palomar/TripleSpec, SOAR/ARCoIRIS) and with well-measured , , and magnitudes, we re-evaluated the relative scaling of the spectral orders to ensure they align with photometry. We computed synthetic spectrophotometric and colors by integrating the spectra over the relevant filter profiles from 2MASS , UKIRT/WFCam , or VISTA 222Filter profiles were obtained from the Spanish Virtual Observatory (SVO) Filter Profile Service (Rodrigo et al., 2012; Rodrigo & Solano, 2020) at http://svo2.cab.inta-csic.es/theory/fps/. using SPLAT tools, with inter-order gaps filled in using a BT-Settl atmosphere model (Allard et al., 2012) at an appropriate temperature, surface gravity, and metallicity (see Section 4.4). For the majority of our sources, the inter-order flux calibration is generally consistent with the photometric uncertainties. However, for J18101010 we found a 50% reduction in -band flux is required to align the spectrum with its UKIDSS photometry (Schneider et al., 2020). We used the appropriately scaled spectrum of this source for subsequent analysis.
Name | Designation | Lit. | Instrument | SNR | [Fe/H] | Ref | |
(J2000) | (mag) | SpT | (dex) | ||||
Metallicity Benchmarks | |||||||
HD 3651B | J003918.91+211516.8aaCoordinates from Faherty et al. (2009). | 16.310.03 | T8 | IRTF/SpeX | 73 | +0.14 | [1]; [2] |
HD 65486B | J075801322538587 | 16.120.08 | T4.5 | IRTF/SpeX | 34 | 0.28 | [30]; [31],[34]ddData re-reduced from the IRSA Legacy Archive. |
Ross 458C | J130041.40+122114.3 | 16.690.02 | T8.5p | Magellan/FIRE | 67 | +0.25 | [28]; [29] |
HD 126053B | J142320.90+011635.3 | 18.710.05 | T8p | Gemini/GNIRS | 24 | 0.38 | [25] |
Gliese 570D | J145716.21212216.5 | 15.320.05 | T7.5 | IRTF/SpeX | 37 | +0.05 | [3]; [4] |
GJ 576B | J150457.34+053757.0 | 16.590.02 | sdT5.5 | Gemini/GNIRS | 56 | 0.37 | [5]; [6] |
VLT/X-Shooter | 21 | [7] | |||||
G 204-38B | J175805.45+463319.8 | 16.150.09 | T6.5 | IRTF/SpeX | 75 | 0.04 | [8]; [9] |
Wolf 1130C | J200519.70+542428.5 | 19.640.09 | sdT8 | Keck/NIRSPEC | 78 | 0.70 | [10]; [11] |
Keck/MOSFIRE | [10] | ||||||
HN Peg B | J214428.67+144606.0bbCoordinates from Pan-STARRS DR1 (Chambers et al., 2016) epoch 2011 Nov 25. | 15.860.03 | T2.5 | IRTF/SpeX | 114 | 0.08 | [12]; [12] |
Wolf 940B | J214638.70001042.3 | 18.020.06 | T8 | Magellan/FIRE | 69 | +0.08 | [26]; [6] |
Metal-poor Comparison Sources | |||||||
J00041336 | J000458.59133655.2 | 16.840.17 | T2p | IRTF/SpeX | 18 | [19] | |
J00042604 | J000430.67260403.6 | 16.490.13 | sdT2 | IRTF/SpeX | 40 | [19] | |
J0013+0634 | J001354.85+063445.4 | 19.750.05 | T8p | Gemini/GNIRS | 22 | [18] | |
J0021+1552 | J002135.97+155226.8 | 17.880.04 | T4p | VLT/X-Shooter | 9 | [7] | |
J0055+5947ccAlso identified as a T subdwarf candidate by our selection criteria. | J005559.94+594744.4 | 17.900.04 | T8 | Keck/NIRES | 13 | [22] | |
J0057+2013 | J005757.95+201302.8 | 16.320.10 | sdL7 | IRTF/SpeX | 40 | [24] | |
J03012319 | J030119.47231921.8 | 16.640.14 | sdT1 | IRTF/SpeX | 25 | [19] | |
J03095016 | J030919.92501614.0 | 17.170.03 | T7p | Magellan/FIRE | 37 | [19] | |
J03485620 | J034858.85562016.9 | 16.520.15 | T3p | Magellan/FIRE | 22 | [19] | |
J04145854 | J041451.86585455.3 | 19.630.11 | esdT | Magellan/FIRE | 6 | [13] | |
J0433+1009ccAlso identified as a T subdwarf candidate by our selection criteria. | J043309.36+100902.3 | 18.000.03 | T8 | Keck/NIRES | 32 | [15] | |
J0532+8246 | J053310.53+824620.5 | 15.180.06 | esdL7 | Keck/LRIS | 1.6 | [14],[16] | |
J06166407 | J061643.51640719.8 | 16.350.01 | esdL6 | VLT/X-Shooter | 26 | 1.6 | [16] |
J06456646 | J064529.11664550.7 | 15.520.01 | sdL8 | VLT/X-Shooter | 47 | [17] | |
J0833+0052 | J083338.11+005206.2 | 20.280.10 | T9p | Gemini/GNIRS | 15 | [18] | |
J08453305ccAlso identified as a T subdwarf candidate by our selection criteria. | J084506.49330532.1 | 16.850.01 | T7 | Magellan/FIRE | 82 | [15] | |
J08500221 | J085038.99022155.1 | 15.440.04 | sdL7 | IRTF/SpeX | 83 | [19] | |
J0911+2146ccAlso identified as a T subdwarf candidate by our selection criteria. | J091105.03+214645.4 | 17.750.03 | T9 | Magellan/FIRE | 29 | [15] | |
J0937+2931 | J093735.65+293127.2 | 14.650.04 | T6p | IRTF/SpeX | 128 | [8] | |
J09392448 | J093936.13244844.0 | 15.670.01 | T8p | IRTF/SpeX | 17 | 0.24 | [8],[35] |
J09530943ccAlso identified as a T subdwarf candidate by our selection criteria. | J095316.33094318.8 | 17.010.03 | T5.5 | IRTF/SpeX | 24 | [15],[34]ddData re-reduced from the IRSA Legacy Archive. | |
J10193911 | J101944.47391150.9 | 16.030.10 | sdT3 | SOAR/ARCoIRIS | 20 | [19] | |
J10350711 | J103534.52071148.5 | 16.390.09 | sdL7 | IRTF/SpeX | 33 | [19] | |
J1055+5443 | J105512.03+544328.3 | 18.870.07 | sdT8/Y0 | Keck/NIRES | 29 | [15],[20] | |
J11301158 | J113019.20115811.5 | 17.220.03 | sdT5 | SOAR/ARCoIRIS | 11 | [15] | |
J1158+0435 | J115821.40+043446.3 | 15.610.06 | sdL7 | IRTF/SpeX | 103 | [32] | |
J1307+1511 | J130710.04+151102.4 | 18.140.04 | sdL8 | VLT/X-Shooter | 10 | [17] | |
J1316+0755 | J131609.68+075553.9 | 19.290.12 | sdT6.5 | Gemini/GNIRS | 13 | [5] | |
J13380229 | J133836.94022912.2 | 17.370.03 | sdL7 | Magellan/FIRE | 7 | [16] | |
J1416+1348B | J141623.96+134837.3 | 17.260.02 | T7.5p | IRTF/SpeX | 48 | 0.35 | [21],[35] |
J1553+6933 | J155347.98+693400.5 | 19.090.07 | sdT4 | Keck/NIRES | 19 | [22],[33] | |
J18101010ccAlso identified as a T subdwarf candidate by our selection criteria. | J181005.99101001.8 | 17.270.02 | esdT | Palomar/TripleSpec | 33 | 1.5 | [13],[23] |
GTC/OSIRIS+EMIR | [23] | ||||||
J21056235 | J210529.52623604.6 | 16.850.14 | sdT1.5 | Gemini/Flamingos | 69 | [24] | |
J2218+1146ccAlso identified as a T subdwarf candidate by our selection criteria. | J221859.41+114642.9 | 17.960.03 | T7p | Keck/NIRES | 51 | [22] |
Note. — All designations are from the CatWISE2020 catalog (Marocco et al., 2021) and are proper motion corrected to epoch 2015 May 28 except for noted sources.
References. — The first reference corresponds to the source of the spectral data; the second reference corresponds to the source of the metallicity (for benchmarks): [1] Burgasser (2007a); [2] Rice & Brewer (2020); [3] Burgasser et al. (2004); [4] Aguilera-Gómez et al. (2018); [5] Burningham et al. (2014); [6] Arentsen et al. (2019); [7] Zhang et al. (2019b); [8] Burgasser et al. (2006a); [9] Schweitzer et al. (2019); [10] Mace et al. (2013a); [11] Mace et al. (2018); [12] Luhman et al. (2007); [13] Schneider et al. (2020); [14] Burgasser et al. (2003); [15] Kirkpatrick et al. (2021b); [16] Zhang et al. (2017); [17] Zhang et al. (2018b); [18] Pinfield et al. (2014); [19] Greco et al. (2019); [20] Robbins et al. (2023); [21] Burgasser et al. (2010b); [22] Meisner et al. (2020); [23] Lodieu et al. (2022); [24] Luhman & Sheppard (2014); [25] Pinfield et al. (2012); [26] Burgasser et al. (2011); [27] Marfil et al. (2021); [28] Burgasser et al. (2010c); [29] Rosenthal et al. (2021); [30] Deacon et al. (2012a); [31] Soubiran et al. (2022); [32] Kirkpatrick et al. (2010); [33] Meisner et al. (2021); [34] This paper; [35] Line et al. (2017).
4 Spectral Analysis
4.1 Comparison to Spectral Templates
Figures 3 and 4 display the reduced spectral data for our subdwarf candidates in both native resolution and smoothed to 125. The smoothed spectra are compared to low resolution spectra of L and T dwarf near-infrared spectral standards from SPLAT333Spectral data for late-L and T dwarf standards are from Burgasser et al. (2004, 2006b, 2007b); Looper et al. (2007); and Bardalez Gagliuffi et al. (2014). as defined by Kirkpatrick et al. (2010) and Burgasser et al. (2006b). A best-fit standard for each source was selected based on comparison over the 0.9-1.4 range using a statistic (see Eqn. 4); these standard types are listed in Table 4. For most of the sources, including the companions GJ 576B and LHS 6176B, we see close agreement with T6–T9 standards, with minor deviations in the -band (1 ) and -band (2.1 ) peaks consistent with slightly subsolar metallicities (Burgasser et al., 2006a). A handful of sources appear to be late-type L or early-type T dwarfs based on weak or absent 1.6 CH4 absorption, and several exhibit even more suppressed -band peaks relative to their matched standards. For our higher resolution Keck/NIRES data, we also see variety in the strengths of the 1.2436 m and 1.2526 m K I doublet, in many cases weaker than the corresponding NIRES dwarf standard (Theissen et al., 2022). We quantify these spectral deviations below.
4.2 Spectral Indices
For all of our spectra, we evaluated ten near-infrared spectral indices that sample molecular features (H2O and CH4 bands) and overall spectral shape, defined in Burgasser et al. (2006a, b) and Bardalez Gagliuffi et al. (2019). Tables 4 and A summarize the index values for our candidates and comparison sources, with uncertainties determined through Monte Carlo sampling of spectral flux errors. We also list index-based spectral classifications based on the index versus spectral type relationships of Burgasser (2007b) which are defined for spectral types L0–T9. Note that these relationships are defined for solar-metallicity dwarfs and are inaccurate for exceptionally metal-poor sources. Indeed, classifications from matched standards (Figure 3) and from indices can differ by more than 3 subtypes (e.g., J0532+8246).
SpTaaSpectral type as inferred from direct comparison to spectral standards (“Std”) and indices (“Ind”). | ||||||||||||
Name | [CH4J] | [CH4H] | [CH4K] | [H2OJ] | [H2OH] | [H2OK] | [Y/J] | [K/H] | [K/J] | [Hdip] | Std | Ind |
(numerator) | 1.315–1.335 | 1.635–1.675 | 2.215–2.255 | 1.14–1.165 | 1.48–1.52 | 1.975–1.995 | 1.005–1.045 | 2.06–2.10 | 2.06–2.10 | 1.61–1.64 | ||
(denominator) | 1.26–1.285 | 1.56–1.60 | 2.08–2.12 | 1.26–1.285 | 1.56–1.60 | 2.08–2.12 | 1.25–1.29 | 1.25–1.29 | 1.56–1.60 | 1.56–1.59,1.66–1.69 | ||
Benchmark Companions | ||||||||||||
HD 3651B | 0.250.01 | 0.140.01 | 0.040.02 | 0.040.01 | 0.170.01 | 0.320.02 | 0.420.01 | 0.290.01 | 0.140.01 | 0.250.02 | T8 | T7.7 |
HD 65486B | 0.640.02 | 0.500.02 | 0.260.07 | 0.320.01 | 0.370.02 | 0.560.09 | 0.640.02 | 0.300.02 | 0.140.01 | 0.590.03 | T4.5 | |
LHS 6176B | 0.260.01 | 0.160.01 | 0.120.01 | 0.060.01 | 0.160.01 | 0.390.02 | 0.520.01 | 0.220.01 | 0.110.01 | 0.310.01 | T8 | T7.5 |
Ross 458C | 0.190.01 | 0.120.01 | 0.070.01 | 0.050.01 | 0.190.01 | 0.360.01 | 0.350.01 | 0.450.01 | 0.260.01 | 0.230.01 | T8 | T8.1 |
HD 126053B | 0.270.02 | 0.140.01 | 0.100.05 | 0.050.01 | 0.220.02 | 0.370.05 | 0.540.03 | 0.110.01 | 0.050.01 | 0.400.03 | T7 | T7.5 |
GJ 570D | 0.280.01 | 0.130.01 | 0.110.02 | 0.060.01 | 0.200.01 | 0.340.02 | 0.410.01 | 0.240.01 | 0.110.01 | 0.340.02 | T8 | T7.5 |
GJ 576B (GNIRS) | 0.430.01 | 0.370.01 | 0.210.05 | 0.160.01 | 0.300.02 | 0.470.04 | 0.520.01 | 0.220.01 | 0.100.01 | 0.670.03 | T6 | T5.6 |
(XS) | 0.430.02 | 0.400.02 | 0.130.15 | 0.150.02 | 0.290.02 | 0.500.09 | 0.540.04 | 0.190.03 | 0.080.01 | 0.640.02 | T6 | T5.5 |
(NIRES) | 0.440.01 | 0.400.01 | 0.250.01 | 0.150.01 | 0.280.01 | 0.490.01 | 0.450.01 | 0.250.01 | 0.100.01 | 0.610.01 | T6 | T5.5 |
G 20439B | 0.410.01 | 0.250.01 | 0.100.01 | 0.100.01 | 0.240.01 | 0.400.01 | 0.410.01 | 0.400.01 | 0.200.01 | 0.440.01 | T6 | T6.4 |
Wolf 1130C | 0.820.03 | 0.420.01 | bbMissing spectral data or skewed by excess noise. | 0.010.01 | 0.390.01 | bbMissing spectral data or skewed by excess noise. | 0.740.11 | bbMissing spectral data or skewed by excess noise. | bbMissing spectral data or skewed by excess noise. | 0.750.01 | T6 | T5.4 |
HN PegB | 0.630.01 | 0.840.01 | 0.530.01 | 0.380.01 | 0.430.01 | 0.540.01 | 0.550.01 | 0.460.01 | 0.310.01 | 0.980.01 | T3 | T2.4 |
Wolf 940B | 0.120.02 | 0.080.03 | 0.030.09 | 0.030.02 | 0.140.04 | 0.290.08 | bbMissing spectral data or skewed by excess noise. | 0.240.01 | 0.130.01 | 0.150.06 | T8 | T8.6 |
Subdwarf Candidates | ||||||||||||
J0045+7958 | 0.820.09 | 1.120.06 | 0.900.05 | 0.670.11 | 0.600.06 | 0.620.04 | 0.570.17 | 0.630.03 | 0.500.04 | 1.020.05 | T0 | L8.6: |
J0055+5947 | 0.350.02 | 0.270.02 | 0.160.05 | 0.110.03 | 0.250.04 | 0.360.08 | 0.600.06 | 0.230.01 | 0.110.01 | 0.460.03 | T6 | T6.4 |
J0140+0150 (SpeX) | 0.590.02 | 0.580.03 | 0.270.05 | 0.280.02 | 0.340.03 | 0.470.05 | 0.450.02 | 0.440.02 | 0.250.01 | 0.820.04 | T4 | T4.2 |
(NIRES) | 0.590.01 | 0.560.01 | 0.260.01 | 0.240.01 | 0.320.01 | 0.470.01 | 0.440.01 | 0.440.01 | 0.240.01 | 0.780.01 | T4 | T4.5 |
J0411+4714 | 0.270.01 | 0.160.01 | 0.100.01 | 0.080.01 | 0.210.01 | 0.460.01 | 0.460.01 | 0.240.01 | 0.100.01 | 0.330.01 | T7 | T7.4 |
J0429+3201 | 0.850.02 | 0.960.01 | 0.770.01 | 0.560.01 | 0.550.01 | 0.590.01 | 0.900.03 | 0.560.01 | 0.450.01 | 1.030.01 | T1 | T1.1 |
J0433+1009 | 0.220.01 | 0.160.01 | 0.100.02 | 0.090.01 | 0.220.01 | 0.340.02 | 0.400.03 | 0.340.01 | 0.190.01 | 0.310.02 | T8 | T7.3 |
J0623+0715 | 0.820.02 | 0.640.01 | 0.580.01 | 0.500.01 | 0.600.02 | 0.770.02 | 0.720.03 | 0.340.01 | 0.250.01 | 0.890.01 | T3 | T3.4 |
J0659+1615 | 0.870.01 | 0.950.01 | 0.760.01 | 0.620.01 | 0.580.01 | 0.610.01 | 0.650.02 | 0.540.01 | 0.440.01 | 1.030.01 | T1 | T1.0 |
J0758+5711 | 0.350.01 | 0.260.01 | 0.170.01 | 0.120.01 | 0.230.01 | 0.450.02 | 0.560.02 | 0.270.01 | 0.140.01 | 0.470.01 | T7 | T6.4 |
J0843+2904 | 0.930.03 | 1.070.04 | 0.900.03 | 0.830.04 | 0.720.04 | 0.750.03 | 0.660.03 | 0.740.03 | 0.670.03 | 1.000.03 | L6 | L5.9 |
J08453305 | 0.380.01 | 0.270.01 | 0.220.02 | 0.120.01 | 0.280.01 | 0.420.01 | 0.460.01 | 0.240.01 | 0.120.01 | 0.460.01 | T7 | T6.2 |
J0911+2146 | 0.190.01 | 0.130.01 | bbMissing spectral data or skewed by excess noise. | 0.040.01 | 0.170.01 | 0.560.05 | 0.370.02 | 0.190.01 | 0.100.01 | 0.290.02 | T8 | T8.2 |
J11101747 | 0.200.01 | 0.100.01 | 0.040.01 | 0.050.01 | 0.170.01 | 0.320.01 | 0.460.01 | 0.320.01 | 0.190.01 | 0.220.01 | T8 | T8.2 |
J1130+3139 | 0.450.01 | 0.380.01 | 0.260.01 | 0.140.01 | 0.260.01 | 0.470.02 | 0.580.01 | 0.270.01 | 0.110.01 | 0.630.01 | T6 | T5.6 |
J1138+7212 | 0.310.02 | 0.230.03 | 0.180.08 | 0.130.03 | 0.260.04 | 0.580.10 | 0.560.04 | 0.210.01 | 0.090.01 | 0.400.05 | T7 | T6.5 |
J12042359 | 0.310.01 | 0.240.01 | 0.100.02 | 0.090.01 | 0.210.02 | 0.330.02 | 0.500.03 | 0.250.01 | 0.130.01 | 0.420.02 | T7 | T6.8 |
J1304+2819 | 0.920.02 | 1.020.01 | 0.900.01 | 0.750.02 | 0.680.01 | 0.690.01 | 0.560.02 | 0.630.01 | 0.440.01 | 0.980.01 | L9 | L7.4 |
J13080321 | 0.220.01 | 0.120.01 | 0.120.01 | 0.030.01 | 0.160.01 | 0.360.01 | 0.500.01 | 0.240.01 | 0.110.01 | 0.250.01 | T8 | T7.9 |
J1401+4325 | 0.420.02 | 0.360.01 | 0.270.03 | 0.090.01 | 0.270.02 | 0.620.04 | 0.380.02 | 0.290.01 | 0.100.01 | 0.600.02 | T7 | T5.6 |
J1458+1734 | 0.210.01 | 0.100.01 | 0.050.01 | 0.010.01 | 0.180.01 | 0.360.01 | 0.400.02 | 0.300.01 | 0.150.01 | 0.230.01 | T8 | T8.2 |
J15152157 | 0.430.01 | 0.280.01 | 0.230.02 | 0.120.01 | 0.180.01 | 0.390.02 | 0.510.03 | 0.230.01 | 0.100.01 | 0.500.02 | T6 | T6.1 |
J15242620 | 0.870.02 | 0.940.01 | 0.770.01 | 0.670.02 | 0.720.02 | 0.720.01 | 0.480.03 | 0.480.01 | 0.360.01 | 1.010.01 | L9 | L9.3 |
J1710+4537 | 0.400.01 | 0.310.01 | 0.260.02 | 0.140.01 | 0.260.01 | 0.440.02 | 0.410.02 | 0.300.01 | 0.150.01 | 0.510.01 | T6 | T5.9 |
J1801+4717 | 0.440.01 | 0.430.01 | 0.280.03 | 0.180.01 | 0.330.02 | 0.370.03 | 0.650.03 | 0.340.01 | 0.130.01 | 0.700.02 | T5 | T5.1 |
J20130326 | 0.400.01 | 0.340.01 | 0.220.02 | 0.130.01 | 0.300.01 | 0.600.03 | 0.580.02 | 0.250.01 | 0.120.01 | 0.570.02 | T6 | T5.8 |
J2021+1524 | 0.860.01 | 1.040.01 | 0.910.01 | 0.620.01 | 0.560.01 | 0.590.01 | 0.530.01 | 0.560.01 | 0.390.01 | 1.030.01 | T0 | L9.8 |
J21120529 | 0.820.01 | 0.880.01 | 0.770.01 | 0.610.01 | 0.600.01 | 0.680.01 | 0.580.02 | 0.420.01 | 0.250.01 | 1.010.01 | T1 | T1.4 |
J2112+3030 | 0.780.01 | 0.740.01 | 0.510.01 | 0.530.01 | 0.500.01 | 0.610.01 | 0.930.03 | 0.390.01 | 0.280.01 | 0.930.01 | T3 | T2.6 |
J2218+1146 | 0.370.01 | 0.240.01 | 0.140.02 | 0.110.01 | 0.300.01 | 0.510.03 | 0.410.02 | 0.260.01 | 0.100.01 | 0.460.02 | T7 | T6.4 |
J22510740 | 0.280.01 | 0.190.01 | 0.140.02 | 0.060.01 | 0.160.01 | 0.340.03 | 0.690.03 | 0.260.01 | 0.160.01 | 0.390.02 | T7 | T7.1 |
Metal-poor Comparison Sources | ||||||||||||
J00041336 | 0.760.02 | 0.950.03 | 0.850.05 | 0.570.02 | 0.590.04 | 0.630.04 | 0.550.02 | 0.440.02 | 0.260.01 | 0.960.03 | T1 | T0.6 |
J00042604 | 0.700.01 | 0.750.01 | 0.570.02 | 0.470.01 | 0.520.02 | 0.590.02 | 0.570.01 | 0.410.01 | 0.270.01 | 0.900.02 | T3 | T2.5 |
J0013+0634 | 0.260.02 | 0.320.04 | 0.210.09 | 0.090.05 | 0.170.02 | bbMissing spectral data or skewed by excess noise. | 0.500.04 | 0.090.01 | 0.040.01 | 0.380.06 | T7 | T6.2 |
J0021+1552 | 0.670.04 | 0.570.03 | 0.260.17 | 0.390.05 | 0.460.04 | 0.670.14 | 0.520.11 | 0.270.04 | 0.170.03 | 0.830.05 | T4 | T4.0 |
J0057+2013 | 0.840.01 | 1.020.01 | 0.950.02 | 0.660.01 | 0.590.01 | 0.700.01 | 0.570.01 | 0.530.01 | 0.380.01 | 0.970.01 | T0 | L8.7 |
J03012319 | 0.820.02 | 0.980.02 | 0.730.02 | 0.670.02 | 0.670.02 | 0.630.02 | 0.740.02 | 0.490.01 | 0.330.01 | 1.010.02 | T1 | L9.6 |
J03095016 | 0.200.01 | 0.160.01 | 0.160.08 | 0.050.01 | 0.230.01 | 0.510.05 | 0.560.02 | 0.140.01 | 0.060.01 | 0.300.02 | T7 | T7.7 |
J03485620 | 0.540.02 | 0.590.02 | 0.300.04 | 0.390.02 | 0.470.02 | 0.610.04 | 0.500.02 | 0.340.02 | 0.200.01 | 0.790.03 | T4 | T4.1 |
J04145854 | 0.900.06 | 0.580.07 | bbMissing spectral data or skewed by excess noise. | 0.220.07 | 0.810.09 | bbMissing spectral data or skewed by excess noise. | 0.910.09 | 0.040.04 | 0.020.02 | 0.890.09 | T4 | T3.9 |
J0532+8246 | 0.850.01 | 0.830.01 | 0.810.01 | 0.830.01 | 0.860.01 | 0.880.01 | 0.560.01 | 0.350.01 | 0.170.01 | 1.000.01 | L8 | L4.6 |
J06166407 | 0.860.03 | 0.820.02 | 0.950.08 | 0.890.03 | 0.920.03 | 0.840.06 | 0.590.03 | 0.310.02 | 0.150.01 | 0.970.02 | T0 | T1.1 |
J06456646 | 0.850.01 | 1.010.01 | 0.870.02 | 0.680.01 | 0.630.01 | 0.660.01 | 0.520.02 | 0.480.01 | 0.340.01 | 1.030.01 | T0 | L9.0 |
J08500221 | 0.890.01 | 1.060.01 | 0.950.01 | 0.770.01 | 0.700.01 | 0.710.01 | 0.480.01 | 0.550.01 | 0.400.01 | 0.980.01 | L6 | L6.5 |
J0833+0052 | 0.090.02 | 0.120.03 | bbMissing spectral data or skewed by excess noise. | 0.040.04 | 0.070.04 | 0.290.09 | 0.510.05 | 0.170.02 | 0.100.01 | 0.300.05 | T8 | |
J0937+2931 | 0.430.01 | 0.310.01 | 0.180.01 | 0.150.01 | 0.310.01 | 0.530.02 | 0.530.01 | 0.170.01 | 0.070.01 | 0.530.01 | T6 | T5.8 |
J09392448 | 0.300.01 | 0.140.01 | 0.100.04 | 0.030.01 | 0.170.01 | 0.520.05 | 0.510.02 | 0.130.01 | 0.060.01 | 0.330.02 | T8 | T7.7 |
J09530943 | 0.500.02 | 0.370.03 | 0.210.10 | 0.210.02 | 0.180.05 | 0.580.13 | 0.450.02 | 0.250.03 | 0.100.01 | 0.500.06 | T6 | T5.1 |
J10193911 | 0.660.01 | 0.550.01 | 0.290.01 | 0.440.01 | 0.400.01 | 0.500.01 | 0.480.01 | 0.360.01 | 0.200.01 | 0.770.01 | T4 | T4.0 |
J10350711 | 0.890.02 | 1.070.02 | 0.940.02 | 0.720.01 | 0.630.02 | 0.660.02 | 0.580.01 | 0.560.01 | 0.380.01 | 1.040.02 | L9 | L7.7 |
J1055+5443 | 0.050.01 | 0.060.01 | 0.050.03 | 0.020.02 | 0.110.02 | 0.270.03 | 0.790.05 | 0.310.01 | 0.230.01 | 0.100.03 | T9 | |
J11301158 | 0.490.03 | 0.330.03 | 0.690.46 | 0.210.06 | 0.380.05 | 1.110.27 | 0.490.05 | 0.130.02 | 0.060.01 | 0.620.06 | T6 | T5.4 |
J1158+0435 | 0.860.01 | 1.040.01 | 0.890.01 | 0.710.01 | 0.640.01 | 0.690.01 | 0.470.01 | 0.500.01 | 0.320.01 | 1.010.01 | L7 | L8.0 |
J1307+1510 | 0.790.05 | 1.030.04 | 0.950.12 | 0.660.06 | 0.580.03 | 0.650.06 | 0.490.07 | 0.520.04 | 0.350.03 | 1.040.03 | T1 | T1.1 |
J1316+0755 | 0.400.02 | 0.350.03 | bbMissing spectral data or skewed by excess noise. | 0.130.02 | 0.260.04 | bbMissing spectral data or skewed by excess noise. | 0.690.05 | 0.100.01 | 0.050.01 | 0.510.05 | T7 | T5.7 |
J13380229 | 0.860.06 | 1.060.08 | 0.830.07 | 0.680.05 | 0.620.05 | 0.700.07 | 0.440.04 | 0.430.03 | 0.280.02 | 1.120.07 | T0 | L9.2: |
J1416+1348B | 0.300.01 | 0.210.01 | 0.000.18 | 0.040.01 | 0.170.01 | 0.350.15 | 0.600.01 | 0.070.01 | 0.040.01 | 0.420.02 | T7 | T7.0 |
J1553+6934 | 0.860.03 | 0.630.02 | 0.650.04 | 0.420.03 | 0.660.03 | 0.860.07 | 0.560.05 | 0.270.01 | 0.160.01 | 0.890.02 | T4 | T3.6 |
J18101010 | 0.980.05 | 0.750.05 | bbMissing spectral data or skewed by excess noise. | 0.530.02 | 1.060.07 | bbMissing spectral data or skewed by excess noise. | 0.740.01 | 0.370.05 | 0.320.04 | 0.950.05 | T3 | T2.2 |
J21056235 | 0.800.01 | 0.850.01 | bbMissing spectral data or skewed by excess noise. | 0.550.01 | 0.600.01 | bbMissing spectral data or skewed by excess noise. | 0.530.01 | bbMissing spectral data or skewed by excess noise. | bbMissing spectral data or skewed by excess noise. | 0.980.01 | T2 | T1.8 |
Note. — Spectral indices are defined as flux ratios, computed as the integral of the spectrum over the defined wavelength ranges in m. For the [H-dip] index, the denominator is the median of the fluxes computed in the two specified ranges. Further details are provided in Burgasser et al. (2006a, b) and Bardalez Gagliuffi et al. (2019).
4.3 K I Equivalent Widths
The moderate resolution of a subset of our spectra facilitates measurement of line equivalent widths (EWs) of the K I doublets at 1.1692/1.1778 m and 1.2436/1.2526 m (Figure 3). We measured these lines by integrating the unsmoothed spectral flux density relative to a linear fit of the local continuum over 0.02 m windows outside the line regions, and sampled both flux uncertainties and small variations in the integration window through Monte Carlo methods. Table 5 lists the measured EWs. In cases where measurements were less than three times their uncertainties, we report 3 upper limits.
Name | SpTaaFinal spectral type as discussed in Section 6. | Instrument | 1.1692 | 1.1778 | 1.2436 | 1.2526 |
(Å) | (Å) | (Å) | (Å) | |||
Metallicity Benchmarks | ||||||
LHS 6176B | d/sdT7.5 | NIRES | 16.02.4 | 9 | 1.0 | 1.2 |
HD 126053B | sdT7.5 | GNIRS | 63 | 93 | 2.5 | 2.0 |
GJ 576B | d/sdT5.5 | XS | 7 | 13.32.2 | 1.3 | 1.5 |
NIRES | 14.72.2 | 10.21.5 | 1.60.2 | 1.80.2 | ||
Wolf 1130C | (e)sdT6: | Keck | 19 | 14 | 1.7 | 1.6 |
Subdwarf Candidates | ||||||
J0045+7958 | d/sdL9 | TSpec | 25 | 6.31.1 | 6.91.2 | 7.81.0 |
J0055+5947 | d/sdT6.5 | NIRES | 35 | 26 | 2.1 | 1.7 |
J0140+0150 | rT4.5 | NIRES | 12.81.4 | 13.01.5 | 7.10.5 | 8.90.5 |
J0411+4714 | d/sdT7.5 | NIRES | 14.82.5 | 8 | 1.0 | 1.1 |
J0429+3201 | T1 | NIRES | 8.30.9 | 8.50.8 | 1.7 | 3.60.5 |
J0433+1009 | T8 | NIRES | 41 | 65 | 1.6 | 3.30.4 |
J0623+0715 | sdT3 | NIRES | 6.71.2 | 5.50.8 | 1.1 | 1.2 |
J0659+1615 | T1 | NIRES | 5.40.6 | 7.30.5 | 2.20.3 | 4.60.3 |
J0758+5711 | T6.5 | NIRES | 12.41.9 | 7.71.5 | 2.00.3 | 2.10.3 |
J11101747 | rT8 | NIRES | 12 | 22 | 1.0 | 1.7 |
J1130+3139 | d/sdT5.5 | NIRES | 13.92.0 | 8.91.5 | 2.00.3 | 0.8 |
J1138+7212 | d/sdT7 | GNIRS | 67 | 50 | 1.3 | 1.2 |
J12042359 | d/sdT7 | NIRES | 163 | 153 | 3.20.4 | 1.6 |
J1304+2819 | d/sdL9: | NIRES | 4.80.6 | 9.50.6 | 6.50.6 | 5.30.7 |
J13080321 | d/sdT8 | NIRES | 51 | bbAffected by excess noise in local continuum. | 1.1 | 1.8 |
J1401+4325 | d/sdT5.5 | NIRES | 18 | 9 | 1.6 | 1.8 |
J1458+1734 | T8 | NIRES | 47 | bbAffected by excess noise in local continuum. | 1.4 | 1.5 |
J15242620 | sdT0 | NIRES | 8.91.1 | 9.40.9 | 5.00.7 | 3.70.7 |
J1710+4537 | T6 | NIRES | 143 | 11.11.9 | 1.0 | 1.0 |
J1801+4717 | d/sdT5 | NIRES | 173 | 9.81.7 | 1.4 | 1.7 |
J20130326 | d/sdT6 | NIRES | 10 | 12.62.3 | 1.2 | 1.1 |
J2021+1524 | d/sdL9 | NIRES | 9.70.7 | 12.40.7 | 6.30.6 | 8.00.5 |
J21120529 | sdT1 | NIRES | 7.90.8 | 10.40.6 | 5.30.5 | 5.90.4 |
J2112+3030 | d/sdT2.5 | NIRES | 5.80.9 | 7.00.8 | 1.1 | 3.00.4 |
J2218+1146 | d/sdT6.5 | NIRES | 12 | 8 | 3.20.4 | 1.4 |
J22510740 | d/sdT7 | NIRES | 47 | bbAffected by excess noise in local continuum. | 1.7 | 1.8 |
Metal-poor Comparison Sources | ||||||
J0021+1552 | d/sdT4 | XS | 76 | 6 | 2.3 | 1.9 |
J0532+8246 | esdL8: | NIRES | 10.10.7 | 13.40.7 | 7.00.7 | 6.30.4 |
J06166407 | esdT0: | XS | 7.30.7 | 10.90.8 | 7.80.8 | 7.20.4 |
J06456646 | d/sdT0 | XS | 10.60.6 | 12.20.5 | 6.80.5 | 8.30.3 |
J0833+0052 | d/sdT9 | GNIRS | 54 | 65 | 8 | 5 |
J1019-3911 | T4 | ARCoIRIS2 | 9.01.1 | 6.30.9 | 3.00.4 | 0.9 |
J1055+5443 | Y0 | NIRES | bbAffected by excess noise in local continuum. | bbAffected by excess noise in local continuum. | 4 | 3 |
J11301158 | sdT5.5 | ARC | 57 | 12 | 3 | 3 |
J1307+1510 | T1 | XS | 9 | 10 | 8.11.0 | 8.71.0 |
J1316+0755 | sdT6.5 | GNIRS | 62 | 11 | 2.5 | 2.5 |
J1553+6933 | sdT4 | NIRES | 5 | 4 | 4 | 3 |
J18101010 | esdT3: | TSpec | 5 | 4 | 3 | 4 |
In Figure 5, we compare these measurements to a compilation of local L and T dwarf spectra from Martin et al. (2017), which is based on 2000 data from the Keck Near-InfraRed SPECtrometer (NIRSPEC; McLean et al. 1998). The dwarf data follow previously-identified trends for K I absorption, with 1.1692 m and 1.1778 m EWs rising from late-L to late-T (largely driven by the drop in continuum as H2O absorption increases), 1.2436 m EWs falling over the same range, and 1.2526 m EWs rising from late-L to mid-T then dropping to zero for the latest T dwarfs. In contrast, many of the previously-observed T subdwarfs and metal-poor T benchmarks, and several of our candidates, are either well below these trends or lack detectable K I absorption to limits of 1–2 Å. This is particularly clear for the 1.2526 m line, which sits near the -band flux peak and shows the greatest dichotomy between dwarfs and subdwarfs for types T1–T7. One notable exception in our candidate sample is J0140+0150, which has exceptionally strong absorption at 1.2436 m for its T4.5 classification.
4.4 Spectral Model Fits
Model | [M/H] | C/O | Other Parameters | Ref | |||
(K) | (cm/s2) | (dex) | (cm2/s) | ||||
LOWZ | 500–1600 | 4.5–5.25 | 2.5 to +1.0 | 0.10–0.85 | 1 to +10 | [1] | |
B06 | 700–2200 | 4.5–5.5 | 0.5 to +0.5 | Cloud model f100 | [2] | ||
Settl | 500–2200 | 4.5–5.5 | 0.5 to +0.5 | [/Fe] = 0 to +0.2 | [3] | ||
Dusty | 1000–2200 | 4.5–5.5 | 2.5 to +0.0 | [/Fe] = 0 to +0.6 | [3] | ||
ATMOaaATMO models are further constrained to the K I broadening profile from Allard et al. (2016) and the condensate rainout prescription described by Goyal et al. (2019); see Phillips et al. (2020). | 600–2000 | 4.5–5.5 | 0.0 | 4–8, CEbbChemical Equilibrium (no mixing). | = 8–0; = 1.0–1.1 | [4] | |
Sonora | 500–2200 | 4.5–5.5 | 0.5 to +0.5 | 0.22–0.83 | [5] | ||
ElfOwl | 525–2200 | 4.5–5.5 | 1.0 to +1.0 | 0.22–0.83 | 2–8 | [6] | |
SAND | 700–2200 | 4.5–6.0 | 2.4 to +0.3 | [/Fe] = 0 to +0.4 | [7] |
To obtain initial estimates of the physical properties of our sources, we compared low-resolution versions of our spectra to eight sets of spectral models from Burrows et al. (2006, hereafter B06), Allard et al. (2012, hereafter Settl and Dusty), Phillips et al. (2020, hereafter ATMO), Marley et al. (2021, hereafter Sonora), Meisner et al. (2021, hereafter LOWZ), Mukherjee et al. (2024, hereafter ElfOwl), and Alvarado et al. (2024, hereafter SAND). The parameters and parameter ranges explored for these models are summarized in Table 4.4. These models were chosen as they span both low temperatures and solar and subsolar metallicities, with the LOWZ models currently providing the best metallicity coverage for T subdwarf atmospheres. The observed () and model () spectra were resampled onto a common wavelength scale with a constant resolution of = 150 over the range 0.9–2.45 m using a median filter,444Smoothed uncertainties were computed as median({})/ for the uncertainty values {} within the sampling region. and noisy data at the edge of the observed spectral range and in the 1.35–1.45 and 1.8–1.95 telluric bands were masked from the fit. To identify the best-fitting model, we used a goodness-of-fit statistic,
where is the number of unmasked flux values, is the observed uncertainty spectrum, and the optimal scale factor is computed as
(cf. Cushing et al. 2008). We report here the reduced chi-square, , as a normalized quality of fit. For sources with measured parallaxes, we scaled the observed spectra to absolute fluxes using their magnitudes, which allows us to estimate the source radius from the scale factor
All other spectra are scaled to apparent magnitudes. No attempt was made to adjust for relative velocity shifts between the observed and model spectra, which are expected to be minimal at this low resolution.
For each model set, an initial fit was made to the individual grid models, and the best-fit (lowest ) model was used as an initial estimate of the atmosphere parameters. We employed a simple Metropolis-Hastings Markov Chain Monte Carlo (MCMC) algorithm (Metropolis et al., 1953; Hastings, 1970) to explore the proximate parameter space. We used a single chain of 2,000 steps for each spectrum and model set comparison, varying the continuous variables of effective temperature (), surface gravity (), and solar-scaled metallicity ([M/H]) for all models, as well as carbon/oxygen abundance ratio (C/O) for LOWZ, Sonora, and ElfOwl models; alpha element enrichment ([/Fe]) for Settl, Dusty, and SAND models; and adiabatic coefficient (Tremblin et al., 2019) and minimum cloud deck pressure for ATMO models. We also varied the disequilibrium mixing diffusion rate for the LOWZ, ATMO, and ElfOwl models as a discrete value. Models were linearly interpolated in logarithmic flux units on a logarithmic parameter grid (i.e,. ). In each iteration, we drew new parameters for the continuous variables using a normal distribution centered on the last parameter values in the chain with pre-defined standard deviations.555These standard deviations were = 30 K, = 0.1 dex, = 0.05 dex, = 0.05 dex, = 0.05 dex, and = 0.01. The discrete variable was randomly drawn from the available values. Sequential fits () were compared using the criterion
where MIN is the minimum of all values in the chain and is a number drawn from a uniform distribution between 0 and 0.5. If the new fit satisfied this criterion, these parameters were added to the chain; otherwise the previous parameters were added. We also enforced a limit on values exceeding 2MIN[], at which point the chain reverted to the minimum parameter set. We verified that all fits converged using the convergence diagnostic defined by Geweke (1992), requiring that the means of the variable parameters in the first 10% and last 50% of each chain differ by less than three times their combined variance. We also visually confirmed convergence, and examined each fit to ensure it provided an accurate reproduction of the data.
Name | SpT | Instrument | [M/H] | C/O | R | MIN | Model | |||
(K) | (cm/s2) | (dex) | (cm2/s) | (R⊙) | ||||||
Metallicity Benchmarks | ||||||||||
HD 3651B | T8 | SpeX | 856 | 4.56 | 0.01 | 0.48 | 1 | 0.070 | 21 | LOWZ |
834 | 4.53 | 0.00 | 0.67 | 4 | 0.077 | 15 | ElfOwl | |||
HD 65486B | d/sdT4.5 | SpeX | 993 | 4.98 | 0.53 | 0.38 | 1 | 0.099 | 5 | LOWZ |
LHS 6176B | d/sdT7.5 | NIRES | 846 | 4.52 | 0.21 | 0.49 | 1 | 0.061 | 15 | LOWZ |
Ross 458C | rT8.5 | FIRE | 790 | 4.61 | 0.95 | 0.38 | 1 | 0.080 | 319 | LOWZ |
812 | 4.61 | 0.78 | 1.11 | 2 | 0.075 | 298 | ElfOwl | |||
HD 126053B | sdT7.5 | GNIRS | 654 | 4.98 | 0.35 | 0.56 | 2 | 0.076 | 40 | LOWZ |
746 | 4.23 | 0.53 | 1.20 | 2 | 0.055 | 33 | ElfOwl | |||
GJ 570D | T7.5 | SpeX | 837 | 4.62 | 0.11 | 0.54 | 1 | 0.061 | 10 | LOWZ |
GJ 576B | sdT5.5 | GNIRS | 912 | 4.71 | 0.47 | 0.37 | 1 | 0.096 | 6 | LOWZ |
XS | 912 | 4.74 | 0.53 | 0.32 | 1 | 0.098 | 4 | LOWZ | ||
NIRES | 966 | 5.25 | 0.24 | 0.39 | -1 | 0.080 | 62 | LOWZ | ||
995 | 5.19 | 0.21 | 0.86 | 0.078 | 29 | Sonora | ||||
G 20439B | T6.5 | SpeX | 984 | 4.76 | 0.43 | 0.46 | 1 | 0.069 | 26 | LOWZ |
Wolf 1130C | (e)sdT6: | Multiple | 978 | 5.03 | 0.29 | 0.60 | 10 | 0.018 | 97 | LOWZ |
700aaReaches parameter limit for given model set. | 5.57 | 0.65 | 0.21bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.056 | 76 | SAND | ||||
HN PegB | T2.5 | SpeX | 1284 | 4.76 | 0.11 | 0.32 | 1 | 0.074 | 90 | LOWZ |
Wolf 940B | T8.5 | FIRE | 636 | 5.20 | 0.33 | 0.81 | 10 | 0.069 | 6 | LOWZ |
Subdwarf Candidates | ||||||||||
J0045+7958 | d/sdL9 | TSpec | 1600aaReaches parameter limit for given model set. | 4.73 | 0.79 | 0.64 | 1 | 6 | LOWZ | |
1637 | 5.47 | 0.05 | 0.21bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 4 | SAND | |||||
J0055+5947 | d/sdT6.5 | NIRES | 641 | 5.00 | 0.11 | 0.73 | 1 | 0.143 | 30 | LOWZ |
738 | 5.07 | 0.04 | 0.73 | 4 | 0.094 | 18 | ElfOwl | |||
J0140+0150 | rT4.5 | SpeX | 1149 | 4.77 | 0.21 | 0.49 | 1 | 4 | LOWZ | |
NIRES | 1148 | 5.06 | 0.40 | 0.48 | 1 | 9 | LOWZ | |||
J0411+4714 | d/sdT7.5 | NIRES | 860 | 4.63 | 0.20 | 0.59 | 1 | 10 | LOWZ | |
843 | 4.54 | 0.12 | 0.68 | 4 | 4 | ElfOwl | ||||
J0429+3201 | T1 | NIRES | 1598aaReaches parameter limit for given model set. | 4.74 | 0.28 | 0.59 | -1 | 12 | LOWZ | |
1648 | 5.36 | 0.08 | 11 | B06 | ||||||
J0433+1009 | T8 | NIRES | 953 | 4.80 | 0.12 | 0.60 | 1 | 4 | LOWZ | |
J0623+0715 | sdT3 | NIRES | 1277 | 4.81 | 0.68 | 0.49 | 1 | 6 | LOWZ | |
J0659+1615 | T1 | NIRES | 1600aaReaches parameter limit for given model set. | 5.20 | 0.03 | 0.48 | 2 | 21 | LOWZ | |
1536 | 5.72 | 0.03 | 0.08bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 6 | SAND | |||||
J0758+5711 | T6.5 | NIRES | 953 | 4.66 | 0.17 | 0.47 | 1 | 6 | LOWZ | |
J0843+2904 | L6: | SpeX | 1600aaReaches parameter limit for given model set. | 4.69 | 0.79 | 0.66 | 1 | 12 | LOWZ | |
1663 | 5.47 | 0.01 | 0.08bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 2 | SAND | |||||
J08453305 | d/sdT6.5 | FIRE | 933 | 4.51 | 0.23 | 0.43 | 1 | 173 | LOWZ | |
952 | 4.50 | 0.35 | 0.53 | 2 | 121 | ElfOwl | ||||
J0911+2146 | d/sdT8 | FIRE | 746 | 4.70 | 0.18 | 0.67 | 1 | 68 | LOWZ | |
750 | 4.55 | 0.00 | 0.56 | 4 | 34 | ElfOwl | ||||
J11101747 | rT8 | NIRES | 909 | 4.74 | 0.20 | 0.67 | 1 | 26 | LOWZ | |
J1130+3139 | d/sdT5.5 | NIRES | 948 | 5.20 | 0.21 | 0.37 | 1 | 12 | LOWZ | |
1033 | 4.72 | 0.17 | 0.53 | 4 | 7 | ElfOwl | ||||
J1138+7212 | d/sdT7 | GNIRS | 854 | 4.62 | 0.37 | 0.54 | 1 | 0.113 | 9 | LOWZ |
900 | 4.54 | 0.36 | 1.05 | 2 | 0.102 | 5 | ElfOwl | |||
J12042359 | d/sdT7 | NIRES | 930 | 4.73 | 0.15 | 0.47 | 1 | 3 | LOWZ | |
J1304+2819 | d/sdL9: | NIRES | 1600aaReaches parameter limit for given model set. | 4.63 | 0.89 | 0.63 | 1 | 33 | LOWZ | |
1803 | 5.57 | 0.20 | 0.06bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 8 | SAND | |||||
J13080321 | d/sdT8 | NIRES | 827 | 4.55 | 0.11 | 0.50 | 1 | 10 | LOWZ | |
J1401+4325 | d/sdT5.5 | NIRES | 896 | 5.19 | 0.13 | 0.39 | 1 | 4 | LOWZ | |
J1458+1734 | T8 | NIRES | 854 | 4.61 | 0.06 | 0.72 | 1 | 8 | LOWZ | |
790 | 4.14 | 0.13 | 0.75 | 4 | 3 | ElfOwl | ||||
J15152157 | d/sdT6 | FIRE | 832 | 4.66 | 0.06 | 0.57 | 10 | 62 | LOWZ | |
952 | 4.68 | 0.08 | 0.87 | 4 | 36 | ElfOwl | ||||
J15242620 | sdT0 | NIRES | 1600aaReaches parameter limit for given model set. | 4.81 | 0.38 | 0.44 | 1 | 63 | LOWZ | |
1565 | 5.35 | 0.69 | 0.20bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 9 | SAND | |||||
J1710+4537 | T6 | NIRES | 984 | 5.17 | 0.02 | 0.45 | 1 | 3 | LOWZ | |
J1801+4717 | d/sdT5 | NIRES | 973 | 5.13 | 0.09 | 0.36 | 1 | 3 | LOWZ | |
J20130326 | d/sdT6 | NIRES | 956 | 4.86 | 0.21 | 0.42 | 1 | 3 | LOWZ | |
J2021+1524 | d/sdL9 | NIRES | 1600aaReaches parameter limit for given model set. | 4.66 | 0.14 | 0.46 | 1 | 61 | LOWZ | |
1611 | 5.59 | 0.35 | 0.34bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 8 | SAND | |||||
J21120529 | sdT1 | NIRES | 1594aaReaches parameter limit for given model set. | 4.63 | 0.92 | 0.15 | 1 | 27 | LOWZ | |
1520 | 5.63 | 0.74 | 0.38bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 4 | SAND | |||||
J2112+3030 | d/sdT2.5 | NIRES | 1325 | 5.00 | 0.43 | 0.40 | 1 | 6 | LOWZ | |
J2218+1146 | d/sdT6.5 | NIRES | 890 | 5.22 | 0.10 | 0.55 | 1 | 3 | LOWZ | |
J22510740 | d/sdT7 | NIRES | 930 | 4.71 | 0.00 | 0.48 | 1 | 5 | LOWZ | |
Metal-poor Comparison Sources | ||||||||||
J00041336 | d/sdT2: | SpeX | 1567aaReaches parameter limit for given model set. | 4.63 | 0.83 | 0.14 | 1 | 9 | LOWZ | |
2011 | 4.80 | 0.56 | 0.20bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 7 | SAND | |||||
J00042604 | T3 | SpeX | 1325 | 4.83 | 0.40 | 0.36 | 2 | 0.087 | 19 | LOWZ |
J0013+0634 | sdT7.5 | GNIRS | 658 | 4.91 | 0.00 | 0.23 | 10 | 105 | LOWZ | |
727 | 5.03 | 0.61 | 0.18bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 80 | Settl | |||||
J0021+1552 | d/sdT4 | XS | 1149 | 4.88 | 0.49 | 0.37 | 1 | 2 | LOWZ | |
J0057+2013 | d/sdL9 | SpeX | 1590aaReaches parameter limit for given model set. | 4.73 | 0.24 | 0.49 | 10 | 30 | LOWZ | |
1859 | 4.80 | 0.15 | 0.10bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 14 | Settl | |||||
J03012319 | d/sdT1 | SpeX | 1550aaReaches parameter limit for given model set. | 4.79 | 0.65 | 0.26 | 1 | 0.056 | 18 | LOWZ |
1488 | 5.65 | 0.44 | 0.18bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.063 | 12 | SAND | ||||
J03095016 | d/sdT7.5 | FIRE | 747 | 4.61 | 0.30 | 0.57 | 2 | 0.087 | 48 | LOWZ |
795 | 4.13 | 0.36 | 0.70 | 2 | 0.077 | 38 | ElfOwl | |||
J03485620 | d/sdT4 | FIRE | 1195 | 5.25 | 0.00 | 0.54 | 2 | 15 | LOWZ | |
J04145854 | esdT6: | FIRE | 699 | 4.88 | 1.34 | 0.21 | 1 | 4 | LOWZ | |
J0532+8246 | esdL8: | Multiple | 1600aaReaches parameter limit for given model set. | 4.56 | 1.56 | 0.11 | -1 | 0.088 | 57 | LOWZ |
1718 | 5.05 | 1.53 | 0.33bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.077 | 9 | SAND | ||||
J06166407 | esdT0: | XS | 1598aaReaches parameter limit for given model set. | 4.58 | 1.77 | 0.11 | 2 | 0.106 | 21 | LOWZ |
2181 | 5.20 | 1.13 | 0.07bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.056 | 6 | Dusty | ||||
J06456646 | d/sdT0 | XS | 1595aaReaches parameter limit for given model set. | 4.54 | 0.63 | 0.33 | 1 | 0.033 | 73 | LOWZ |
1594 | 5.49 | 0.46 | 0.29bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.063 | 65 | SAND | ||||
J0833+0052 | d/sdT9 | GNIRS | 528 | 5.03 | 0.08 | 0.42 | 2 | 0.050 | 13 | LOWZ |
J08500221 | d/sdL6 | SpeX | 1600aaReaches parameter limit for given model set. | 4.57 | 0.24 | 0.64 | 10 | 0.131 | 175 | LOWZ |
1852 | 5.53 | 0.17 | 0.05bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.099 | 46 | SAND | ||||
J0937+2931 | sdT6 | SpeX | 859 | 4.70 | 0.59 | 0.49 | 1 | 0.088 | 33 | LOWZ |
J09392448 | d/sdT8 | SpeX | 682 | 4.75 | 0.35 | 0.73 | 1 | 0.080 | 6 | LOWZ |
J09530943 | d/sdT6 | SpeX | 881 | 4.81 | 0.43 | 0.52 | 1 | 7 | LOWZ | |
J10193911 | T4 | ARC | 1200 | 5.25 | 0.00 | 0.55 | 1 | 58 | LOWZ | |
J10350711 | d/sdL9 | SpeX | 1600aaReaches parameter limit for given model set. | 4.65 | 0.35 | 0.37 | 10 | 29 | LOWZ | |
1695 | 5.59 | 0.01 | 0.17bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 10 | SAND | |||||
J1055+5443 | Y0 | NIRES | 670 | 4.77 | 0.70 | 0.34 | 2 | 0.021 | 8 | LOWZ |
J11301158 | sdT5.5 | ARC | 918 | 4.86 | 0.56 | 0.44 | 1 | 25 | LOWZ | |
J1158+0435 | d/sdL8 | SpeX | 1600aaReaches parameter limit for given model set. | 4.66 | 0.56 | 0.25 | 1 | 0.085 | 112 | LOWZ |
1870 | 5.50 | 0.19 | 0.09bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.062 | 30 | SAND | ||||
J1307+1510 | T1 | XS | 1549aaReaches parameter limit for given model set. | 4.79 | 0.14 | 0.39 | 2 | 3 | LOWZ | |
1633 | 5.45 | 0.42 | 0.35bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 1 | SAND | |||||
J1316+0755 | sdT6.5 | GNIRS | 765 | 4.69 | 0.65 | 0.29 | 1 | 24 | LOWZ | |
J13380229 | sdL9 | FIRE | 1592aaReaches parameter limit for given model set. | 4.60 | 0.71 | 0.25 | 1 | 2 | LOWZ | |
1639 | 5.51 | 0.50 | 0.22bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 1 | SAND | |||||
J1416+1348B | sdT7 | SpeX | 635 | 4.62 | 0.67 | 0.38 | 1 | 0.102 | 11 | LOWZ |
J1553+6933 | sdT4 | NIRES | 1225 | 4.78 | 0.96 | 0.42 | 1 | 2 | LOWZ | |
J18101010 | esdT3: | Multiple | 1300 | 5.00 | 0.91 | 0.34 | -1 | 0.020 | 32 | LOWZ |
869 | 5.87 | 1.27 | 0.34bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.081 | 9 | SAND | ||||
J21056235 | d/sdT2: | Flamingos | 1282 | 4.72 | 0.32 | 0.66 | 10 | 65 | LOWZ |
Note. — Model parameters are listed for best-fit LOWZ models by default (Meisner et al., 2021), Sources that reached the LOWZ parameter limit or had a significantly better fits ( 5) to an alternate model are also listed with the alternative model parameters. Model names are given in shorthand notation as B06 (Burrows et al., 2006, assuming cloud model f100), Settl and Dusty (Allard et al., 2012), ATMO (Phillips et al., 2020), Sonora (Marley et al., 2021), ElfOwl (Mukherjee et al., 2024), or or SAND (Alvarado et al., 2024). Instrument names are given in shorthand notation as ARC = SOAR/ARCoIRIS, TSpec = APO/TripleSpec or Palomar/TripleSpec, and XS = VLT/X-Shooter. Multiple instruments contribute to spectra used for Wolf 1130C (Keck/NIRSPEC and Keck/MOSFIRE), J0532+8246 (Keck/LRIS and Keck/NIRES), and J18101010 (GTC/OSIRIS, GTC/EMIR, and Palomar/TripleSpec)
Table 7 lists the best-fit model parameters for each of the spectra in our sample, while Figures 6 through 9 display model fits for representative sources. We find excellent agreement with the LOWZ models for most of our spectra, with temperatures appropriate for their spectral classifications. In particular, we find good agreement between the metallicities inferred for our benchmark companions and the metallicities reported for their primaries (Figure 6). Nearly all of the sources with measured parallaxes also have inferred radii consistent with evolved brown dwarfs, with a few cases where radii are either too large (e.g., J0055+5947) or too small (e.g., J1316+0755 and J18101010), likely reflecting inferred temperatures that are too low or too high, respectively. Fits to three sources (GJ 576B, J04145854, and J1316+0755) yield unusually small C/O ratios, which likely offset low temperatures since low C/O reduces the relative strength of CH4 absorption. The poorest fits for the LOWZ models occur for the late-L and early-T dwarfs in our sample, a consequence of the = 1600 K maximum for this model set.
For sources that hit the parameter limits for the LOWZ models or have better fits among the other seven model sets ( 5), we also report in Table 7 the parameters inferred for the best fits among all models. For the highest-temperature sources, these are typically from the SAND models which yield higher temperatures outside the LOWZ parameter range, and generally higher surface gravities and lower metallicities. An illustrative case is the extreme L subdwarf J0532+8246 (Figure 7) for which the SAND models provide an excellent fit to strong FeH absorption at 1 m and the smooth near-infrared continuum, while inferring a radius (R = 0.077 R⊙) in line with theoretical models. On the other, other SAND model fits to sources with parallax measurements (e.g., J1158+0435) yield radii too small to be physical, suggesting that the inferred temperatures are too high in these cases. SAND models also provide improved fits for some of the lower-temperature metal-poor sources, including Wolf 1130C for which the inferred metallicity ([M/H] = 0.65) is consistent with the metallicity of the system’s M dwarf primary ([Fe/H] = 0.700.12; Mace et al. 2018); and the extreme T subdwarf J18101010 (Figure 8), for which the inferred temperature ( = 869 K) and radius (0.081 R⊙) are consistent with luminosity-based estimates from Lodieu et al. (2022). The better agreement in near-infrared spectral morphology and inferred parameters for these sources suggest that the SAND models are well-suited for metal-poor L and T subdwarfs. Other low-temperature T dwarfs with solar/near-solar metallicities are better fit by the ElfOwl models, which have up-to-date opacities for key molecules such as CH4 and thus better reproduce the strong absorption features present in late-type T dwarf spectra. Nevertheless, these fits yield similar atmosphere parameters as LOWZ ( 100 K, 0.5, 0.2).
Turning to our subdwarf candidates, several show evidence of modest subsolar metallicity (0.2 [M/H] 0.5) based on the model fits, with three sources—J0623+0715, J15242620, and J21120529—being the most metal-poor ([M/H] 0.7). Remarkably, the candidate J0140+0150 and the previously reported J1055+5443 (Robbins et al., 2023) are found to be significantly metal-rich ([M/H] +0.4) despite their selection as high velocity objects. These sources are discussed in further detail in Section 5.
4.5 Radial Velocities
The moderate resolution of our Keck/NIRES data facilitates measurement of the radial velocities of our targets via the forest of H2O, CH4, and (for late-L and early T dwarfs) CO molecular features that are marginally resolved in the data. Our approach is based on the forward modeling technique using pre-telluric-corrected spectra, as described in Burgasser et al. (2015) and Hsu et al. (2021). We performed separate fits in -band (1.11–1.16 m), -band (1.26–1.31 m), -band (1.52–1.59 m), and -band (2.03–2.23 m), regions that encompass strong molecular features in both source and telluric spectra. We also conducted additional fits in the 1.235–1.285 m, 1.95–2.40 m, and 2.26–2.40 m regions for our earliest-type sources, which encompass K I and CO absorption. Our spectral data model was defined as
where is a high-resolution ( = 50,000) BT-Settl atmosphere model (Allard et al., 2012), with = 5.0 (cgs) and based on the fitting values in Table 7. This model is evaluated at wavelengths where is the barycentric velocity correction for the source at the time of observation, is the heliocentric velocity constrained to be within 200 km/s, and is the speed of light. is an empirical telluric absorption template from Livingston & Wallace (1991) scaled by an exponent to account for varying airmass. is a fifth-order polynomial continuum correction to the ratio of smoothed versions of the data and the combination . is a constant flux offset. is a Gaussian instrumental broadening profile of width , where is the average wavelength over the fitting range and is a free parameter. As is typically of order 100 km/s, we ignore stellar rotational broadening in this analysis. We also fit for a constant wavelength offset between the wavelength grid of the data and that of the data model to account for slight offsets (up to 50 Å) in the original wavelength calibration. These adjustments significantly improve our RV precisions and motivate our use of pre-telluric-corrected spectra (cf. Hsu et al. 2021).
The data model is defined by five free parameters (RV, , , , ) and six continuum coefficients that are determined at each fitting step. We used a similar reduced statistic as Eqn. 4 to assess the quality of the fit,
where is the number of unmasked data points and . Note that the scaling factor in Eqns. 4 and 5 is incorporated into the continuum correction . The five free parameters were first optimized using the Nelder-Mead algorithm (Nelder, 1965; Gao & Han, 2012) as implemented in the scipy optimize package (Virtanen et al., 2020). These parameters were then used to seed an MCMC model with a single chain of 5,000 steps for each source and wavelength range, with verification of convergence using the Geweke (1992) diagnostic for the RV values and visual confirmation of the chain. We also examined each fit to ensure the presence of distinct spectral structure from both stellar and telluric absorption, and rejected those fits lacking sufficient structure or are in very low signal-to-noise regions.
Figure 10 illustrates fits for GJ 576B, demonstrating how the combination of telluric and stellar absorption are well reproduced by the forward model. For our highest signal-to-noise data, these fits provide robust constraints on RV ( 5 km/s, equivalent to roughly 0.2 pixels at native resolution) and consistency between the four fit regions. For lower signal-to-noise data, RV errors increase to tens of km/s. Variations in the RV fit value across the spectral regions are generally consistent with the measurement uncertainties.
Table 4.5 summarizes the individual fit RVs, with median values and uncertainties computed as weighted 16%, 50%, and 84% quantiles, using the weighting factor
For sources with at least two viable RV measurements, we computed a weighted mean RV and uncertainty by adopting the larger of the lower and upper RV error estimates, , as a weighting factor:
The sums are over the = 2–4 combined measurements, and the form of Eqn. 12 accounts for the size of the individual measurement uncertainties. The final RVs are reported in Table 4.5.
Y | J | H | K | Adopted | Prior | ||||||
Name | RV | RV | RV | RV | RV | RV | Ref | ||||
(km/s) | (km/s) | (km/s) | (km/s) | (km/s) | (km/s) | ||||||
Metallicity Benchmarks | |||||||||||
LHS 6176B | 10 | 4 | 9 | 5 | 2 | 2 | 87 | 11.72.1ccBased on fits to alternative spectral range 2.26-2.40 m. | [1] | ||
GJ 576B | 32 | 7 | 23 | 12 | 24 | 7 | 25 | 5 | (254)ddDue to unusually good seeing at the time of observation, offsets of tens of km/s are possible; value should be treated with skepticism and is reported here only for completeness. | 84.790.12 | [2] |
Subdwarf Candidates | |||||||||||
J0055+5947 | 38 | 1 | 30 | 1 | 27 | 4 | 3425 | ||||
J0140+0150 | 84 | 2 | 96 | 2 | 90 | 2 | 86 | 4 | 917 | ||
J0411+4714 | 37 | 137 | 17 | 175 | 40 | 118 | 348 | ||||
J0429+3201 | 117aaBased on fits to alternative spectral range 1.235–1.280 m. | 2 | 63bbBased on fits to alternative spectral range 1.95–2.40 m. | 6 | (7021)ddDue to unusually good seeing at the time of observation, offsets of tens of km/s are possible; value should be treated with skepticism and is reported here only for completeness. | ||||||
J0433+1009 | 21 | 2 | 15 | 2 | 21 | 2 | 217 | ||||
J0623+0715 | 14 | 2 | 18 | 2 | 2bbBased on fits to alternative spectral range 1.95–2.40 m. | 3 | 127 | ||||
J0659+1615 | 41 | 1 | 46 | 2 | 29 | 2 | 38ccBased on fits to alternative spectral range 2.26-2.40 m. | 7 | 369 | ||
J0758+5711 | 81 | 3 | 75 | 2 | 72 | 2 | 797 | ||||
J11101747 | 46 | 3 | 55 | 5 | 44 | 5 | 475 | ||||
J1130+3139 | 64 | 3 | 76 | 3 | 76 | 2 | 68 | 2 | 728 | ||
J12042359 | 41 | 2 | 76 | 3 | 57 | 2 | 4715 | ||||
J1304+2819 | 45 | 1 | 35 | 2 | 51ccBased on fits to alternative spectral range 2.26-2.40 m. | 6 | 509 | ||||
J13080321 | 40 | 3 | 52 | 3 | 51 | 2 | 426 | ||||
J1401+4325 | 68 | 3 | 79 | 2 | 7121 | ||||||
J1458+1734 | 19 | 4 | 16 | 3 | 24 | 4 | 196 | ||||
J15242620 | 34 | 1 | 28 | 4 | 2917 | ||||||
J1710+4537 | 1 | 3 | 2 | 2 | 13 | 3 | 38 | ||||
J1801+4717 | 27 | 2 | 33 | 2 | 22 | 4 | 2917 | ||||
J20130326 | 20 | 2 | 22 | 2 | 23 | 3 | 1813 | ||||
J2021+1524 | 11 | 1 | 10 | 2 | 7ccBased on fits to alternative spectral range 2.26-2.40 m. | 5 | 57 | ||||
J21120529 | 42 | 2 | 24aaBased on fits to alternative spectral range 1.235–1.280 m. | 2 | 37bbBased on fits to alternative spectral range 1.95–2.40 m. | 4 | 3714 | ||||
J2112+3030 | 125 | 4 | 12562 | ||||||||
J2218+1146 | 23 | 5 | 22 | 3 | 15 | 3 | 43 | 2 | 229 | ||
J22510740 | 133 | 4 | 197 | 3 | 14025 | ||||||
Comparison Sources | |||||||||||
J0532+8246 | 155 | 99 | 205aaBased on fits to alternative spectral range 1.235–1.280 m. | 56 | 128bbBased on fits to alternative spectral range 1.95–2.40 m. | 42 | 16039 | 1721 | [3] | ||
J1055+5443 | 36 | 4 | 3 | 3 | 75 | 4 | 3922 | ||||
J1553+6933 | 110 | 5 | 11090 |
For two of the three sources with previously published RVs, LHS 6176B (from its primary) and J0532+82346, we find excellent agreement with our Keck/NIRES RV measurements. For GJ 576B, we measure a precise RV = 254 km/s that is nevertheless highly discrepant with its primary RV of 84.790.12 km/s (Halbwachs et al., 2018). As the latter value is validated by Gaia DR3 measurements (83.90.2 km/s; Gaia Collaboration et al. 2021), this deviation cannot be explained by unresolved multiplicity in the primary, and is of too large of a magnitude to be explained by unresolved multiplicity of the brown dwarf companion.666The shortest-period very-low-mass binary currently known, LP 413-53AB with = 17 hr, has an RV semi-amplitude of 24 km/s (Hsu et al., 2023). Instead, this discrepancy is likely due to an observational bias. Both LHS 6176B and J0429+3201 were observed in conditions of excellent seeing (04), smaller than the fixed slit width of NIRES. A slight offset of the source from the slit center by a single pixel (015, approximately 1/3 of a resolution element or 40 km/s) would shift the dispersed spectrum by an amount sufficient to explain this discrepancy. Unfortunately, this systematic bias is uncorrectable with the present data. We therefore exclude the RV of this source and that of J0429+3201 from our analysis.
Name | daaDistance based on parallax measurement of source or primary companion, or estimated from Eqns.2 and 3. | RV | U | V | W | Probabilities | Ref | |||||
(pc) | (mas/yr) | (mas/yr) | (km/s) | (km/s) | (km/s) | (km/s) | (km/s) | thin | thick | halo | ||
Metallicity Benchmarks | ||||||||||||
HD 3651BbbAnalysis based on kinematics of the primary component of this system. | 11.1080.006 | 461.950.07 | 369.620.03 | 31.1500.017 | 32.8980.005 | 51.1490.015 | 7.2050.004 | 15.6640.008 | 99% | 1% | 0% | [1],[2] |
HD 65486BbbAnalysis based on kinematics of the primary component of this system. | 18.4650.006 | 362.4090.012 | 245.7860.016 | 38.3260.013 | 8.190.12 | 46.170.05 | 4.240.10 | 22.8160.006 | 99% | 1% | 0% | |
LHS 6176BbbAnalysis based on kinematics of the primary component of this system. | 19.6070.010 | 234.470.03 | 360.660.02 | 39.980.02 | 122 | 38.70.9 | 18.31.3 | 13.71.3 | 99% | 1% | 0% | [1],[3] |
Ross 458CbbAnalysis based on kinematics of the primary component of this system. | 11.5070.015 | 628.720.18 | 33.470.13 | 34.340.05 | 12.30.3 | 18.460.07 | 6.020.07 | 3.90.3 | 99% | 1% | 0% | |
HD 126053BbbAnalysis based on kinematics of the primary component of this system. | 17.4610.011 | 2243.530.05 | 478.280.03 | 43.690.03 | 19.210.14 | 32.780.07 | 3.0520.017 | 32.430.12 | 93% | 7% | 0% | |
GJ 570DbbAnalysis based on kinematics of the primary component of this system. | 5.8860.002 | 1031.470.07 | 1724.620.06 | 56.040.02 | 26.9870.002 | 59.360.01 | 9.7920.006 | 25.2440.019 | 93% | 7% | 0% | [1],[2] |
GJ 576BbbAnalysis based on kinematics of the primary component of this system. | 19.0460.009 | 607.630.02 | 506.510.02 | 71.410.03 | 84.790.12 | 47.250.07 | 62.790.03 | 49.810.09 | 0% | 99% | 0% | [1],[4] |
G 20439BbbAnalysis based on kinematics of the primary component of this system. | 13.9870.003 | 15.070.02 | 578.140.02 | 38.3430.008 | 31.1500.002 | 33.7040.008 | 7.1640.002 | 0.2880.002 | 99% | 1% | 0% | [1],[5] |
Wolf 1130CbbAnalysis based on kinematics of the primary component of this system. | 16.5850.007 | 1159.520.04 | 904.010.03 | 115.580.05 | 33.21.3ccCenter of mass motion of the Wolf 1130AC binary based on orbit fits (Mace et al., 2018). | 117.460.05 | 31.71.2 | 42.20.2 | 18% | 81% | 0% | [1],[6] |
HN PegBbbAnalysis based on kinematics of the primary component of this system. | 18.1330.011 | 231.110.03 | 113.200.03 | 22.120.014 | 16.72300.0013 | 3.3400.006 | 8.8480.005 | 3.5100.012 | 99% | 1% | 0% | [1],[2] |
Wolf 940BbbAnalysis based on kinematics of the primary component of this system. | 12.3860.008 | 769.370.05 | 505.680.03 | 54.050.03 | 28.40.3 | 21.270.12 | 33.110.18 | 17.720.17 | 92% | 8% | 0% | |
Subdwarf Candidates | ||||||||||||
J0055+5947bbAnalysis based on kinematics of the primary component of this system. | 22.840.03 | 40825 | 2425 | 443 | 3425 | 4415 | 1623 | 43 | 96% | 4% | 0% | [1],[7] |
J0140+0150 | 272 | 97113 | 9019 | 12311 | 917 | 1178 | 458 | 537 | 0% | 99% | 1% | [7] |
J0411+4714 | 222 | 44025 | 14326 | 495 | 348 | 408 | 136 | 263 | 97% | 3% | 0% | [7] |
J0429+3201ddRV and velocities may be systematically biased by slit-seeing effect and should be treated with caution. | 353 | 3031 | 45831 | 758 | (7021) | (7521) | (628) | (257) | 15% | 85% | 0% | [7] |
J0433+1009 | 293 | 17434 | 38436 | 587 | 217 | 36 | 467 | 126 | 83% | 17% | 0% | [7] |
J0623+0715 | 252 | 732 | 98037 | 11711 | 127 | 5622 | 8212 | 466 | 0% | 99% | 1% | [7] |
J0659+1615 | 313 | 1158 | 4458 | 686 | 369 | 559 | 275 | 424 | 43% | 57% | 0% | [7] |
J0758+5711 | 272 | 12031 | 34233 | 466 | 797 | 756 | 35 | 345 | 84% | 16% | 0% | [7] |
J11101747 | 262 | 2428 | 39527 | 495 | 475 | 374 | 235 | 555 | 28% | 71% | 0% | [7] |
J1130+3139 | 323 | 118535 | 147638 | 28425 | 728 | 828 | 26524 | 328 | 0% | 0% | 100% | [7] |
J12042359 | 313 | 26337 | 47338 | 799 | 4715 | 66 | 411 | 8210 | 1% | 98% | 0% | [7] |
J1304+2819 | 313 | 35619 | 1822 | 525 | 509 | 314 | 134 | 609 | 49% | 51% | 0% | [7] |
J13080321 | 212 | 45426 | 18929 | 505 | 426 | 495 | 164 | 165 | 96% | 4% | 0% | [7] |
J1401+4325 | 343 | 17233 | 65934 | 11011 | 7121 | 628 | 10411 | 2521 | 0% | 98% | 2% | [7] |
J1458+1734 | 323 | 47620 | 16317 | 757 | 196 | 376 | 144 | 626 | 38% | 62% | 0% | [7] |
J15242620 | 404 | 37238 | 10745 | 7310 | 2917 | 315 | 5710 | 4410 | 35% | 65% | 0% | [7] |
J1710+4537 | 363 | 40436 | 8041 | 699 | 38 | 128 | 269 | 6110 | 39% | 61% | 0% | [7] |
J1801+4717 | 484 | 10956 | 47863 | 11117 | 2917 | 10218 | 1917 | 1015 | 56% | 44% | 0% | [7] |
J20130326 | 293 | 64939 | 62540 | 12412 | 1813 | 8313 | 8111 | 508 | 1% | 99% | 1% | [7] |
J2021+1524 | 222 | 5473 | 4413 | 726 | 57 | 697 | 257 | 313 | 90% | 10% | 0% | [7] |
J21120529 | 303 | 55624 | 74825 | 13112 | 3714 | 8912 | 9510 | 428 | 0% | 99% | 1% | [7] |
J2112+3030 | 393 | 43115 | 7915 | 808 | 12562 | 2815 | 10261 | 10015 | 0% | 66% | 34% | [7] |
J2218+1146 | 272 | 22431 | 26733 | 456 | 229 | 14 | 318 | 168 | 92% | 8% | 0% | [7] |
J22510740 | 293 | 69641 | 16341 | 9710 | 14025 | 11710 | 6611 | 8819 | 0% | 93% | 7% | [7] |
Metal-poor Comparison Sources | ||||||||||||
J0532+8246 | 24.60.3 | 2038.80.6 | 1663.00.5 | 3063 | 1721 | 552 | 3282 | 662 | 0% | 0% | 100% | [1],[8] |
J0937+2931 | 6.120.07 | 9736 | 12985 | 47.10.5 | 4.30.4 | 42.10.4 | 21.50.5 | 18.90.4 | 98% | 2% | 0% | [9],[10] |
J1055+5443 | 6.90.7 | 15192 | 2232 | 505 | 3922 | 5012 | 16 | 2118 | 93% | 7% | 0% | [11],[7] |
J1416+1348BbbAnalysis based on kinematics of the primary component of this system. | 9.2820.019 | 86.70.3 | 128.00.2 | 6.7990.017 | 425 | 72 | 18.020.12 | 315 | 93% | 7% | 0% | [1],[12] |
J1553+6933 | 343 | 145133 | 132729 | 32128 | 11090 | 29931 | 2973 | 13762 | 0% | 4% | 96% | [7] |
J18101010 | 8.90.6 | 10274 | 2644 | 453 | 464 | 624 | 23 | 433 | 91% | 9% | 0% | [13] |
References. — [1] Gaia EDR3 (Gaia Collaboration et al., 2021); [2] Soubiran et al. (2018); [3] Steinmetz et al. (2020); [4] Halbwachs et al. (2018); [5] Fouqué et al. (2018); [6] Mace et al. (2018); [7] This paper; [8] Reiners & Basri (2006); [9] Faherty et al. (2012); [10] Hsu et al. (2021); [11] Kirkpatrick et al. (2021b); [12] Scholz (2010b); [13] Lodieu et al. (2022)
Using the measured proper motions and radial velocities, and parallax or estimated distances, we computed Galactic , , velocities in the Local Standard of Rest (LSR) assuming a LSR solar velocity of (11.10, 12.24, 7.25) km/s (Schönrich et al., 2010). Uncertainties were propagated by the Monte Carlo method assuming normal distributions. These values are listed in Table 4.5 and plotted in Figure 11 along with the local late-L and T dwarf kinematic sample of Hsu et al. (2021). As expected for our kinematic selection, this sample is more widely dispersed than the local dwarf population, with dispersions of = 63 km/s, = 60 km/s, and = 48 km/s, and a mean azimuthal drift velocity of = 17 km/s. Two candidates, J1130+3139 and J2112+3030, stand out for their large negative velocities, with J1130+3139 showing retrograde motion with respect to the Galactic rest frame. The candidates J12042359, J2112+3030, and J22510740 show large vertical velocities with 75 km/s. Our RV measurement and estimated distance of the T subdwarf J1553+6933 places this source in a particularly unusual position in velocity space, with a total LSR speed of 330 km/s, comparable to J0532+8246 but with distinct component velocities. Overall, we find that the majority of our targets lie outside the 1 velocity sphere for thin disk sources defined by Anguiano et al. (2020).
From the velocities, we computed the kinematic probabilities of membership in the Galactic thin disk (), thick disk (), and halo () following Bensby et al. (2003), assuming . High probability of thin disk membership does not rule out thick disk or halo membership (cf. J18101010), while high probability of thick disk or halo membership likely rules out thin disk membership. Three sources in our sample, the previously identified subdwarfs J0532+8246 and J1553+6933, and the subdwarf candidate J1130+3139, all have high probabilities of halo membership (90%) due to their extreme LSR velocities; while eight of our subdwarf candidates have thick disk probabilities exceeding 90%.
5 Assessment of T Subdwarf Candidates
5.1 New T Subdwarfs
With a thorough assessment of the spectral and kinematic properties of our candidate T subdwarfs, we qualitatively evaluated their likelihood of being a metal-poor brown dwarf by considering four criteria: (1) the degree of deviation of the near-infrared spectrum from the dwarf standards, (2) the presence and strength of 1.25 m K I absorption compared to solar-metallicity field dwarfs for sources with medium-resolution data and spectral types T7, (3) the metallicity inferred from model fitting, and (4) kinematics and the probability of thick disk or halo membership for sources with RV measurements. We treated these criteria in a holistic manner, noting that some criteria vary across spectral type (e.g, metallicity deviations are more pronounced among cloudy late-L and early-T dwarfs) and some may not clarify subdwarf status (e.g., thin disk kinematics does not rule out halo membership). We identified three broad categories for our candidates: likely subdwarfs (sd and esd), “mild” subdwarfs (d/sd), and not subdwarfs. Table 10 summarizes our assessments for both candidates and metal-poor comparison sources.
Assessment Metrics | Spectral Type | ||||||
Name | deviant? | K I? | model? | UVW? | aaMetallicity index from infrared indices; see Section 6.2 and Table 13. | Prior | Adopted |
Confirmed T Subdwarfs | |||||||
J0623+0715 | Y | Y | Y | y | 0.790.09 | sdT3 | |
J15242620 | y? | y | Y | y | 0.740.09 | sdT0 | |
J21120529 | Y | n | Y | y | 0.790.13 | sdT1 | |
Mild T Subdwarfs | |||||||
J0045+7958 | y | N | y? | 0.990.17 | d/sdL9: | ||
J0055+5947 | y? | y | n? | n | 0.780.11 | T8 | d/sdT6.5 |
J0411+4714 | y | y | y | n | 0.950.18 | d/sdT7.5 | |
J08453305 | y | y | 0.850.11 | T7 | d/sdT6.5 | ||
J0911+2146 | y | y? | 0.700.14 | T9 | d/sdT8 | ||
J09530943 | y | y | 0.780.13 | T5.5 | d/sdT6 | ||
J1130+3139 | y | Y | y | Y | 0.860.11 | d/sdT5.5 | |
J1138+7212 | Y | y | 0.760.17 | d/sdT7 | |||
J12042359 | y | y | y | y | 0.900.15 | d/sdT7 | |
J1304+2819 | y | n | n | y? | 0.900.13 | d/sdL9: | |
J13080321 | y | y | n | 1.080.22 | d/sdT8 | ||
J1401+4325 | y? | Y | y | y | 0.870.08 | d/sdT5.5 | |
J15152157 | y? | n | y? | 0.710.11 | d/sdT6 | ||
J1801+4717 | n | Y | y? | n? | 0.930.08 | d/sdT5 | |
J20130326 | y | Y | y | y | 0.790.08 | d/sdT6 | |
J2021+1524 | Y | n | Y | n | 0.940.15 | d/sdL9 | |
J2112+3030 | y? | Y | y | Y | 0.960.13 | d/sdT2.5 | |
J2218+1146 | y | y | y? | n | 0.850.09 | T7p | d/sdT6.5 |
J22510740 | y? | y | n? | y | 0.990.19 | d/sdT7 | |
Not Subdwarfs | |||||||
J0140+0150 | n | N | N | y | 1.110.11 | rT4.5 | |
J0429+3201 | n | y | y | 1.040.11 | T1 | ||
J0433+1009 | n | n? | n? | 1.120.17 | T8 | T8 | |
J0659+1615 | n | n | y? | y? | 0.970.10 | T1 | |
J0758+5711 | n | n | y | n? | 0.930.12 | T6.5 | |
J0843+2904 | n | n? | 0.930.11 | L6: | |||
J11101747 | n | n | y | 1.280.14 | rT8 | ||
J1458+1734 | n | n? | y? | 1.200.17 | T8 | ||
J1710+4537 | n | y | n? | y? | 1.000.11 | T6 | |
Metal-poor Comparison Sources | |||||||
HD 65486B | n | Y | n | 0.950.11 | T4.5 | d/sdT4.5 | |
LHS 6176B | y | y | y | n | 0.760.12 | T8p | d/sdT7.5 |
HD 126053B | Y | Y | n | 0.380.05 | T8p | sdT7.5 | |
GJ 576B | y | Y | y | y | 0.780.05bbWeighted average from multiple spectra. | sdT5.5 | d/sdT5.5 |
Wolf 1130C | Y | Y | y | y | T8p | (e)sdT6: | |
J00041336 | y | Y | 0.910.11 | T2p | d/sdT2: | ||
J00042604 | n | y | 0.910.06 | sdT2 | T3 | ||
J0013+0634 | Y | n | 0.330.18 | T8p | sdT7.5 | ||
J0021+1552 | n | Y | y | 0.810.13 | T4p | d/sdT4 | |
J0057+2013 | Y | y? | sdL7 | d/sdL9 | |||
J03012319 | y | y | 0.890.11 | sdT1 | d/sdT1 | ||
J03095016 | y | y | 0.490.08 | T7p | d/sdT7.5 | ||
J03485620 | y | n | 0.890.08 | T3p | d/sdT4 | ||
J04145854 | Y | Y | 0.320.16 | esdT | esdT6: | ||
J0532+8246 | Y | n | Y | Y | 0.580.07 | esdL7 | esdL8: |
J06166407 | Y | n | Y | 0.540.06 | esdL6 | esdT0: | |
J06456646 | y | N | y? | sdL8 | d/sdT0 | ||
J0833+0052 | y | y? | 0.620.13 | T9p | d/sdT9 | ||
J08500221 | y | y | sdL7 | d/sdL6 | |||
J0937+2931 | Y | Y | n | 0.640.18 | T6p | sdT6 | |
J09392448 | y | y | 0.450.09 | T8p | d/sdT8 | ||
J10193911 | y? | n | 1.000.17 | sdT3 | T4 | ||
J10350711 | Y | y | sdL7 | d/sdL9 | |||
J1055+5443 | N | N | n | sdT8 | Y0 | ||
J11301158 | Y | Y | Y | 0.450.18 | sdT5.5 | sdT5.5 | |
J1158+0435 | Y | y | sdL7 | d/sdL8 | |||
J1307+1510 | n | N | y? | sdL8 | T1 | ||
J1316+0755 | Y | Y | 0.370.17 | sdT6.5 | sdT6.5 | ||
J13380229 | y | y | sdL7 | sdL9 | |||
J1416+1348B | Y | Y | n | 0.230.06 | T7.5p | sdT7 | |
J1553+6933 | Y | Y | Y | Y | 0.650.05 | sdT4 | sdT4 |
J18101010 | Y | Y | Y | n | 0.490.07 | esdT | esdT3: |
J21056235 | y? | y | 0.890.05 | sdT1.5 | d/sdT2: |
Note. — The qualitative assessment metrics are as follows: “Y”: strong positive evidence, “y”: marginal positive evidence, “n”: marginal negative evidence, “N”: strong negative evidence, and “…”: unmeasured or unable to evaluate for this source (e.g., K I for T8 sources and low resolution spectra).
From our candidate sample, we identify J0623+0715, J15242620, and J21120529 as high probability early-type T subdwarfs, as all three exhibit strong deviations from dwarf templates, [M/H] 0.7 from model fits, and thick disk kinematics. J0623+0715 also lacks detectable K I absorption, while J15242620 and J21120529 have weak K I absorption compared to dwarf L/T transition objects. We identify a further 19 candidates as mild subdwarfs, predominantly late-type T dwarfs with -band suppression, modestly subsolar metallicities ([M/H] 0.3) from model fits, and a mixture of population assignments. The remaining 9 candidates appear to be normal late-L and T dwarfs, although a few of these sources warrant further scrutiny as discussed below.
We also assessed the subdwarf status of our comparison sample following the same rubric. While the majority are found to be true subdwarfs, several sources appear to be normal T dwarfs. We reclassify J00042604 and J10193911, classified sdT2 and sdT3 by Greco et al. (2019), as T3 and T4 dwarfs based on their excellent agreement with the dwarf standards. We find J1307+1510, classified sdL8 by Zhang et al. (2018b), to be well-matched to the T1 dwarf standard with strong K I absorption. We also confirm the analysis of Robbins et al. (2023) that J1055+5443 is not an sdT8, as photometrically estimated by Kirkpatrick et al. (2021b), but rather a nearby (6.90.7 pc) Y0 dwarf with unusual colors. Indeed, our spectral model fits suggests this source is highly metal-rich: [M/H] = +0.70 from LOWZ model fits and [M/H] = +0.78 from equivalently matching ElfOwl model fits. However, both fits predict small radii (0.02 R⊙) that are not physical, indicating an overestimated temperature and a likely unreliable metallicity. Our inferred radial velocity of 3922 km/s for this source has insufficient precision to confirm or rule out membership in the 180 Myr Crius 197 association (Moranta et al., 2022) as proposed by Robbins et al. (2023).
5.2 The Remarkable Galactic Orbits of J1130+3139 and J1553+6933
Our analysis identified three sources that stand out in their Galactic motion. The L subdwarf J0532+8246 has previously been recognized as an object on a retrograde Galactic orbit (Burgasser et al., 2008b). The subdwarf candidate J1130+3139 was found to have similar component velocities that place it on a nearly radial Galactic orbit. The orbits for both objects are illustrated in Figure 12, and were computed using the galpy package (Bovy, 2015) assuming the MWPotential2014 potential, solar motion from Schönrich et al. (2010), a solar radius of 8 kpc, and a local circular velocity of 220 km/s. These figures display 100 orbit realizations sampling uncertainties in distance, proper motion, and radial velocity, all of which dominate over possible model uncertainties for J1130+3139 (e.g., choice of Galactic potential). Both sources are currently near apogal, with J0532+8246 reaching an inner radius of 2.6 kpc and J1130+3139 reaching an inner radius of 0.4–1.8 kpc, well within the bulge. Both sources are also closely aligned in velocity space with the Thamnos streams identified by Koppelman et al. (2019), whose members span a broad range of metallicities that encompass those inferred from model fits for these two sources.
The other T subdwarf with unusual kinematics is J1553+6933, previously reported in Meisner et al. (2021). This source’s velocities are distinct from all of the other sources in our sample. Its large inferred Galactic radial () velocity places it on an orbit that reaches R 25 kpc at apogal, with distribution tails extending as far as 200 kpc from the Galactic center. However, this source also has the most poorly measured RV in our sample, +11090 km/s, due to the lack of clear spectral features across the near-infrared region and very low S/N. Nevertheless, with a 300 km/s (similar to J0532+8246), and a velocity that aligns it with the prograde Helmi streams (Helmi et al., 1999), this source warrants further kinematic study.
5.3 Metal-rich Thick-Disk Sources
Three sources in our sample, J0140+0150, J11101747, and J1458+1734 show evidence of kinematic membership in the Galactic thick disk population but have spectra best fit by solar or supersolar metallicity models. J0140+0150, with 99% thick disk membership probability, is found to be the most metal-rich at [M/H] = +0.40, and has the strongest K I lines in the sample (see Table 5 and Figure 5). J11101747 is one of the few sources to show excess flux at -band compared to its closest-match dwarf standard (Figure 3). The nature of these sources suggests that they may be members of the high velocity thick disk, which have high LSR speeds ( 220 km/s) and supersolar metallicities of up to [Fe/H] +0.5 (Yan et al., 2020; Zhu et al., 2021). Bonaca et al. (2017) have demonstrated that these stars likely formed in the inner Milky Way and scattered outward through radial migration. Indeed, the Galactic orbit of J0140+0150 is modestly eccentric (e = 0.45), extending over a radius range of 4–11 kpc (Figure 12). All three sources have relatively large vertical velocities, 50 km/s, pushing them to maximum displacements of 1–2 kpc. While not formally part of the “high” velocity tail, these sources demonstrate that thick disk kinematics are not sufficient evidence for subsolar metallicity.
6 Establishing a T Dwarf Metallicity Classification System
6.1 Building a Temperature and Metallicity Grid
Our comprehensive spectral sample of solar- and subsolar-metallicity T dwarfs provides the opportunity to establish an empirical classification system for cold brown dwarfs that encapsulates both temperature and metallicity effects. We followed prior approaches to qualifying and quantifying metallicity effects in M dwarfs (e.g., Gizis 1997; Lépine et al. 2007; Zhang et al. 2019a) and L dwarfs (e.g., Burgasser et al. 2007a; Zhang et al. 2017, 2018b; Greco et al. 2019). In particular, we aimed to define T subdwarf metallicity classes that roughly align with the metallicities of the existing M and L subdwarf classes.
Gizis (1997) originally proposed two metallicity classes for metal-poor M subdwarfs with [M/H] 1.2 for sdMs and [M/H] 2 for edMs. Subsequent expansion of this system into dM, sdM, esdM, and usdM classes by Lépine et al. (2007), and calibration of metallicity scales using companions of metal-poor FGK primaries and improved spectral models have refined these estimates. A recent analysis of over 1500 M dwarf and subdwarf optical spectra by Hejazi et al. (2020) with classifications on the Lépine et al. (2007) system yields median metallicities of [M/H] = 0.150.27 for dM, [M/H] = 0.750.43 for sdM, [M/H] = 1.420.40 for esdM, and [M/H] = 1.880.18 for usdM (see also Woolf et al. 2009; Lodieu et al. 2019). Zhang et al. (2017) used model fits to optical and infrared spectra to argue for a similar metallicity scale for their defined L dwarf metallicity classes: [M/H] 0.3 for dL, 1.0 [M/H] 0.3 for sdL, 1.7 [M/H] 1.0 for esdL, and [M/H] 1.7 for usdL sources. Given the range of metallicities inferred from our own model fits of the T dwarfs examined here, we restrict our metallicity grid to dT, sdT, and esdT classes, and introduce a mild subdwarf class d/sdT to identify sources with modest subsolar metallicities which are nevertheless distinguishable from solar-metallicity dwarfs (cf. Catchpole et al. 1967; Burgasser et al. 2007a; Bowler et al. 2009).
We adopted a qualitative empirical approach to define temperature and metallicity subclasses, based on principles inherent to the Morgan et al. (1943) stellar classification system. Our system is defined by a grid of standard spectra covering a common wavelength range and resolution (see Morgan & Keenan 1973; Kirkpatrick 2005). We resampled our entire spectral sample onto a common wavelength scale spanning 1.0–2.4 m at a fixed resolution = 300, and normalized all data at the peak flux between 1.2–1.3 m. At this resolution, we are primarily focused on the behavior of broad molecular absorption bands and overall spectral shape, rather than atomic lines. To anchor our system to the existing T dwarf classification system (Burgasser et al., 2006b), we used numerical subtypes that parallel the evolution of molecular features through the dwarf track, notably the 1.6 m CH4 and 1.15 m H2O+CH4 absorption features, and the gradual narrowing of the 1.25 m and 1.6 m spectral peaks. Metallicity types were then segregated based on deviations associated with metallicity effects: unusually blue near-infrared spectral energy distributions driven by reduced cloud opacity in late-L and early-T dwarfs, and enhanced H2 in later T dwarfs; a “muting” of H2O and CH4 bands beyond 1.3 m as H2 dominates near-infrared opacities; and a broadening of the 1.1 m spectral trough, possibly driven by enhanced pressure-broadening of the 0.77 m K I alkali line wings. Mild subdwarfs have spectra similar to their dwarf counterparts but with a slightly enhanced 1.05 m peak and suppressed 2.1 m peak. Extreme subdwarfs have highly distinct spectra from their closest dwarf counterpart, which were assigned by matching the depth of the 1.1 m trough.
We iteratively evaluated these variations by direct comparison between spectra. We then built a two-dimensional grid of representative standards that spanned our available sample. Selection of these standards favored sources with high signal-to-noise spectra and/or accessible declinations when more than one option was available. These standards are listed in Table 11 and displayed in sequence in Figure 13. There are gaps among subclasses in the sd and esd sequences where a clear standard was not available. However, as illustrated in Figure 14, those subtypes with representatives across metallicity classes display continuous variations in spectral features tied to metallicity (e.g., near-infrared slope, breadth of the 1.15 m feature), while maintaining agreement with the dwarf standards in spectral features tied to temperature (e.g., strength of the 1.6 m CH4 band, depth of the 1.15 m H2O+CH4 feature). The highly distinct spectra of the extreme subdwarfs J0532+8246, J06166407, J18101010, and J04145854 make their numerical classifications, anchored to the 1.1 m trough, highly uncertain and are indicated as such with a “:” suffix.
Dwarfs (d)aaDwarf standards from Kirkpatrick et al. (1999, 2010) for L dwarfs and Burgasser et al. (2006b) for T dwarfs; see these references for full source designations. | Mild Subdwarfs (d/sd) | Subdwarfs (sd) | Extreme Subdwarfs (esd) | ||||
SpT | Name | SpT | Name | SpT | Name | SpT | Name |
dL8 | J1632+1904 | d/sdL8 | J1158+0435 | sdL8 | esdL8 | J0532+8246 | |
dL9 | J02554700 | d/sdL9 | J2021+1524 | sdL9 | J13380229 | esdL9 | |
dT0 | J1207+0244 | d/sdT0 | J06456646 | sdT0 | J15242620 | esdT0 | J06166407 |
dT1 | J08370000 | d/sdT1 | J03012319 | sdT1 | J21120529 | esdT1 | |
dT2 | J12540122 | d/sdT2 | J00041336 | sdT2 | esdT2 | ||
dT3 | J12091004 | d/sdT2.5 | J2112+3030 | sdT3 | J0623+0715 | esdT3 | J18101010 |
dT4 | J2254+3123 | d/sdT4 | J0021+1552 | sdT4 | J1553+6933 | esdT4 | |
dT5 | J1503+2525 | d/sdT5.5 | GJ 576B | sdT5.5 | J11301158 | esdT5 | |
dT6 | J1624+0029 | d/sdT6 | J2013-0326 | sdT6 | J0937+2931 | esdT6 | J04145854 |
dT7 | J0727+1710 | d/sdT7.5 | LHS 6176B | sdT7 | J1416+1348B | esdT7 | |
dT8 | J04150935 | d/sdT8 | J09392448 | sdT7.5 | J0013+0634 | esdT8 | |
dT9 | J07220540 | d/sdT9 | J0833+0052 | sdT9 | esdT9 |
We used these standards to determine revised classifications for the remainder of our candidate and comparison sample, listed in Table 10. Note that this analysis resulted in several significant changes in the classifications of previously identified metal-poor late L and T dwarfs. In addition to reclassifying the previously reported subdwarfs J00042604, J10193911, and J1307+1510 as normal early T dwarfs, we reassigned several sources as mild subdwarfs given their relatively weak metallicity signatures. We also reclassified the L subdwarfs J06166407 and J06456646 as early T subdwarfs based on the presence of weak CH4 absorption in their - or -band spectra. For J06166407, this is a substantial shift from its prior esdL6 classification (Zhang et al., 2017), but we assert this reclassification is warranted given the emergent CH4 feature at 1.6 m present in its spectrum, a defining trait for T dwarfs (Burgasser et al., 2006b). For sources with incomplete 1–2.5 m coverage (e.g., Wolf 1130C, J21056235), we indicate uncertainty in their classifications with a “:” suffix.
As with any empirical stellar classification system, the relationship between and spectral type can vary with metallicity, as discussed below. Nevertheless, our model fits indicate that model-fit metallicities of the metallicity class standards align with the metallicity ranges for M and L subdwarf classes, with uncertainty-weighted means of [M/H] = 0.360.09 for d/sdT standards, [M/H] = 0.630.09 for sdT standards, and [M/H] = 1.320.13 for esdT standards. We present this initial grid as a starting point, with the expectation of future expansion and/or reorganization as the known T subdwarf population expands.
6.2 Defining a Metallicity Index
While direct comparison to standards is the optimal method for determining metallicity classifications, metallicity-sensitive spectral indices can be used as estimators of the quantified metallicity of a source. For M subdwarfs, metallicity-sensitive spectral features in the optical (e.g., CaH/TiO; Woolf & Wallerstein 2006; Lépine et al. 2007; Woolf et al. 2009; Dhital et al. 2012; Neves et al. 2014) or infrared (e.g., Na/H2O/CO/FeH; Rojas-Ayala et al. 2012; Mann et al. 2013; Terrien et al. 2012; Newton et al. 2014) are calibrated with sources of known metallicity, typically companions to FGK dwarfs and resolved binaries (Mann et al., 2013; Montes et al., 2018; Qiu et al., 2024). We applied the same approach to our T dwarf sample. We first identified spectral indices in the near-infrared that clearly segregate spectroscopically-identified T subdwarfs and low-metallicity benchmark companions from the predominantly solar-metallicity local T dwarf population. The ten spectral indices reported in Section 4.2 provide 45 unique index pairings, of which six demonstrate consistent separation between dwarfs and subdwarfs. These pairs as listed in Table 12 and displayed in Figures 15 and 16. We combined two index pairings, [H2O-J] versus [Y/J] and [CH4-J] versus [Y/J] into a composite index. For the three pairings with ordinate indices [H2O-H] and [Y/J], the subdwarfs were found to lie above the dwarf sequence, consistent with weakened H2O absorption at 1.5 m and a brighter -band peak, both noted previously as metallicity indicators. The other three pairings with the ordinate index [K/H] have subdwarfs lying below the dwarf sequence, consistent with suppressed -band peaks due to enhanced CIA H2 absorption. Burgasser et al. (2006b) had previously proposed use of the [K/J] color index as an indicator of enhanced CIA H2 absorption, but probable errors in the order-to-order flux calibration of the multi-order spectral data in our sample (cf. Section 3.6) results in higher scatter for this color index, hence we excluded it from our metallicity index set.
Abscissa (x) | Ordinate (y) | Solar Metallicity Sequence Polynomial Fit | |||||
Index | Range | Index | |||||
[CH4-H] | 0.55–1.00 | [H2O-H] | 0.52753311 | 0.28280498 | 0.37088006 | 0.053 | |
[H2O-J] | 0.05–0.60 | [H2O-H] | 0.44401774 | 0.38492182 | 0.20196172 | 0.033 | |
[H2O-J] | 0.05–0.65 | [K/H] | 1.23813044 | 0.23225536 | 0.31791332 | 0.097 | |
[H-dip] | 0.20–0.90 | [K/H] | 0.01818876 | 0.17387556 | 0.22533516 | 0.046 | |
aaNot used as a metallicity index. | [CH4-H] | 0.10–0.65 | [K/H] | 0.78127813 | 0.92425434 | 0.1014752 | 0.043 |
aaNot used as a metallicity index. | [H2O-J],[CH4-J] | 0.15–0.75 | [Y/J] | 0.19243998 | 0.16534304 | 0.3675336 | 0.048 |
Note. — indicates an average of two indices, or the value of one index if the other is unmeasured. Fit relations for the dwarf standards are defined as for indices (x,y) over the ranges specified. The quantity measures the root mean square deviation of y-values about this relation. See Figures 15 and 16 for visualization of these fits.
For each of these pairings, we fit second-order polynomials to the index values of the local T dwarf sample over an abscissa range that had the clearest offset from the T subdwarfs. In some cases, the abscissa range encompasses a subset of our sample; e.g., only early-type or only late-type T dwarfs. Fit parameters and scatter about the best fits are listed in Table 12 and displayed in Figures 15 and 16.
To guide our interpretation of these trends, we also show the same spectral indices measured on the LOWZ atmosphere models over the temperature range 600 K 1600 K, surface gravities 4.0 5.25 (cgs), metallicities 2.5 [M/H] +0.5, C/O = 0.55, and = 2. We find that five of these index pairs show good separation of metallicity and temperature and reasonable overlap with our measured values. The [Y/J] versus H2O-J,CH4-J pair shows the most discrepant values between models and measurements, and poor metallicity separation in the models for [M/H] 0.5, so we reject this pair as a viable metallicity diagnostic. Of the remaining indices, [H2O-H] versus [CH4-H], [H2O-H] versus [H2O-J], and [K/H] versus [H2O-J] show reasonably orthogonal metallicity and temperature trends in the models, although there is disagreement between models and measurements for the last index for the late-L and early-T dwarfs, likely due to unincorporated cloud opacity effects. The [K/H] versus [H-dip] and [K/H] versus [CH4-H] pairings show greater degeneracy between metallicity and temperature, while [K/H] versus [H-dip] shows greater dynamic range; we therefore prioritize this index pair over [K/H] versus [CH4-H].
Based on these analyses, we defined four metallicity indices, , as the ratio between the metallicity-sensitive index [H2O-H] or [K/H] and their expected values based on the solar-metallicity dwarf sequences:
Here, the notation is the polynomial fit for the dwarf sequence for the index pairings and listed in Table 12. These relationships are defined such that subsolar metallicity objects are expected to have .
Prior | Model | |||||||
Name | SpT | [Fe/H] | [M/H]aaBased on best-fit spectral model; see Table 7. | |||||
Benchmark Companions | ||||||||
HD 3651B | T8 | 1.090.20 | 1.090.20 | +0.14 | 0.00 | |||
HD 65486B | d/sdT4.5 | 1.010.11 | 0.800.29 | 0.890.16 | 0.950.11 | 0.28 | 0.53 | |
LHS 6176B | d/sdT7.5 | 0.710.34 | 0.770.13 | 0.760.12 | 0.30 | 0.21 | ||
Ross 458C | rT8.5 | 1.710.30 | 1.710.30 | +0.25 | +0.95 | |||
HD 126053B | sdT7.5 | 0.380.08 | 0.380.08 | 0.38 | 0.35 | |||
GJ 570D | T7.5 | 0.850.16 | 0.850.16 | +0.05 | 0.11 | |||
GJ 576B (GNIRS) | d/sdT5.5 | 0.920.12 | 0.710.26 | 0.630.10 | 0.750.16 | 0.37 | 0.47 | |
(XS) | d/sdT5.5 | 0.940.14 | 0.610.41 | 0.560.10 | 0.710.21 | 0.37 | 0.53 | |
(NIRES) | d/sdT5.5 | 0.960.11 | 0.810.43 | 0.740.09 | 0.830.13 | 0.37 | 0.21 | |
G 204-39B | T6.5 | 1.010.15 | 1.320.21 | 1.110.19 | 0.04 | +0.43 | ||
HN PegB | T2.5 | 1.170.13 | 0.950.07 | 1.130.33 | 1.010.11 | 0.08 | 0.11 | |
Subdwarf Candidates | ||||||||
J0055+5947 | d/sdT6.5 | 0.990.21 | 0.750.34 | 0.750.12 | 0.810.15 | +0.04 | ||
J0140+0150 (SpeX) | rT4.5 | 1.140.22 | 1.030.17 | 1.250.40 | 1.160.16 | 1.120.11 | +0.21 | |
(NIRES) | rT4.5 | 1.160.16 | 0.990.10 | 1.170.16 | 1.070.11 | +0.40 | ||
J0411+4714 | d/sdT7.5 | 1.120.17 | 0.770.32 | 0.830.14 | 0.950.18 | 0.12 | ||
J0429+3201 | T1 | 1.070.10 | 1.010.06 | 0.960.17 | 1.020.06 | |||
J0433+1009 | T8 | 1.080.16 | 1.210.40 | 1.100.16 | +0.12 | |||
J0623+0715 | sdT3 | 0.680.09 | 0.840.06 | 0.670.14 | 0.860.12 | 0.790.09 | 0.68 | |
J0659+1615 | T1 | 1.000.09 | 0.820.14 | 0.950.11 | 0.03 | |||
J0758+5711 | T6.5 | 1.090.16 | 0.870.14 | 0.970.12 | 0.17 | |||
J08453305 | d/sdT6.5 | 0.920.11 | 0.790.28 | 0.790.18 | 0.850.11 | 0.35 | ||
J0911+2146 | d/sdT8 | 0.700.17 | 0.700.17 | 0.00 | ||||
J11101747 | rT8 | 1.200.24 | 1.200.24 | +0.20 | ||||
J1130+3139 | d/sdT5.5 | 1.030.15 | 0.790.11 | 0.880.12 | 0.17 | |||
J1138+7212 | d/sdT7 | 1.000.24 | 0.670.34 | 0.700.12 | 0.760.17 | 0.36 | ||
J12042359 | d/sdT7 | 1.130.20 | 0.810.39 | 0.820.15 | 0.910.18 | 0.15 | ||
J13080321 | d/sdT8 | 0.910.18 | 0.910.18 | 0.11 | ||||
J1401+4325 | d/sdT5.5 | 0.880.13 | 0.860.14 | 0.870.10 | 0.13 | |||
J1458+1734 | T8 | 1.120.18 | 1.120.18 | +0.13 | ||||
J15152157 | d/sdT6 | 0.730.31 | 0.710.11 | 0.710.11 | 0.08 | |||
J15242620 | sdT0 | 0.800.08 | 0.800.08 | 0.69 | ||||
J1710+4537 | T6 | 1.010.13 | 0.980.40 | 0.950.15 | 0.980.10 | +0.02 | ||
J1801+4717 | d/sdT5 | 0.870.10 | 0.940.12 | 0.910.09 | 0.09 | |||
J20130326 | d/sdT6 | 0.880.11 | 0.760.12 | 0.830.10 | 0.21 | |||
J2021+1524 | d/sdL9 | 0.850.13 | 0.850.13 | 0.35 | ||||
J21120529 | sdT1 | 0.890.08 | 0.650.14 | 0.790.13 | 0.74 | |||
J2112+3030 | d/sdT2.5 | 0.900.11 | 1.050.07 | 0.720.16 | 0.950.14 | 0.43 | ||
J2218+1146 | d/sdT6.5 | 0.830.13 | 0.860.15 | 0.840.09 | 0.10 | |||
J22510740 | d/sdT7 | 0.840.39 | 0.880.17 | 0.880.16 | 0.00 | |||
Metal-Poor Comparison Sources | ||||||||
J00041336 | d/sdT2: | 0.980.11 | 0.960.08 | 0.750.13 | 0.910.11 | 0.56 | ||
J00042604 | T3 | 0.880.10 | 0.930.07 | 0.850.18 | 0.910.06 | 0.40 | ||
J0013+0634 | sdT7.5 | 1.370.36 | 0.310.06 | 0.320.15 | 0.61 | |||
J0021+1552 | d/sdT4 | 0.830.15 | 0.920.13 | 0.650.22 | 0.720.14 | 0.810.13 | 0.44 | |
J03012319 | d/sdT1 | 0.890.09 | 0.890.09 | 0.37 | ||||
J03095016 | d/sdT7.5 | 0.960.16 | 0.500.09 | 0.630.22bbIndividual values differ by more than the uncertainties; reported value and uncertainty is the unweighted mean and standard deviation of the individual values. | 0.36 | |||
J03485620 | d/sdT4 | 0.830.13 | 0.900.09 | 0.820.35 | 0.920.13 | 0.890.08 | 0.00 | |
J04145854 | esdT6: | 0.480.10 | 0.380.08 | 0.120.14 | 0.100.10 | 0.270.16bbIndividual values differ by more than the uncertainties; reported value and uncertainty is the unweighted mean and standard deviation of the individual values. | 1.34 | |
J0532+8246 | esdL8: | 0.580.07 | 0.580.07 | 1.6 | 1.53 | |||
J06166407 | esdT0: | 0.540.06 | 0.540.06 | 1.6 | 1.13 | |||
J0833+0052 | d/sdT9 | 0.620.12 | 0.620.12 | 0.08 | ||||
J0937+2931 | sdT6 | 0.880.09 | 0.550.33 | 0.520.08 | 0.640.18 | 0.59 | ||
J09392448 | d/sdT8 | 0.450.10 | 0.450.10 | 0.24 | 0.35 | |||
J09530943 | d/sdT6 | 0.770.24 | 0.790.16 | 0.780.13 | 0.43 | |||
J10193911 | T4 | 1.150.15 | 0.790.24 | 0.980.13 | 1.000.17 | 0.00 | ||
J11301158 | sdT5.5 | 0.790.15 | 0.390.15 | 0.370.10 | 0.520.19bbIndividual values differ by more than the uncertainties; reported value and uncertainty is the unweighted mean and standard deviation of the individual values. | 0.56 | ||
J1316+0755 | sdT6.5 | 1.010.23 | 0.340.18 | 0.330.06 | 0.560.32bbIndividual values differ by more than the uncertainties; reported value and uncertainty is the unweighted mean and standard deviation of the individual values. | 0.65 | ||
J1416+1348B | sdT7 | 0.230.05 | 0.230.05 | 0.35 | 0.67 | |||
J1553+6933 | sdT4 | 0.600.08 | 0.660.07 | 0.620.17 | 0.680.09 | 0.650.05 | 0.96 | |
J18101010 | esdT3: | 0.430.06 | 0.500.05 | 0.690.18 | 0.490.07 | 1.5 | 1.27 | |
J21056235 | d/sdT2: | 0.860.09 | 0.910.05 | 0.890.05 | 0.32 |
Table 13 lists the individual values for our spectral sample, as well as the combined index , while Figure 17 displays the distributions of these indices. Uncertainties for the individual values were estimated from the original spectral index uncertainties and the scatter in the dwarf sequence fits using Monte Carlo methods, and are typically 0.1–0.2. Note that shows larger uncertainties due to the greater scatter of its dwarf sequence fit. We rejected all values of with uncertainties greater than 0.4, then computed as the uncertainty-weighted mean (Eqn. 10). For some sources the scatter among values exceeded their individual uncertainties (i.e., ), in which case we adopted a straight mean with uncertainty equal to the standard deviation of the values. The combined index has a typical uncertainty of 0.1–0.2. Metallicity index values for sources with multiple spectra (GJ 576B and J0140+0150) agree within these uncertainties. Several sources in our sample, including the benchmark companions Wolf 940B and Wolf 1130C, fall outside the fit ranges in Table 12 or had incomplete spectral coverage and thus values were unable to be measured.
We found excellent alignment between and our metallicity classifications. Sources with dwarf classifications, including our local T dwarf sample, have = 1 within 10-15%, while subdwarfs and extreme subdwarfs typically having and , respectively. All three of our new subdwarf discoveries, J0623+0715, J15242620, and J21120529 have 0.8. On the other hand, there are outliers among our late-type T dwarfs, notably the d/sdT9 J0833+0052 with , the d/sdT8 J09392448 with , and the sdT7 standard J1416+1348B with , all smaller than expected for their metallicity classifications. In addition, the peculiar red T8.5 Ross 458C, which has an unusually bright -band peak, has a large . These sources have only the index measured, which is based on the [K/H] versus [H-dip] index pairing. As noted above, the LOWZ models show a coupling between temperature and metallicity for these index pairs at low temperatures (Figure 16) and, in the case of Ross 458C, may not account for enhanced cloud opacity (Burgasser et al., 2010c; Morley et al., 2012). Thus, the discrepant values of these sources, and potentially other very late-type T dwarfs, may reflect a temperature bias in this particular metallicity index.
To further assess the reliability of the index, Figure 18 compares this index to total metallicity [M/H] based on the best-fit spectral models (Table 7), and to the iron abundances [Fe/H] of the primaries of our benchmark companions (Tables 2 and 3.6). The majority of our sample shows a roughly linear trend between and [M/H], with a Pearson r correlation of 0.460.07. The dispersion becomes larger at the lowest metallicities and values, and may reflect the temperature-metallicity coupling for the index for very late-type T dwarfs. Excluding the L dwarfs in our sample and the unusually red T8.5p dwarf Ross 458C, we find a roughly linear trend between and [M/H] for 57 sources:
spanning 0.24 1.2 and 1.34 [M/H] +0.43, with a metallicity scatter of 0.29 dex. We emphasize that this is a model-dependent correlation, and may be skewed by inaccuracies in the model fitting. However, this trend also accurately reproduces the correlation between and [Fe/H] measurements for the primaries of our benchmark companions, with two key exceptions: the aforementioned Ross 458C and the sdT7.5 HD 1260538B, which has a low value of for its primary [Fe/H] = 0.38. Excluding these sources, we find a tight correlation among the 9 remaining benchmarks (Pearson r = 0.520.23) with a similar linear trend
for 0.71 1.11 and 0.37 [M/H] +0.14, with a metallicity scatter of 0.12 dex. The small number of benchmark companions with both [Fe/H] and measurements, and the limited metallicity range of this sample, warrants some caution in the use of this relationship. Nevertheless, these comparisons indicate an overall self-consistency between the metallicity classifications, the index, and the actual metallicities of our sources, whether inferred from spectral model fits or stellar primaries.
7 Discussion
7.1 Effective Temperature Trends for T Subdwarfs
The classification of a star is a proxy for its physical parameters, including temperature, luminosity, surface gravity, and (in the case of metallicity classes) composition. For cool stars and brown dwarfs, there are numerous relations linking dwarf numerical spectral types to luminosities and effective temperatures, the latter either requiring a measured or estimated radius inferred from atmosphere model fitting (Marocco et al., 2013; Pecaut & Mamajek, 2013; Filippazzo et al., 2015; Dupuy & Liu, 2017; Kirkpatrick et al., 2021b; Sanghi et al., 2023). There are well-documented metallicity offsets in the spectral type/temperature relations of M and L dwarfs, with subdwarfs typically found to be warmer by several 100 K than their equivalently-classified dwarf counterparts (Burgasser & Kirkpatrick, 2006; Burgasser et al., 2007a; Rajpurohit et al., 2014; Zhang et al., 2017). It is thus useful to explore how effective temperature trends vary with metallicity in the T dwarf regime.
Our first approach was to compare effective temperatures from our model fits to assigned spectral classifications, as illustrated in Figure 19. We see the expected decline of with later spectral type, although many of our sources—including our local comparison sample—are warmer than previously-established dwarf empirical relations. These offsets are particularly large for the late-L and early-T dwarfs. For example, our model fit temperature for J06166407, = 2181 K, is over 800 K warmer than the equivalent solar-metallicity T0 dwarf ( 1250 K; Filippazzo et al. 2015). As this source has a measured parallax, we can also deduce that its model-fit temperature is overestimated based on the non-physically small radius required to match its absolute fluxes. Similar scaling issues are seen for other sources in our sample with trigonometric parallaxes (Table 7).
We therefore considered an alternative estimate for based on bolometric magnitude
where = 5772 K, = 6.9571010 cm, and = 4.74 are the temperature, radius and bolometric magnitude of the Sun (Mamajek et al., 2015). Ignoring interstellar absorption, the bolometric magnitude can be inferred from the absolute magnitude in filter via the bolometric correction (BC)
While empirical BCs for solar-metallicity T dwarfs exist in the literature (e.g., Liu et al. 2010; Filippazzo et al. 2015; Sanghi et al. 2023), we chose to compute BCs from the LOWZ atmosphere models to incorporate metallicity effects. For models computed in surface fluxes,
(Casagrande & VandenBerg, 2014), where is the apparent magnitude computed directly from the model spectra. As our available photometry is on the Vega system, we computed spectrophotometric model magnitudes as
using a model spectrum of Vega from Kurucz (1993) and filter profiles from the Spanish Virtual Observatory Filter Profile Service777http://svo2.cab.inta-csic.es/theory/fps/. (Rodrigo et al., 2012; Rodrigo & Solano, 2020). Appendix B provides a list of bolometric corrections for various infrared filters for the LOWZ models; here, we focus on and photometry for sources with measured parallaxes. As the bolometric corrections are inherently dependent on the physical properties of the atmosphere, we assumed = 5.0 (cgs), = 2 (cgs), C/O = 0.55, and a metallicity based on the classification: [M/H] = 0 for dT, [M/H] = 0.25 for d/sdT, [M/H] = 0.75 for sdT, and [M/H] = 1.5 for esdT. We also assumed a theoretical radius of = 0.077 R⊙ based on the solar-metallicity evolutionary models of Burrows et al. (2001) and Baraffe et al. (2003), and assumed a 10% uncertainty on this value. We iteratively converged to a self-consistent bolometric temperature starting from the model fit temperature, and estimated the uncertainty in as the greater of the combined formal uncertainty from both filters,
(here is parallax, x {,}) and half of the difference between the temperatures inferred from and .
MKO | WISE | Adopted | Model | ||||||||
Name | SpT | aaBolometric corrections computed from the LOWZ models assuming = 5.0 (cgs), = 2 (cgs), C/O = 0.55, a metallicity based on the source classification ([M/H] = 0 for dwarfs, [M/H] = 0.25 for mild subdwarfs, [M/H] = 0.75 for subdwarfs, and [M/H] = 1.5 for extreme subdwarfs), and the closest temperature to in the model grid. See Appendix B for the full list of bolometric corrections. | aaBolometric corrections computed from the LOWZ models assuming = 5.0 (cgs), = 2 (cgs), C/O = 0.55, a metallicity based on the source classification ([M/H] = 0 for dwarfs, [M/H] = 0.25 for mild subdwarfs, [M/H] = 0.75 for subdwarfs, and [M/H] = 1.5 for extreme subdwarfs), and the closest temperature to in the model grid. See Appendix B for the full list of bolometric corrections. | ||||||||
(mag) | (mag) | (mag) | (K) | (mag) | (mag) | (mag) | (K) | (K) | (K) | ||
T Dwarf Benchmarks | |||||||||||
HD 3651B | T8 | 16.080.03 | 2.73 | 18.81 | 80180 | bbThe proximity of a bright stellar companion prevents WISE detection of this source. | 80180 | 824 | |||
HD 65486B | d/sdT4.5 | 14.790.08 | 2.29 | 17.08 | 1193119 | 12.530.02 | 5.15 | 17.68 | 1039104 | 111677 | 993 |
LHS 6176B | d/sdT7.5 | 16.560.03 | 2.61 | 19.17 | 73874 | 13.030.02 | 5.44 | 18.47 | 86787 | 80264 | 846 |
Ross 458C | rT8.5 | 15.020.05 | 2.73 | 17.75 | 1023102 | 13.590.02 | 5.52 | 19.10 | 75075 | 887137 | 812 |
HD 126053B | sdT7.5 | 17.500.05 | 2.20 | 19.69 | 65465 | 13.570.03 | 5.93 | 19.49 | 68569 | 67034 | 746 |
GJ 570D | T7.5 | 16.470.05 | 2.67 | 19.15 | 74274 | 13.290.01 | 5.34 | 18.63 | 83684 | 78947 | 837 |
GJ 576B | d/sdT5.5 | 15.190.02 | 2.71 | 17.90 | 98899 | 12.910.02 | 5.32 | 18.23 | 91792 | 95348 | 995 |
G 20439B | T6.5 | 15.420.09 | 2.66 | 18.09 | 94795 | 13.110.01 | 5.18 | 18.28 | 90591 | 92646 | 984 |
Wolf 1130C | (e)sdT6: | 18.540.09 | 1.25 | 19.79 | 64064 | 13.770.05 | 5.97 | 19.74 | 64865 | 64432 | 700ccFit at limit of model parameters. |
Wolf 940B | T8.5 | 17.560.06 | 2.36 | 19.92 | 62162 | 14.400.06 | 6.05 | 20.45 | 55055 | 58635 | 636 |
HN PegB | T2.5 | 14.570.03 | 2.19 | 16.75 | 1288129 | bbThe proximity of a bright stellar companion prevents WISE detection of this source. | 1288129 | 1284 | |||
T and Y Subdwarfs | |||||||||||
J03012319 | d/sdT1 | 14.490.26 | 2.24 | 16.74 | 1292130 | 11.950.22 | 4.49 | 16.44 | 1383139 | 133896 | 1488 |
J03095016 | d/sdT7.5 | 16.290.13 | 2.66 | 18.96 | 77578 | 12.750.13 | 5.27 | 18.02 | 96196 | 86893 | 795 |
J0532+8246 | esdL8: | 13.230.06 | 2.26 | 15.48 | 1725173 | 11.730.03 | 4.73 | 16.47 | 1376138 | 1550175 | 1730 |
J06166407 | esdT0: | 12.80.7 | 2.3 | 15.10 | 1884197 | 11.70.7 | 4.58 | 16.2 | 1452152 | 1668288 | 2181ddModel fit produces non-physically small radius (R 0.05 R⊙), suggesting an overestimated . |
J06456646 | d/sdT0 | 14.180.12 | 2.04 | 16.21 | 1458146 | 11.960.12 | 4.49 | 16.45 | 1380138 | 141981 | 1594 |
J0833+0052 | d/sdT9 | 19.860.15 | 1.62 | 21.49 | 43343 | 14.490.12 | 6.13 | 20.63 | 52853 | 48048 | 528ddModel fit produces non-physically small radius (R 0.05 R⊙), suggesting an overestimated . |
J08500221 | d/sdL6.5 | 12.700.11 | 1.96 | 14.66 | 2086209 | 10.360.10 | 4.23 | 14.58 | 2121212 | 2104115 | 1852 |
J0937+2931 | sdT6 | 15.710.04 | 2.57 | 18.28 | 90691 | 12.740.03 | 5.33 | 18.07 | 95295 | 92947 | 859 |
J09392448 | d/sdT8 | 17.130.12 | 2.37 | 19.50 | 68368 | 13.100.12 | 5.44 | 18.54 | 85385 | 76885 | 682 |
J1158+0435 | d/sdL8 | 13.470.16 | 1.96 | 15.43 | 1745175 | 11.220.16 | 4.23 | 15.45 | 1739174 | 1742107 | 1870 |
J1416+1348B | sdT7 | 17.420.02 | 2.20 | 19.62 | 66667 | 12.940.04 | 5.59 | 18.54 | 85485 | 76094 | 635 |
J18101010 | esdT3: | 17.520.16 | 1.42 | 18.95 | 77778 | 12.840.16 | 5.50 | 18.34 | 89390 | 83558 | 869 |
J15341043 | (e)sdY: | 23.40.3 | 0.56ddModel fit produces non-physically small radius (R 0.05 R⊙), suggesting an overestimated . | 22.9 | 31432 | 15.090.19 | 6.26eeUsing the BC value for = 500 K and [M/H] = 1.5. | 21.35 | 44745 | 38167 |
Table 14 lists the resulting bolometric temperatures, which are displayed in Figure 19. These temperatures are generally lower than those inferred from the model fits, in some cases by hundreds of degrees Kelvin (e.g., J06166407), in alignment with the underestimated radii from the model fits for these sources. Nevertheless, the bolometric temperatures for the late-L and early-T subdwarfs remain warmer than the dwarf sequence. The mid- and late-type T subdwarfs, on the other hand, appear to be cooler than their equivalently-classified dwarf counterparts, notably so for J18101010 ( = 790 K compared to 1150 K for a typical T3 dwarf) and Wolf 1130C ( = 645 K compared to 940 K for a typical T6 dwarf). This apparent reversal in the metallicity-induced temperature offset may be caused by a flattening of the H2O and CH4 features by enhanced H2 absorption, or reduced absorption from these molecules due to reduced C and O abundances. Both cases lead to earlier spectral classifications at a given temperature. We also note that this temperature reversal occurs around the L/T transition, where dwarf spectra evolve due to changes in photospheric cloud properties (or equivalently the adiabatic gradient; Tremblin et al. 2016), producing a flat temperature/spectral type relation; i.e., large changes in spectral morphology with minimal change in (Burgasser et al., 2002a; Dupuy & Liu, 2012). The absence of significant condensate opacity in metal-poor brown dwarf atmospheres could also modify the relationship between temperature and the evolution of spectral morphology. This discussion highlights the non-trivial systematic effects that may be present when inferring subdwarf temperatures from either spectral types or model fitting, and emphasizes the need for additional parallax measurements to infer robust atmospheric properties.
7.2 Prospects for Growing the Sample of T Subdwarfs
To deepen our understanding of the formation, evolution, and atmosphere properties of metal-poor brown dwarfs, it is necessary to expand the known sample of T and Y subdwarfs. Increasing the local sample (e.g., d 20 pc) is a challenge given the sparsity of metal-poor stars in general and metal-poor brown dwarfs in particular, the latter estimated to be as low as 0.1% (Schneider et al., 2020). Nevertheless, this study’s focus on relatively bright sources () accessible from the northern hemisphere means that several promising T subdwarf candidates remain to be characterized in the BYW sample alone (cf. Brooks et al. 2022; Meisner et al. 2023).
Current and forthcoming deep imaging and spectroscopic surveys are also expected to substantially increase the number of halo brown dwarfs with photometric, spectroscopic, and astrometric characterization. In addition to the individual discoveries now being made in deep pointings with JWST to limiting magnitudes of 30 AB (Nonino et al., 2023; Glazebrook et al., 2023; Wang et al., 2023; Burgasser et al., 2024; Hainline et al., 2024; Holwerda et al., 2024), the recently-launched Euclid mission is expected to detect of order 104 thick disk and 103 halo T dwarfs in its 15,000 deg2 wide-field survey, reaching 24.5 AB (Laureijs et al., 2011; Zhang et al., 2019b; Solano et al., 2021). A significant fraction of these sources will also be observed with Euclid’s slitless spectroscopic mode, yielding 1.25–1.85 m spectra at 250 for potentially hundreds of metal-poor brown dwarfs. Starting in the late 2020s, the Nancy Grace Roman Space Telescope’s High Latitude Wide Area Imaging Survey, spanning 2,000 deg2 to 27 AB (Troxel et al., 2021), is expected to detect 5–10 more sources than Euclid (Zhang et al., 2019b). This survey is also matched to a spectroscopic survey that will acquire 1.0–1.93 m spectra at 400–800, covering many of the metallicity-sensitive features examined in this study. These infrared surveys will be complemented by optical photometry from the Vera Rubin Observatory Large Survey of Space and Time (LSST; Ivezić et al. 2019). While this survey will detect few halo brown dwarfs due to its poor sensitivity at low temperatures (10s to 100s over the Southern sky; Zhang et al. 2019b), it will provide useful color limits and -band dectections for refining candidate samples.
The multi-epoch astrometry of all three surveys will enable identification of candidate nearby halo brown dwarfs through reduced proper motion constraints, candidates which can be spectroscopically confirmed with JWST or planned 30-meter class ground-based facilities. In short, the limited sample presented here is a small fraction of the halo brown dwarf population expected to be discovered and explored over the next decade, primarily by space-based infrared surveys that continue the legacy of WISE.
8 Summary
This work has presented a comprehensive study of the spectroscopic properties of currently known metal-poor T dwarfs, augmented with new discoveries from the Backyard Worlds: Planet 9 program. Our key results are as follows:
Using photometric and astrometric data from several wide-field infrared surveys, most notably the CatWISE2020 catalog, and imposing color and reduced proper motion constraints, we identified a sample of 95 high-probability, bright ( 18.5) T subdwarf candidates visible from the northern hemisphere, which includes several previously reported metal-poor brown dwarfs (Schneider et al., 2020; Meisner et al., 2020; Kirkpatrick et al., 2021b, 2024).
Near-infrared spectroscopic follow-up with Keck/NIRES, IRTF/SpeX, APO/TSpec, Magellan/FIRE, and Gemini-N/GNIRS of a subset of these candidates has allowed us to identify three new T subdwarfs: the sdT0 J15242620, the sdT1 J21120529, and the sdT3 J0623+0715. We also identify 19 new mildly metal-poor T dwarfs. These identifications are based on metallicity-sensitive spectral features, including a broadened 1.1 m absorption band, suppressed 2.1 m peak emission, and weak or absent K I lines at 1.17 and 1.25 m.
Atmosphere model fits to the spectra of candidate and previously reported late-L and T subdwarfs and benchmark companions allowed us to validate the metal-poor nature of our candidates, as well as identify several prior discoveries as normal dwarfs. Comparing to several sets of spectral models, we find that the LOWZ models (Meisner et al., 2021) typically provide the best fits for all brown dwarfs with 800 K 1600 K, the SAND models (Alvarado et al., 2024) perform better at higher temperatures and low metallicities, and the ElfOwl models (Mukherjee et al., 2024) perform better at lower temperatures and near-solar metallicities. Notably, the SAND models provide self-consistent fits to the spectra of extreme L subdwarfs (e.g., J0532+8246) and T subdwarfs (e.g., J18101010), indicating that these models are well-suited for metal-poor brown dwarfs.
By forward-modeling moderate-resolution Keck/NIRES spectral data, we measured the radial velocities of 27 candidate and comparison late-L and T subdwarfs. The corresponding velocities are more dispersed than the local T dwarfs, and several sources have velocities consistent with the thick disk or halo populations. Two subdwarfs, the esdL8 J0532+8246 and the d/sdT5.5 J1130+3139, have retrograde Galactic orbits potentially consistent with the Thamnos stream; while the sdT4 J1553+6933 has a motion potentially consistent with the Helmi stream. We also identify three sources with thick disk kinematics that appear to be metal-rich, and may be part of the high-velocity thick disk that originated in the inner Milky Way.
We used our spectral sample to construct a metallicity classification system for T dwarfs anchored to spectral standards. Our scheme encompasses classes of mild subdwarfs (d/sdT), subdwarfs (sdT), and extreme subdwarfs (esdT), with estimated metallicities of [M/H] 0.4, [M/H] 0.6, and [M/H] 1.3, respectively, matching the metallicity scales of M and L subdwarfs. We have defined a near-infrared index to quantify metallicity, , and demonstrated that this index correlates with the metallicities inferred from spectral model fitting and the iron abundances measured from the primaries of a small sample of benchmark T dwarf companions.
We examined the relationship between effective temperature and spectral type for brown dwarfs of differing metallicities, and find that late-L and early-T subdwarfs are generally warmer and mid- and late-T subdwarfs generally cooler than equivalently-classified solar-metallicity dwarfs. The latter offset may be due to the enhanced H2 absorption and reduced C and O abundances weakening the molecular bands used for spectral classification, while the switch from warmer to colder may be driven by a reduction in condensate opacity in subdwarfs.
The sample explored here encompasses nearly all of the lowest-temperature, metal-poor brown dwarfs currently known, but it is in no way complete. In addition to T subdwarf candidates identified at southern declinations and fainter magnitudes by the BYW program, additional yields from JWST, Euclid, LSST and the Nancy Grace Roman Space Telescope are expected to dramatically expand our sample of T subdwarfs from tens to thousands. Adding these discoveries to the known T subdwarf sample will provide a more robust assessment of the diversity of metal-poor brown dwarf atmospheres, including variations in individual elemental abundances that are tied to distinct Milky Way populations. Both near-infrared spectroscopy and parallax measurements for these sources will be crucial, the latter providing an essential constraint on bolometric luminosity to calibrate spectral model fits. Indeed, our study demonstrates that advances in metal-poor atmosphere modeling, as illustrated by the SAND models, yield more reliable fits and self-consistent radii for absolute flux densities. We anticipate that our assessment of the currently-known T subdwarf population and this first approach at T dwarf metallicity classification will require revision as future searches, observations, and modeling are undertaken.
Appendix A IRTF/SpeX Comparison Sample
The following Tables provide the measured spectral indices (Table A) and LOWZ model fit parameters (Table A) for a sample of 184 late-L and T dwarfs with low-resolution near-infrared spectra obtained with the IRTF/SpeX instrument (Rayner et al., 2003) and curated in the SPLAT archive (Burgasser & Splat Development Team, 2017). Data references for the individual spectra are included in the Table endnotes. The model fits for these data were conducted exclusively on the individual grid models without interpolation.
Designation | SpTaaSpectral type as inferred from direct comparison to spectral standards. | [CH4-J] | [CH4-H] | [CH4-K] | [H2O-J] | [H2O-H] | [H2O-K] | [Y/J] | [K/H] | [K/J] | [H-dip] | Ref. | |
J00001354+2554180 | T4.5 | 0.5490.015 | 0.5480.006 | 0.2590.005 | 0.3270.002 | 0.3490.005 | 0.4800.004 | 0.4760.005 | 0.3520.002 | 0.1970.001 | 0.7380.008 | 1.000.09 | [3] |
J00062785+1857288 | L7 | 1.0110.021 | 1.0560.018 | 1.0010.013 | 0.7270.018 | 0.7810.017 | 0.8090.013 | 0.5080.013 | 0.7110.011 | 0.6910.011 | 0.9950.012 | 0.930.12 | [29] |
J00150206+2959323 | L7 | 0.9820.013 | 1.0470.016 | 0.9240.009 | 0.7730.011 | 0.7430.012 | 0.7700.011 | 0.5850.011 | 0.6580.008 | 0.5830.007 | 0.9830.010 | 0.850.10 | [12] |
J00164396+2304267 | T0.0 | 0.7040.012 | 1.0050.013 | 0.7980.015 | 0.5590.010 | 0.5340.011 | 0.6750.013 | 0.5780.010 | 0.5730.009 | 0.4490.006 | 1.0100.012 | 1.040.06 | [28] |
0.7520.012 | 1.0460.015 | 0.7960.011 | 0.5550.011 | 0.5640.012 | 0.6680.012 | 0.6020.013 | 0.5940.008 | 0.4650.006 | 1.0450.011 | 0.990.07 | [29] | ||
J00185151+5153306 | L7.0 | 0.8450.020 | 1.1220.019 | 1.0110.013 | 0.6620.017 | 0.7100.018 | 0.7980.011 | 0.4810.016 | 0.8290.017 | 0.8930.016 | 1.0150.016 | 1.200.15 | [28] |
J00311004+5749363 | L8 | 0.9000.009 | 1.0300.009 | 0.8430.007 | 0.6530.007 | 0.6580.007 | 0.6870.007 | 0.5820.009 | 0.6550.007 | 0.5910.005 | 1.0280.007 | 0.990.15 | [20] |
0.9400.007 | 1.0730.007 | 0.8370.005 | 0.6530.005 | 0.6560.007 | 0.6780.005 | 0.6000.006 | 0.6510.005 | 0.5700.004 | 1.0210.005 | 0.980.15 | [34] | ||
J00325937+1410371 | L8 | 0.8840.014 | 1.0320.014 | 0.8400.013 | 0.6350.013 | 0.6730.014 | 0.6930.012 | 0.5910.014 | 0.6560.009 | 0.6260.009 | 1.0330.011 | 0.990.13 | [11] |
0.8760.010 | 1.0340.010 | 0.8380.009 | 0.6950.010 | 0.6640.009 | 0.7140.009 | 0.6690.013 | 0.6550.007 | 0.5700.006 | 1.0230.009 | 0.950.13 | [11] | ||
J00345157+0523050 | T6.5 | 0.4200.007 | 0.2270.004 | 0.1200.007 | 0.1020.001 | 0.2090.006 | 0.4270.009 | 0.4500.005 | 0.2170.002 | 0.1000.001 | 0.4850.015 | 0.820.21 | [1] |
J00440332+0228112 | L7p (red) | 0.8570.012 | 1.2150.017 | 1.1400.010 | 0.6640.011 | 0.7030.012 | 0.8880.013 | 0.4370.010 | 0.8340.012 | 0.8140.010 | 1.0170.011 | 1.210.18 | [33] |
0.9360.021 | 1.2120.026 | 1.1100.015 | 0.6600.022 | 0.7290.020 | 0.8720.013 | 0.4540.018 | 0.8300.016 | 0.7680.019 | 0.9960.016 | 1.180.19 | [32] | ||
J00454256+3611391 | T5 | 0.4080.009 | 0.3930.008 | 0.2010.013 | 0.1970.003 | 0.3250.007 | 0.4570.010 | 0.5550.007 | 0.3790.005 | 0.1750.002 | 0.6160.013 | 1.020.13 | [21] |
J00492826+0440575 | L9 | 0.9160.008 | 1.0310.008 | 0.8540.006 | 0.7280.007 | 0.7080.007 | 0.6950.007 | 0.5990.006 | 0.6620.004 | 0.6390.005 | 1.0110.006 | 0.910.15 | [16] |
J00550564+0134365 | L7 | 0.9200.009 | 1.1520.011 | 1.1010.007 | 0.7590.007 | 0.7860.009 | 0.8700.009 | 0.4910.006 | 0.7350.006 | 0.6240.005 | 1.0070.008 | 0.930.12 | [33] |
J00575764+2013040 | sdL7 | 0.8370.012 | 1.0240.015 | 0.9480.017 | 0.6570.011 | 0.5870.014 | 0.7000.016 | 0.5690.014 | 0.5330.009 | 0.3820.007 | 0.9670.012 | 0.850.13 | [25] |
0.7170.006 | 1.0490.006 | 0.9080.006 | 0.6940.004 | 0.6700.008 | 0.6680.006 | 0.5230.006 | 0.5980.004 | 0.3920.003 | 0.9890.004 | 0.870.11 | [23] | ||
J01020186+0355405 | L9 | 0.9370.019 | 1.0500.016 | 0.8150.014 | 0.6370.018 | 0.5920.014 | 0.6560.012 | 0.6280.040 | 0.6660.011 | 0.6320.010 | 1.0420.014 | 1.070.19 | [29] |
J01352531+0205232 | L9.5 | 1.0450.009 | 1.0940.010 | 0.9000.009 | 0.6740.009 | 0.6060.011 | 0.6160.012 | 0.6090.008 | 0.6360.007 | 0.5550.006 | 1.0540.008 | 0.980.13 | [33] |
J01365662+0933473 | T2.5 | 0.7540.004 | 0.8400.003 | 0.5520.003 | 0.4930.001 | 0.4820.003 | 0.5900.002 | 0.5310.004 | 0.4480.001 | 0.3070.001 | 0.9700.003 | 1.030.07 | [9] |
J013836580322212 | T3 | 0.6380.008 | 0.7600.008 | 0.6840.011 | 0.4100.006 | 0.5410.008 | 0.6320.011 | 0.5520.008 | 0.4720.005 | 0.3080.003 | 0.9010.009 | 0.830.06 | [16] |
J014148391601196 | L7: | 0.9840.016 | 1.0800.022 | 0.9420.019 | 0.7200.019 | 0.6510.028 | 0.6900.017 | 0.5620.020 | 0.7150.017 | 0.6360.012 | 0.9610.018 | 1.030.15 | [12] |
J014535200314117 | L9 | 0.6860.011 | 1.0030.010 | 0.8490.008 | 0.7030.009 | 0.7150.010 | 0.7220.010 | 0.6810.010 | 0.6740.007 | 0.6270.006 | 1.0310.008 | 0.940.13 | [33] |
J01501089+3827241 | T0 | 0.8480.010 | 0.9730.006 | 0.7950.006 | 0.6700.005 | 0.6620.006 | 0.6740.005 | 0.5940.007 | 0.5960.004 | 0.5160.003 | 1.0060.005 | 0.890.07 | [16] |
0.9040.032 | 0.9970.021 | 0.8070.013 | 0.7060.025 | 0.6610.018 | 0.6760.012 | 0.6220.030 | 0.6340.012 | 0.5970.016 | 1.0000.015 | 0.930.07 | [31] | ||
J01514155+1244300 | T1 | 0.8210.011 | 0.9490.008 | 0.7070.007 | 0.6710.005 | 0.6250.008 | 0.6150.008 | 0.6080.008 | 0.5120.004 | 0.4310.002 | 1.0260.007 | 0.850.10 | [1] |
J02022929+2305139 | L7 | 0.9740.022 | 1.1880.019 | 1.0500.013 | 0.6690.018 | 0.6760.015 | 0.7530.011 | 0.4670.014 | 0.7980.014 | 0.7780.013 | 1.0040.016 | 1.170.19 | [32] |
J02062527+2640236 | L9p (red) | 0.8900.012 | 1.0760.008 | 0.9350.007 | 0.6800.008 | 0.6830.008 | 0.7310.006 | 0.5740.010 | 0.8160.007 | 0.8540.008 | 1.0310.007 | 1.190.14 | [16] |
J02101025+4008296 | T4.5 | 0.5360.010 | 0.5020.011 | 0.2480.017 | 0.3310.005 | 0.3650.013 | 0.4720.014 | 0.4770.009 | 0.3660.006 | 0.1970.003 | 0.6700.012 | 1.020.07 | [21] |
J022609161610004 | L7p (red) | 0.9650.023 | 1.1980.026 | 1.0670.013 | 0.6860.024 | 0.7480.023 | 0.9000.012 | 0.4250.024 | 0.9480.022 | 0.9950.022 | 0.9920.019 | 1.310.16 | [32] |
J023038850225471 | L8:p | 0.9240.015 | 1.0570.017 | 0.8400.012 | 0.6330.013 | 0.6130.015 | 0.7050.011 | 0.6580.018 | 0.7080.012 | 0.7200.011 | 1.0510.013 | 1.120.17 | [34] |
J024313712453298 | T6 | 0.4030.006 | 0.3320.005 | 0.1570.005 | 0.1410.002 | 0.2820.005 | 0.4550.007 | 0.4460.005 | 0.3860.003 | 0.1910.001 | 0.5380.009 | 1.070.15 | [1] |
J025503574700509 | L8 | 0.8300.013 | 1.0340.004 | 0.8370.005 | 0.6810.003 | 0.6800.005 | 0.6860.003 | 0.5670.005 | 0.6470.004 | 0.5930.003 | 1.0260.004 | 0.940.14 | [3] |
J025725813105523 | L8 | 0.9150.010 | 1.0540.010 | 0.8990.007 | 0.7170.006 | 0.7330.007 | 0.7230.007 | 0.5830.007 | 0.7120.006 | 0.6850.008 | 1.0150.010 | 0.970.10 | [5] |
J03162759+2650277 | T2.5 | 0.6400.010 | 0.7670.009 | 0.5770.012 | 0.4380.005 | 0.4780.009 | 0.5410.011 | 0.5650.009 | 0.4320.005 | 0.2820.003 | 0.9310.009 | 0.950.06 | [31] |
J03230186+5625580 | L7 | 0.9290.012 | 1.1030.010 | 1.0090.009 | 0.6670.012 | 0.6880.010 | 0.7670.007 | 0.4890.012 | 0.7720.008 | 0.7700.008 | 1.0150.008 | 1.130.18 | [25] |
J032440231919056 | L8p (blue) | 0.9500.014 | 1.0200.013 | 0.8840.013 | 0.7330.013 | 0.6870.015 | 0.7150.011 | 0.5870.011 | 0.6300.009 | 0.5720.007 | 0.9990.013 | 0.880.12 | [32] |
J03255322+0425406 | T5.5 | 0.4990.005 | 0.3670.008 | 0.2240.012 | 0.2020.005 | 0.3110.011 | 0.4550.014 | 0.4450.011 | 0.2860.005 | 0.1440.003 | 0.5900.011 | 0.900.09 | [2] |
J03284265+2302051 | L8 | 0.9070.010 | 0.9970.011 | 0.8280.009 | 0.6450.008 | 0.6360.011 | 0.6520.008 | 0.6010.010 | 0.6170.006 | 0.5730.006 | 1.0050.008 | 0.960.07 | [9] |
J03490944+1514360 | L7p | 1.1070.020 | 1.1810.021 | 1.0860.013 | 0.6780.020 | 0.7310.019 | 0.9050.011 | 0.4690.019 | 0.9830.015 | 1.2000.025 | 1.0200.017 | 1.380.20 | [28] |
J04070885+1514565 | T5 | 0.5210.010 | 0.3690.006 | 0.1670.005 | 0.2360.003 | 0.3490.008 | 0.4180.009 | 0.4150.011 | 0.4020.003 | 0.2040.002 | 0.5910.016 | 1.020.16 | [1] |
J041519540935066 | T8 | 0.2320.005 | 0.1060.003 | 0.0540.006 | 0.0380.002 | 0.1590.004 | 0.3190.007 | 0.3480.005 | 0.2740.002 | 0.1500.001 | 0.2200.007 | 1.200.20 | [1] |
J04373705+2331080 | L7.0 | 1.0180.021 | 1.1890.031 | 1.1400.019 | 0.7920.019 | 0.7660.024 | 0.9740.017 | 0.6060.025 | 0.7600.018 | 0.7060.015 | 1.0000.019 | 0.960.11 | [24] |
J045746020207179 | T2 | 0.7080.006 | 0.8220.005 | 0.5220.005 | 0.4560.002 | 0.4800.006 | 0.5830.005 | 0.4870.005 | 0.4720.003 | 0.3050.002 | 0.9180.005 | 0.990.05 | [34] |
J05065020+5236312 | T4.5 | 0.5150.009 | 0.4960.011 | 0.2480.012 | 0.2540.004 | 0.3200.011 | 0.4580.015 | 0.4500.008 | 0.3720.007 | 0.1900.003 | 0.7240.017 | 1.030.07 | [28] |
J051609450445499 | T5.5 | 0.4800.007 | 0.3360.005 | 0.1590.007 | 0.1940.003 | 0.2960.005 | 0.4160.007 | 0.3940.005 | 0.3630.003 | 0.1820.001 | 0.5100.009 | 1.060.10 | [9] |
J055919141404488 | T4.5 | 0.5920.004 | 0.4820.009 | 0.2400.007 | 0.3470.004 | 0.4010.012 | 0.5020.008 | 0.4600.005 | 0.3510.004 | 0.1840.001 | 0.6710.011 | 0.970.07 | [3] |
J06020638+4043588 | T4.5 | 0.5720.009 | 0.4610.008 | 0.2520.016 | 0.3020.005 | 0.3440.009 | 0.4690.013 | 0.4710.008 | 0.3550.005 | 0.1790.002 | 0.7160.014 | 1.010.07 | [6] |
J06073865+2429535 | L8 | 0.9300.004 | 1.0200.002 | 0.8470.002 | 0.6870.001 | 0.6730.003 | 0.7010.002 | 0.5340.003 | 0.6800.002 | 0.6560.002 | 1.0130.002 | 0.990.16 | [22] |
0.8930.006 | 1.0030.003 | 0.8390.002 | 0.7000.002 | 0.6830.004 | 0.7030.003 | 0.5810.004 | 0.6440.002 | 0.5970.002 | 1.0160.002 | 0.920.12 | [19] | ||
J061135130410240 | T0 | 0.8100.015 | 0.9250.015 | 0.7430.018 | 0.5840.013 | 0.6000.015 | 0.6620.017 | 0.5660.016 | 0.5530.010 | 0.4330.007 | 1.0210.017 | 0.950.05 | [16] |
J062720071114280 | T6 | 0.3990.008 | 0.3240.007 | 0.1550.013 | 0.1580.003 | 0.2780.009 | 0.4520.011 | 0.4250.007 | 0.3580.005 | 0.1680.003 | 0.4930.013 | 1.050.11 | [16] |
J06420559+4102000 | L8.0 | 0.8220.011 | 1.0480.012 | 0.8380.008 | 0.5650.009 | 0.6000.011 | 0.6770.008 | 0.5890.011 | 0.7690.010 | 0.7610.008 | 1.0090.012 | 0.960.10 | [28] |
J070534001839256 | L8 | 0.9590.025 | 1.0630.026 | 0.9390.013 | 0.7680.021 | 0.7280.022 | 0.7920.014 | 0.6380.020 | 0.6820.015 | 0.6700.014 | 1.0540.017 | 0.900.11 | [32] |
J07271824+1710012 | T7 | 0.2960.005 | 0.1830.005 | 0.0610.007 | 0.0850.003 | 0.2060.010 | 0.3480.009 | 0.4030.008 | 0.3410.003 | 0.1600.002 | 0.3330.013 | 1.190.12 | [4] |
J07414920+2351275 | T5 | 0.5080.008 | 0.3620.008 | 0.1670.012 | 0.1960.003 | 0.2990.009 | 0.4410.012 | 0.4280.007 | 0.3180.005 | 0.1480.002 | 0.5870.017 | 0.970.07 | [11] |
J07420130+2055198 | T5 | 0.4810.007 | 0.3990.005 | 0.1950.011 | 0.2450.003 | 0.3170.007 | 0.4770.010 | 0.4480.007 | 0.3160.003 | 0.1570.001 | 0.5970.010 | 0.970.08 | [11] |
J07483861+1743329 | L7 | 0.9840.010 | 1.0840.013 | 0.9720.010 | 0.7350.009 | 0.6790.010 | 0.7350.009 | 0.4900.011 | 0.6650.010 | 0.5750.006 | 0.9680.010 | 0.930.14 | [26] |
J07554795+2212169 | T5 | 0.4980.012 | 0.4030.006 | 0.2090.015 | 0.2330.004 | 0.3270.011 | 0.4890.015 | 0.4590.009 | 0.3400.006 | 0.1550.003 | 0.5960.014 | 0.990.08 | [3] |
J07584037+3247245 | T2 | 0.6820.005 | 0.9270.005 | 0.6290.005 | 0.4450.002 | 0.4480.005 | 0.5620.003 | 0.5670.005 | 0.5030.003 | 0.3560.001 | 1.0060.004 | 1.090.12 | [9] |
J08070023+4130268 | L8p | 0.9430.007 | 1.0560.005 | 0.9270.005 | 0.7610.004 | 0.7720.005 | 0.7500.004 | 0.5640.006 | 0.6920.004 | 0.6310.003 | 1.0050.004 | 0.890.11 | [22] |
J081958050335285 | T4 | 0.6340.010 | 0.6450.008 | 0.3600.009 | 0.3850.004 | 0.3780.007 | 0.4950.011 | 0.4790.006 | 0.3460.004 | 0.2060.002 | 0.8340.009 | 1.000.11 | [16] |
J08203013+1037372 | L9.5: | 0.7930.012 | 1.0600.017 | 0.8800.012 | 0.6100.016 | 0.6240.017 | 0.6840.010 | 0.6040.013 | 0.6480.012 | 0.5830.010 | 1.0100.020 | 1.040.17 | [2] |
J08300825+4828482 | L8 | 0.9280.010 | 1.0780.009 | 0.8720.006 | 0.6040.008 | 0.5900.007 | 0.5630.005 | 0.5320.008 | 0.6660.006 | 0.6330.006 | 1.0370.006 | 1.100.17 | [9] |
0.8470.008 | 1.0670.006 | 0.8750.004 | 0.6420.004 | 0.5860.006 | 0.5670.004 | 0.6180.005 | 0.6480.004 | 0.5850.003 | 1.0370.005 | 1.040.15 | [9] | ||
J08304878+0128311 | T4.5 | 0.4990.009 | 0.4570.007 | 0.2320.012 | 0.2590.003 | 0.3410.009 | 0.4840.014 | 0.4540.007 | 0.3330.004 | 0.1640.002 | 0.6750.011 | 0.950.06 | [11] |
J083717180000179 | T1 | 0.8580.012 | 0.9840.012 | 0.7210.013 | 0.5880.010 | 0.5350.012 | 0.6220.013 | 0.6320.013 | 0.5050.007 | 0.3860.006 | 1.0170.010 | 1.070.09 | [34] |
J08523490+4720359 | L9.5: | 0.8670.010 | 1.0220.009 | 0.8450.009 | 0.6390.009 | 0.6470.010 | 0.6720.010 | 0.6280.011 | 0.6390.006 | 0.5490.005 | 1.0160.009 | 0.980.15 | [11] |
J08575849+5708514 | L8 | 0.9490.005 | 1.1060.005 | 0.9260.004 | 0.6390.004 | 0.6350.005 | 0.7000.003 | 0.5250.005 | 0.8190.003 | 0.8710.004 | 1.0280.004 | 1.270.17 | [11] |
J08583467+3256275 | T1 | 0.8100.011 | 0.9950.011 | 0.8990.011 | 0.5890.009 | 0.5510.010 | 0.6240.008 | 0.6500.013 | 0.6420.006 | 0.5540.006 | 0.9800.009 | 1.080.06 | [11] |
0.8130.014 | 0.9790.017 | 0.9050.018 | 0.5650.011 | 0.5970.016 | 0.6240.017 | 0.5700.022 | 0.6300.012 | 0.5540.007 | 0.9960.015 | 0.970.06 | [2] | ||
J09002368+2539343 | L7 | 0.9380.009 | 1.0670.009 | 0.9320.010 | 0.7410.008 | 0.6670.009 | 0.6800.008 | 0.5320.008 | 0.5810.006 | 0.4450.004 | 1.0010.008 | 0.810.10 | [26] |
J09025899+6708331 | L7p (red) | 1.0440.008 | 1.1780.012 | 1.1210.006 | 0.6780.008 | 0.6700.009 | 0.7700.006 | 0.3980.008 | 0.7650.007 | 0.7480.007 | 0.9970.007 | 1.120.18 | [32] |
J09293364+3429527 | L8 | 0.9020.012 | 1.0990.015 | 0.9460.010 | 0.6930.011 | 0.6940.011 | 0.7260.008 | 0.6160.014 | 0.8010.010 | 0.8060.009 | 1.0420.010 | 1.140.17 | [11] |
J09310955+0327331 | L7.51.5 | 0.8260.012 | 1.1020.011 | 0.9770.016 | 0.6720.009 | 0.5850.011 | 0.6530.013 | 0.4890.010 | 0.5320.007 | 0.3460.004 | 1.0310.009 | 0.840.12 | [34] |
J09373487+2931409 | T6p | 0.4300.011 | 0.3060.007 | 0.1830.012 | 0.1480.002 | 0.3050.006 | 0.5320.022 | 0.5320.009 | 0.1700.003 | 0.0740.001 | 0.5330.011 | 0.620.16 | [4] |
J095533360208403 | L7p (red) | 0.9040.022 | 1.0960.026 | 0.9820.015 | 0.6770.026 | 0.7110.026 | 0.7570.013 | 0.5160.023 | 0.8270.019 | 0.8840.021 | 1.0220.024 | 1.180.19 | [32] |
J095608101447065 | L9.5 | 0.8250.011 | 1.0380.012 | 0.8080.009 | 0.5790.010 | 0.5400.011 | 0.6530.009 | 0.6090.010 | 0.6580.008 | 0.5860.006 | 1.0180.009 | 1.070.07 | [31] |
J10071185+1930563 | L8: | 0.8930.026 | 1.0360.014 | 0.8350.020 | 0.6400.015 | 0.6560.016 | 0.6840.016 | 0.6300.016 | 0.6330.015 | 0.5410.015 | 1.0430.018 | 0.970.14 | [2] |
0.9000.011 | 1.0610.014 | 0.8570.012 | 0.6210.011 | 0.6430.013 | 0.6510.010 | 0.5980.014 | 0.7360.010 | 0.7030.008 | 1.0570.012 | 1.150.19 | [11] | ||
J10255227+3212349 | L7.5:: | 0.9030.019 | 1.0270.014 | 0.9430.012 | 0.6900.019 | 0.7030.013 | 0.7150.016 | 0.5540.017 | 0.6930.008 | 0.6560.009 | 0.9910.011 | 0.980.19 | [2] |
0.9130.016 | 1.0520.016 | 0.9080.014 | 0.7280.020 | 0.7360.014 | 0.7400.013 | 0.5870.019 | 0.7040.011 | 0.6980.011 | 1.0040.012 | 0.950.11 | [26] | ||
J10271549+5445175 | L7 | 0.9850.021 | 1.1460.019 | 1.1050.013 | 0.8090.018 | 0.7930.016 | 0.9170.016 | 0.5770.013 | 0.6820.011 | 0.5720.012 | 1.0060.011 | 0.830.08 | [33] |
J10302678+0213057 | L9.5: | 0.9120.015 | 1.0230.013 | 0.8580.013 | 0.6640.011 | 0.6240.012 | 0.7060.012 | 0.5430.014 | 0.6220.008 | 0.5620.008 | 1.0030.013 | 0.960.11 | [11] |
J10430758+2225236 | L8 | 0.9110.015 | 1.1070.010 | 0.8570.009 | 0.6150.010 | 0.6060.010 | 0.6750.007 | 0.5620.013 | 0.7720.007 | 0.7860.007 | 1.0390.009 | 1.250.19 | [5] |
J10433508+1213149 | L9 | 0.9320.011 | 1.0380.010 | 0.8550.008 | 0.6610.011 | 0.6730.011 | 0.6970.008 | 0.6170.010 | 0.6640.008 | 0.6280.008 | 1.0200.009 | 0.980.15 | [12] |
0.9030.007 | 1.0310.006 | 0.8420.005 | 0.6750.005 | 0.6460.006 | 0.6690.005 | 0.5820.007 | 0.6790.004 | 0.6640.004 | 1.0190.004 | 1.020.16 | [11] | ||
J10440942+0429376 | L7 | 0.8900.009 | 1.0310.006 | 0.8740.005 | 0.7100.006 | 0.6840.006 | 0.7060.006 | 0.5450.007 | 0.6670.004 | 0.6310.004 | 1.0170.006 | 0.950.14 | [11] |
J10475385+2124234 | T6.5 | 0.4040.005 | 0.2610.005 | 0.1220.009 | 0.1190.002 | 0.2510.006 | 0.4240.010 | 0.4350.006 | 0.2370.003 | 0.1090.001 | 0.4870.011 | 0.830.12 | [9] |
J10482926+0919373 | T2.5 | 0.7430.013 | 0.8310.010 | 0.5580.015 | 0.4910.008 | 0.4570.012 | 0.5760.014 | 0.5810.012 | 0.4100.006 | 0.2880.004 | 0.9470.013 | 1.060.10 | [11] |
J104915575319061A | L8 | 0.9300.007 | 1.1120.006 | 0.9280.004 | 0.6030.005 | 0.6390.006 | 0.6920.003 | 0.4880.005 | 0.8460.005 | 0.9130.005 | 1.0270.003 | 1.350.24 | [18] |
J104915575319061B | T0.5 | 0.8320.008 | 1.0590.006 | 0.8340.005 | 0.6110.004 | 0.5510.006 | 0.6590.003 | 0.6340.005 | 0.6090.004 | 0.5150.003 | 1.0450.004 | 1.030.19 | [18] |
J10595185+3042059 | T4 | 0.6240.014 | 0.5230.009 | 0.2770.012 | 0.3840.007 | 0.3870.011 | 0.4760.014 | 0.4700.010 | 0.4020.007 | 0.2400.005 | 0.7250.012 | 1.080.07 | [10] |
J112208263512363 | T2 | 0.7860.013 | 0.7810.008 | 0.5050.009 | 0.5300.005 | 0.5370.012 | 0.5470.009 | 0.5470.007 | 0.3860.004 | 0.2830.002 | 0.9260.007 | 0.920.11 | [3] |
J114724212040204 | L7 | 0.9350.029 | 1.2550.030 | 1.1080.014 | 0.5490.027 | 0.6720.026 | 0.7790.012 | 0.4190.027 | 0.9860.028 | 1.0980.029 | 1.0560.025 | 0.860.20 | [30] |
J11555389+0559577 | L7.5 | 0.8890.008 | 1.0610.006 | 0.9110.007 | 0.7160.005 | 0.6910.006 | 0.7020.005 | 0.5580.007 | 0.6470.004 | 0.5550.004 | 1.0110.005 | 0.910.12 | [11] |
J11582077+0435014 | sdL7 | 0.8620.009 | 1.0390.007 | 0.8890.008 | 0.7090.005 | 0.6380.007 | 0.6860.007 | 0.4700.006 | 0.4960.005 | 0.3210.002 | 1.0060.006 | 0.730.09 | [12] |
J12010457+5730042 | L9 | 0.9590.018 | 1.1030.028 | 0.8710.017 | 0.6140.020 | 0.5980.022 | 0.6190.014 | 0.6230.020 | 0.6550.016 | 0.5970.012 | 1.0570.019 | 1.070.16 | [32] |
J12060248+2813293 | T3 | 0.7160.006 | 0.6680.011 | 0.4290.013 | 0.4570.007 | 0.5480.018 | 0.5400.018 | 0.6170.014 | 0.3210.006 | 0.2100.003 | 0.8920.019 | 0.790.10 | [2] |
J12144089+6316434 | T3.5: | 0.5970.015 | 0.7060.016 | 0.4210.016 | 0.2780.005 | 0.3260.011 | 0.5020.013 | 0.4920.007 | 0.3850.007 | 0.2440.003 | 0.9320.016 | 1.120.15 | [2] |
J12195156+3128497 | L8 | 0.8980.009 | 1.0650.009 | 0.8480.005 | 0.6370.006 | 0.6380.009 | 0.6570.005 | 0.5640.007 | 0.7170.006 | 0.6830.005 | 1.0090.006 | 1.110.14 | [11] |
J12314753+0847331 | T5.5 | 0.4460.013 | 0.3570.005 | 0.1540.007 | 0.1820.002 | 0.2640.005 | 0.3700.006 | 0.4510.005 | 0.3220.004 | 0.1550.002 | 0.5650.010 | 1.020.09 | [1] |
J125453930122474 | T2 | 0.7750.009 | 0.9140.006 | 0.5940.005 | 0.4700.005 | 0.4600.007 | 0.5450.006 | 0.5280.006 | 0.4900.004 | 0.3610.003 | 1.0150.005 | 1.090.09 | [1] |
J130729840558146 | L8 | 0.9500.010 | 1.0200.012 | 0.8980.010 | 0.7230.013 | 0.6940.014 | 0.7380.008 | 0.5990.011 | 0.6650.008 | 0.6230.006 | 1.0170.011 | 0.930.12 | [29] |
J13184567+3626138 | L7 | 0.9360.019 | 1.1510.028 | 1.1300.023 | 0.7440.021 | 0.7850.025 | 0.8930.026 | 0.4870.016 | 0.7090.017 | 0.6510.016 | 0.9870.017 | 0.910.10 | [33] |
J13240767+1906259 | T3.5 | 0.6430.007 | 0.6230.006 | 0.3560.009 | 0.4340.003 | 0.4040.006 | 0.4880.009 | 0.6050.007 | 0.3940.004 | 0.2140.002 | 0.8580.010 | 1.050.09 | [13] |
J132629810038314 | L8: | 0.9840.010 | 1.0940.008 | 0.9980.007 | 0.6670.010 | 0.7100.010 | 0.7920.007 | 0.5160.013 | 0.7950.006 | 0.8060.009 | 0.9990.007 | 1.140.20 | [11] |
J134646340031501 | T6.5 | 0.4050.009 | 0.2790.007 | 0.1230.010 | 0.1300.004 | 0.2660.007 | 0.4250.010 | 0.4190.009 | 0.3400.005 | 0.1540.002 | 0.4560.018 | 1.050.11 | [4] |
J13480699+6603284 | L9 | 1.0120.013 | 1.0620.015 | 0.8500.012 | 0.5920.012 | 0.6160.014 | 0.6330.015 | 0.5570.013 | 0.6540.010 | 0.6500.009 | 0.9960.013 | 0.960.07 | [16] |
J13585269+3747137 | T4.5: | 0.5660.005 | 0.3760.011 | 0.3030.017 | 0.2520.003 | 0.3470.010 | 0.5100.036 | 0.4820.012 | 0.2670.006 | 0.1220.003 | 0.6680.012 | 0.840.10 | [2] |
J14002320+4338222 | L7: | 0.9520.013 | 1.0410.008 | 0.8880.007 | 0.7760.009 | 0.7480.008 | 0.7500.006 | 0.5940.010 | 0.7330.006 | 0.7240.007 | 0.9990.007 | 0.950.13 | [11] |
J14321117+3244338 | L7 | 0.9860.026 | 1.1730.022 | 1.0440.014 | 0.7340.021 | 0.6590.021 | 0.7760.014 | 0.4870.021 | 0.7720.013 | 0.7020.018 | 0.9940.016 | 1.100.13 | [32] |
J14564268+6450097 | L7p (red) | 1.0340.016 | 1.1560.015 | 1.0860.011 | 0.7030.015 | 0.6990.014 | 0.7630.010 | 0.4700.015 | 0.7850.011 | 0.8090.013 | 0.9640.010 | 1.110.16 | [32] |
J15031961+2525196 | T5 | 0.4650.013 | 0.4110.004 | 0.2010.006 | 0.2380.001 | 0.3190.006 | 0.4530.005 | 0.4590.005 | 0.3240.002 | 0.1510.001 | 0.5900.006 | 0.980.07 | [1] |
J150502372853452 | T1.5 | 0.7850.010 | 0.8610.007 | 0.6690.009 | 0.5390.005 | 0.5300.009 | 0.6170.009 | 0.5750.007 | 0.4930.005 | 0.3880.004 | 0.9840.008 | 1.000.06 | [13] |
J15131461+4019356 | L8 | 0.9100.017 | 1.0590.022 | 0.9170.018 | 0.6660.020 | 0.6460.021 | 0.7350.016 | 0.6000.027 | 0.7410.016 | 0.7120.015 | 1.0320.021 | 1.120.15 | [22] |
J15150607+4436483 | L7.5: | 0.9510.011 | 1.1060.010 | 0.9440.008 | 0.6790.008 | 0.6910.009 | 0.7320.009 | 0.5040.011 | 0.7390.007 | 0.7040.008 | 1.0180.009 | 1.070.16 | [11] |
J15164306+3053443 | T0.5: | 0.7500.014 | 1.1450.014 | 0.8430.013 | 0.4730.011 | 0.5070.014 | 0.6000.010 | 0.5370.013 | 0.7550.011 | 0.6730.008 | 1.0580.011 | 0.980.13 | [11] |
J15203974+3546210 | T0: | 0.9390.020 | 1.0200.015 | 0.7950.011 | 0.7160.016 | 0.6930.016 | 0.6900.010 | 0.6430.016 | 0.6130.009 | 0.5430.009 | 1.0170.012 | 0.860.11 | [34] |
0.9260.005 | 1.0180.005 | 0.8080.004 | 0.7050.005 | 0.6880.005 | 0.6750.004 | 0.5900.005 | 0.6320.004 | 0.5750.004 | 1.0100.004 | 0.900.13 | [11] | ||
J15210327+0131426 | T2: | 0.7190.009 | 0.8410.013 | 0.4820.014 | 0.4440.008 | 0.4560.013 | 0.5580.013 | 0.5300.010 | 0.4630.008 | 0.3030.005 | 0.9750.012 | 1.040.08 | [11] |
J15232263+3014562 | L8 | 0.9130.014 | 1.0300.011 | 0.8790.009 | 0.6650.011 | 0.6670.011 | 0.7100.009 | 0.5890.013 | 0.7080.008 | 0.7150.008 | 0.9860.009 | 1.050.16 | [11] |
J15261405+2043414 | L7 | 0.9200.008 | 1.0720.006 | 0.9350.005 | 0.7620.005 | 0.7190.007 | 0.7240.005 | 0.5370.007 | 0.6530.005 | 0.5500.004 | 0.9960.007 | 0.870.12 | [1] |
J15335852+4757069 | L8: | 1.0120.027 | 1.0910.025 | 0.9620.015 | 0.6670.022 | 0.7180.018 | 0.6950.015 | 0.5610.019 | 0.7460.015 | 0.7090.017 | 1.0120.020 | 1.070.15 | [32] |
J15400942+3742316 | L9: | 0.8220.022 | 1.0450.019 | 0.8840.015 | 0.6240.016 | 0.6750.017 | 0.6720.010 | 0.6310.014 | 0.7060.015 | 0.6330.013 | 1.0700.013 | 1.070.20 | [2] |
0.8910.010 | 1.0820.011 | 0.8910.009 | 0.6550.009 | 0.6710.009 | 0.6140.008 | 0.6380.010 | 0.6860.006 | 0.6200.006 | 1.0340.008 | 1.020.16 | [11] | ||
J154236300045452 | L7 | 0.8020.012 | 1.1430.021 | 1.0250.022 | 0.7720.012 | 0.6100.018 | 0.7060.021 | 0.6410.015 | 0.5600.014 | 0.3600.008 | 1.0020.017 | 0.800.12 | [15] |
J15450901+3555271 | L7.5 | 0.9880.018 | 1.0170.014 | 0.9070.014 | 0.7360.016 | 0.6960.014 | 0.7490.013 | 0.6030.017 | 0.6900.012 | 0.7080.011 | 0.9920.012 | 0.950.13 | [34] |
J15461460+49321139 | T2.5: | 0.6790.010 | 0.7550.010 | 0.4930.011 | 0.4450.005 | 0.4750.009 | 0.5560.011 | 0.5060.009 | 0.3980.006 | 0.2800.003 | 0.8930.011 | 0.960.07 | [11] |
J15462718332511 | T5.5 | 0.5070.008 | 0.3770.009 | 0.1790.011 | 0.2200.003 | 0.2850.010 | 0.4360.012 | 0.4030.007 | 0.3320.004 | 0.1520.002 | 0.6270.015 | 1.020.09 | [9] |
J16241436+0029158 | T6 | 0.4610.008 | 0.3030.005 | 0.1610.006 | 0.1510.001 | 0.2700.008 | 0.4120.009 | 0.4060.006 | 0.3070.003 | 0.1410.001 | 0.4950.010 | 0.990.08 | [4] |
J16272565+3255246 | T6 | 0.4050.007 | 0.2940.004 | 0.1800.014 | 0.1360.002 | 0.2960.008 | 0.4490.013 | 0.4140.006 | 0.2320.003 | 0.1050.001 | 0.4760.010 | 0.790.09 | [16] |
J16291855+0335355 | T2 | 0.7380.006 | 0.9160.006 | 0.6180.005 | 0.4410.003 | 0.4200.006 | 0.5610.006 | 0.5620.005 | 0.4920.003 | 0.3540.002 | 1.0140.006 | 1.140.11 | [17] |
J16303054+4344032 | L7: | 0.9190.012 | 1.0620.010 | 0.8790.007 | 0.6960.010 | 0.6700.008 | 0.7060.006 | 0.6020.010 | 0.7290.007 | 0.6980.006 | 1.0220.007 | 1.060.17 | [11] |
J16322911+1904407 | L8 | 0.9430.009 | 1.0590.011 | 0.8830.010 | 0.7100.012 | 0.6770.009 | 0.6930.010 | 0.5940.008 | 0.7470.007 | 0.7420.008 | 1.0250.009 | 1.060.13 | [8] |
J16323934+4150048 | T1: | 0.8830.015 | 0.8960.015 | 0.5920.015 | 0.5970.012 | 0.5490.013 | 0.5680.016 | 0.5530.014 | 0.4680.010 | 0.3820.007 | 0.9920.016 | 1.020.10 | [11] |
J16471557+5632083 | L9p (red) | 0.8390.014 | 1.0950.013 | 0.9430.010 | 0.6690.012 | 0.7020.012 | 0.7120.009 | 0.5600.011 | 0.8050.008 | 0.8170.010 | 1.0050.011 | 1.160.17 | [16] |
J17342104+5023508 | T4 | 0.4830.009 | 0.5200.007 | 0.2490.009 | 0.4000.004 | 0.4270.008 | 0.4860.010 | 0.4280.007 | 0.4020.004 | 0.2230.002 | 0.7140.010 | 1.030.09 | [21] |
J17385927+6142421 | L9p | 0.8960.021 | 1.0880.020 | 0.9430.011 | 0.5640.019 | 0.6260.019 | 0.7170.010 | 0.5300.019 | 0.9690.018 | 1.2250.024 | 1.0340.017 | 0.930.17 | [21] |
J17502385+4222373 | T2 | 0.8520.015 | 0.9130.012 | 0.6190.013 | 0.5040.011 | 0.5340.013 | 0.6090.010 | 0.5240.009 | 0.5080.007 | 0.3640.004 | 1.0210.012 | 0.970.07 | [3] |
J17503293+1759042 | T3.5 | 0.7140.010 | 0.6610.008 | 0.3480.013 | 0.4600.004 | 0.4350.009 | 0.5200.012 | 0.5180.008 | 0.3650.006 | 0.2300.003 | 0.8700.012 | 0.990.11 | [1] |
J17545447+1649196 | T5.5 | 0.4820.006 | 0.4050.005 | 0.2160.008 | 0.1910.002 | 0.2990.006 | 0.4300.007 | 0.4550.006 | 0.3630.004 | 0.1750.001 | 0.6300.010 | 1.020.09 | [9] |
J17551028+1803202 | T2 | 0.7550.010 | 0.8500.010 | 0.5480.011 | 0.5840.006 | 0.5270.010 | 0.5370.011 | 0.5970.008 | 0.4220.005 | 0.3080.004 | 0.9680.010 | 1.020.14 | [25] |
J18002660+0134531 | L7.5 | 0.9070.005 | 1.0390.003 | 0.8820.003 | 0.7270.002 | 0.7080.004 | 0.7140.002 | 0.6040.004 | 0.7160.002 | 0.6930.002 | 1.0230.002 | 0.980.13 | [14] |
J180952530448125 | T0.5 | 0.8350.009 | 0.9640.007 | 0.7310.007 | 0.6250.005 | 0.6150.007 | 0.6300.007 | 0.6020.007 | 0.5340.005 | 0.4280.004 | 1.0070.007 | 0.940.09 | [21] |
0.8610.015 | 0.9760.012 | 0.7230.009 | 0.6610.010 | 0.6350.011 | 0.6410.011 | 0.6560.012 | 0.5540.006 | 0.4380.006 | 1.0100.011 | 0.920.08 | [20] | ||
0.8610.014 | 0.9760.014 | 0.7230.010 | 0.6610.010 | 0.6350.012 | 0.6410.009 | 0.6560.012 | 0.5540.007 | 0.4380.006 | 1.0100.011 | 0.920.08 | [28] | ||
J18305856+4542574 | L9 | 0.8410.013 | 1.0550.014 | 0.8350.012 | 0.6560.013 | 0.6690.013 | 0.7240.011 | 0.5980.014 | 0.6920.011 | 0.6490.010 | 1.0570.012 | 1.030.16 | [16] |
J19010601+4718136 | T5 | 0.5530.008 | 0.4330.007 | 0.2480.008 | 0.2820.003 | 0.3530.008 | 0.5010.011 | 0.4900.005 | 0.3310.004 | 0.1700.002 | 0.6750.012 | 0.960.07 | [1] |
J19062474+4508071 | T6 | 0.4520.009 | 0.3230.007 | 0.2000.012 | 0.1870.003 | 0.2940.010 | 0.4710.016 | 0.4410.006 | 0.2830.005 | 0.1290.002 | 0.5430.012 | 0.920.09 | [16] |
J19284135+2356049 | T6 | 0.4160.004 | 0.3470.002 | 0.1450.003 | 0.1730.001 | 0.2720.003 | 0.3730.002 | 0.4260.003 | 0.3640.001 | 0.1720.001 | 0.5370.005 | 1.080.09 | [21] |
0.4420.005 | 0.3420.004 | 0.1580.004 | 0.1630.001 | 0.2780.005 | 0.3750.006 | 0.4160.004 | 0.3540.002 | 0.1690.001 | 0.5480.008 | 1.050.10 | [34] | ||
J20304279+0749347 | T1.5 | 0.8000.004 | 0.8770.003 | 0.5680.003 | 0.6000.002 | 0.5610.003 | 0.5780.002 | 0.5580.003 | 0.4450.001 | 0.3360.001 | 0.9750.003 | 0.870.12 | [21] |
0.7800.008 | 0.8680.007 | 0.5570.008 | 0.6220.007 | 0.5700.007 | 0.5760.009 | 0.6220.009 | 0.4300.004 | 0.3200.003 | 0.9810.007 | 0.840.12 | [20] | ||
J204317691551031 | L9 | 0.8700.013 | 1.0400.017 | 0.7710.015 | 0.6810.016 | 0.6660.015 | 0.6640.015 | 0.6330.014 | 0.6350.014 | 0.5640.011 | 1.0310.014 | 0.930.12 | [11] |
J20435642+6220490 | T1.5 | 0.7650.013 | 0.9330.013 | 0.6220.012 | 0.5610.009 | 0.5300.012 | 0.6050.013 | 0.6660.017 | 0.4530.006 | 0.3350.004 | 0.9960.014 | 1.020.11 | [21] |
J204749590718176 | T0: | 0.8820.014 | 1.0100.011 | 0.7890.012 | 0.6910.011 | 0.6830.011 | 0.6920.010 | 0.6390.014 | 0.6130.008 | 0.5620.007 | 1.0290.011 | 0.880.11 | [11] |
J205201742043119 | L8 | 0.9450.016 | 1.0390.019 | 0.9110.015 | 0.7420.016 | 0.7470.017 | 0.7400.010 | 0.6130.014 | 0.6520.010 | 0.5830.009 | 1.0230.016 | 0.860.11 | [29] |
J211714442940052 | T0: | 0.8540.018 | 0.9630.017 | 0.7540.020 | 0.6640.015 | 0.6460.016 | 0.6240.019 | 0.6770.020 | 0.5630.012 | 0.4980.011 | 1.0310.016 | 0.900.07 | [28] |
J21241387+0059599 | T5 | 0.5840.012 | 0.3830.008 | 0.1960.010 | 0.2390.004 | 0.3070.012 | 0.4120.012 | 0.4360.009 | 0.3770.004 | 0.1890.002 | 0.6340.011 | 1.080.08 | [3] |
J21243865+1849263 | L9 | 0.8490.012 | 1.0410.014 | 0.7980.009 | 0.5390.007 | 0.5280.011 | 0.5960.011 | 0.5900.009 | 0.6420.008 | 0.5760.007 | 1.0500.012 | 1.040.08 | [27] |
J213248981452544 | T4 | 0.7310.012 | 0.6380.014 | 0.3650.017 | 0.4120.008 | 0.3700.014 | 0.5640.022 | 0.5300.011 | 0.3480.007 | 0.2180.004 | 0.8950.020 | 0.980.15 | [34] |
J21392676+0220226 | T1.5 | 0.6640.010 | 1.0150.005 | 0.6910.003 | 0.4220.002 | 0.4540.006 | 0.5750.002 | 0.5140.004 | 0.5720.003 | 0.4050.002 | 1.0310.003 | 0.990.10 | [3] |
J21442847+1446077 | T2.5 | 0.6300.007 | 0.8420.008 | 0.5260.006 | 0.3800.004 | 0.4320.010 | 0.5420.007 | 0.5540.007 | 0.4600.004 | 0.3060.002 | 0.9760.008 | 1.020.12 | [7] |
J215138394853542 | T4 | 0.6150.008 | 0.6040.008 | 0.3070.013 | 0.3310.004 | 0.3500.008 | 0.4710.011 | 0.4170.008 | 0.3780.005 | 0.1920.002 | 0.7700.012 | 1.040.07 | [3] |
J21513979+3402444 | L7p | 0.9680.013 | 1.0490.010 | 0.9430.011 | 0.7530.010 | 0.7480.009 | 0.7700.011 | 0.5760.012 | 0.6990.007 | 0.6610.009 | 0.9990.009 | 0.920.11 | [12] |
J215424941023022 | T4.5 | 0.5850.009 | 0.5280.007 | 0.2820.015 | 0.3530.005 | 0.4010.008 | 0.4960.010 | 0.4550.007 | 0.3900.006 | 0.2120.003 | 0.7580.012 | 1.010.07 | [6] |
J220921832711329 | T2.5 | 0.8240.008 | 0.7190.007 | 0.4630.009 | 0.4600.004 | 0.4630.009 | 0.5570.009 | 0.4880.006 | 0.4300.004 | 0.3010.003 | 0.8730.011 | 1.000.07 | [34] |
J22153705+2110554 | T0 | 0.8140.009 | 0.9420.009 | 0.6910.007 | 0.5760.006 | 0.5720.009 | 0.6690.006 | 0.5020.008 | 0.5300.004 | 0.3820.003 | 0.9960.008 | 0.990.05 | [31] |
0.7850.005 | 0.9460.005 | 0.7030.004 | 0.5810.003 | 0.5710.006 | 0.6570.005 | 0.5320.005 | 0.5240.003 | 0.3690.002 | 1.0020.005 | 1.000.05 | [27] | ||
J22221993+3026014 | L9 | 0.9410.015 | 1.0420.012 | 0.8410.009 | 0.7490.012 | 0.7380.013 | 0.7090.009 | 0.6380.011 | 0.6810.008 | 0.6330.007 | 1.0370.012 | 0.900.11 | [26] |
J22361759+5105519 | T5.5 | 0.4660.005 | 0.4080.004 | 0.1940.004 | 0.2280.002 | 0.3220.006 | 0.4390.005 | 0.4360.005 | 0.3720.002 | 0.1810.001 | 0.6360.008 | 1.030.09 | [20] |
J22393755+1617161 | T3 | 0.6530.007 | 0.7300.006 | 0.5530.008 | 0.4180.004 | 0.4990.007 | 0.5800.008 | 0.5460.006 | 0.4110.004 | 0.2630.002 | 0.8900.007 | 0.910.08 | [16] |
J22541892+3123498 | T4 | 0.6730.010 | 0.5740.005 | 0.3140.005 | 0.3680.003 | 0.3620.004 | 0.4940.005 | 0.4550.007 | 0.3820.002 | 0.2390.002 | 0.8080.010 | 1.050.08 | [1] |
J23032925+3150210 | T2p | 0.7190.009 | 0.7550.008 | 0.4460.011 | 0.4900.005 | 0.4680.009 | 0.5430.011 | 0.5610.008 | 0.3680.004 | 0.2450.002 | 0.9110.010 | 1.000.12 | [31] |
0.7490.013 | 0.7630.012 | 0.4450.013 | 0.4780.007 | 0.4550.012 | 0.5440.014 | 0.5470.009 | 0.3450.006 | 0.2290.004 | 0.9390.015 | 0.990.15 | [29] | ||
J231854971301106 | T5 | 0.6990.008 | 0.4430.008 | 0.1980.007 | 0.3170.005 | 0.3770.009 | 0.4660.012 | 0.3440.007 | 0.4160.005 | 0.2350.003 | 0.7040.013 | 1.060.11 | [34] |
J23230708+0541130 | L8p (blue) | 0.8930.022 | 1.0070.019 | 0.8680.015 | 0.6940.022 | 0.6880.022 | 0.7180.015 | 0.6280.018 | 0.6710.012 | 0.6090.014 | 1.0360.020 | 0.960.12 | [32] |
J23254530+4251488 | L8 | 0.8720.019 | 1.0400.015 | 0.8840.015 | 0.6970.013 | 0.6820.012 | 0.6820.015 | 0.5810.010 | 0.6920.012 | 0.6570.012 | 1.0160.011 | 0.990.14 | [11] |
J232728742730566 | L9 | 0.8480.009 | 0.9860.010 | 0.8060.011 | 0.6030.009 | 0.6360.011 | 0.6780.009 | 0.5540.010 | 0.6370.007 | 0.5900.006 | 0.9940.009 | 0.960.07 | [16] |
J233051160844533 | L8.0 | 0.9150.026 | 1.0340.020 | 0.8630.011 | 0.6930.024 | 0.6390.017 | 0.7040.013 | 0.6340.020 | 0.7130.013 | 0.7410.018 | 1.0130.016 | 1.060.15 | [34] |
J23333346+0251284 | L8p (blue) | 0.9530.017 | 1.0750.016 | 1.0360.013 | 0.7920.018 | 0.8130.014 | 0.7530.011 | 0.4760.013 | 0.7060.010 | 0.6140.010 | 0.9680.013 | 0.860.09 | [32] |
J23352734+4511442 | L9p | 0.9540.014 | 1.1300.015 | 0.9760.009 | 0.6530.013 | 0.6570.012 | 0.7360.007 | 0.5570.015 | 0.8710.011 | 0.9360.010 | 1.0320.013 | 1.320.23 | [28] |
J23391025+1352284 | T5 | 0.5140.009 | 0.4580.006 | 0.2330.012 | 0.2440.003 | 0.2950.009 | 0.4600.009 | 0.4690.007 | 0.3430.004 | 0.1620.002 | 0.6470.014 | 1.030.09 | [3] |
J235654771553111 | T5.5 | 0.5480.010 | 0.3650.006 | 0.1400.009 | 0.2310.004 | 0.3140.006 | 0.4060.012 | 0.3960.007 | 0.3830.006 | 0.1930.002 | 0.5390.009 | 1.080.11 | [3] |
References. — [1] Burgasser et al. (2004);[2] Chiu et al. (2006);[3] Burgasser et al. (2006b);[4] Burgasser et al. (2006a);[5] Siegler et al. (2007);[6] Looper et al. (2007);[7] Luhman et al. (2007);[8] Burgasser (2007b);[9] Burgasser et al. (2008a);[10] Sheppard & Cushing (2009);[11] Burgasser et al. (2010a);[12] Kirkpatrick et al. (2010);[13] Deacon et al. (2011);[14] Gizis et al. (2011);[15] Geissler et al. (2011);[16] Kirkpatrick et al. (2011);[17] Deacon et al. (2012b);[18] Burgasser et al. (2013);[19] Castro et al. (2013);[20] Best et al. (2013);[21] Mace et al. (2013b);[22] Thompson et al. (2013);[23] Kirkpatrick et al. (2014);[24] Bowler et al. (2014);[25] Luhman & Sheppard (2014);[26] Bardalez Gagliuffi et al. (2014);[27] Kellogg et al. (2015);[28] Best et al. (2015);[29] Schneider et al. (2016a);[30] Schneider et al. (2016b);[31] Robert et al. (2016);[32] Schneider et al. (2017);[33] Kellogg et al. (2017);[34] Burgasser et al. (in prep)
Best-fit LOWZ Parameters | ||||||||
Designation | SpT | [M/H] | C/O | Ref. | ||||
(K) | (cm/s2) | (dex) | (cm2/s) | |||||
J00001354+2554180 | T4.5 | 1200 | 5.2 | 0.00 | 0.55 | 2 | 111 | [3] |
J00062785+1857288 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 29 | [28] |
J00150206+2959323 | L7 | 1600 | 4.5 | 0.25 | 0.85 | 10 | 23 | [12] |
J00164396+2304267 | T0.0 | 1600 | 4.5 | 0.00 | 0.55 | 2 | 18 | [27] |
J00164397+2304265 | T0 | 1600 | 4.5 | 0.00 | 0.55 | 2 | 31 | [28] |
J00185151+5153306 | L7.0 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 69 | [27] |
J00311004+5749363 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 38 | [20] |
1600 | 4.5 | 0.50 | 0.85 | 10 | 89 | [33] | ||
J00325937+1410371 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 10 | [11] |
1600 | 4.5 | 0.50 | 0.85 | 10 | 32 | [11] | ||
J00345157+0523050 | T6.5 | 850 | 4.5 | 0.25 | 0.55 | 1 | 187 | [1] |
J00440332+0228112 | L7p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 266 | [32] |
1600 | 4.5 | 0.75 | 0.85 | 10 | 100 | [31] | ||
J00454256+3611391 | T5 | 1100 | 4.5 | 0.25 | 0.55 | 1 | 53 | [21] |
J00492826+0440575 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 33 | [16] |
J00550564+0134365 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 297 | [32] |
J01020186+0355405 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 12 | [28] |
J01352531+0205232 | L9.5 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 143 | [32] |
J01365662+0933473 | T2.5 | 1400 | 5.0 | 0.00 | 0.55 | 2 | 1364 | [9] |
J013836580322212 | T3 | 1400 | 5.0 | 0.00 | 0.55 | 2 | 96 | [16] |
J014148391601196 | L7: | 1600 | 4.5 | 0.50 | 0.85 | 10 | 9 | [12] |
J014535200314117 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 76 | [32] |
J01501089+3827241 | T0 | 1600 | 4.5 | 0.00 | 0.55 | 2 | 97 | [16] |
1600 | 4.5 | 0.25 | 0.85 | 10 | 11 | [30] | ||
J01514155+1244300 | T1 | 1600 | 4.5 | 0.25 | 0.55 | 1 | 26 | [1] |
J02022929+2305139 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 76 | [31] |
J02062527+2640236 | L9p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 120 | [16] |
J02101025+4008296 | T4.5 | 1200 | 5.2 | 0.00 | 0.55 | 1 | 14 | [21] |
J022609161610004 | L7p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 156 | [31] |
J023038850225471 | L8:p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 14 | [33] |
J024313712453298 | T6 | 1000 | 5.0 | 0.25 | 0.55 | 2 | 112 | [1] |
J025503574700509 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 125 | [3] |
J025725813105523 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 61 | [5] |
J03162759+2650277 | T2.5 | 1300 | 5.0 | 0.25 | 0.55 | 2 | 48 | [30] |
J03230186+5625580 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 131 | [24] |
J032440231919056 | L8p | 1600 | 4.5 | 0.50 | 0.85 | 10 | 38 | [31] |
J03255322+0425406 | T5.5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 14 | [2] |
J03284265+2302051 | L8 | 1500 | 4.5 | 0.25 | 0.85 | 10 | 36 | [9] |
J03490944+1514360 | L7p | 1500 | 4.5 | 1.00 | 0.85 | 10 | 220 | [27] |
J04070885+1514565 | T5 | 1100 | 5.0 | 0.25 | 0.55 | 2 | 25 | [1] |
J041519540935066 | T8 | 850 | 4.5 | 0.00 | 0.85 | 1 | 133 | [1] |
J04373705+2331080 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 36 | [23] |
J045746020207179 | T2 | 1400 | 5.2 | 0.25 | 0.55 | 1 | 274 | [33] |
J05065020+5236312 | T4.5 | 1100 | 4.5 | 0.25 | 0.55 | 2 | 20 | [27] |
J051609450445499 | T5.5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 38 | [9] |
J055919141404488 | T4.5 | 1200 | 5.2 | 0.00 | 0.55 | 1 | 64 | [3] |
J06020638+4043588 | T4.5 | 1100 | 4.5 | 0.25 | 0.55 | 2 | 26 | [6] |
J06073865+2429535 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 1542 | [22] |
1600 | 4.5 | 0.50 | 0.85 | 10 | 639 | [19] | ||
J061135130410240 | T0 | 1600 | 4.5 | 0.25 | 0.55 | 2 | 11 | [16] |
J062720071114280 | T6 | 1000 | 4.5 | 0.25 | 0.55 | 2 | 23 | [16] |
J06420559+4102000 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 53 | [27] |
J070534001839256 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 20 | [31] |
J07271824+1710012 | T7 | 950 | 4.5 | 0.25 | 0.55 | 2 | 28 | [4] |
J07414920+2351275 | T5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 23 | [11] |
J07420130+2055198 | T5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 35 | [11] |
J07483861+1743329 | L7 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 32 | [25] |
J07554795+2212169 | T5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 18 | [3] |
J07584037+3247245 | T2 | 1400 | 4.5 | 0.25 | 0.55 | 2 | 391 | [9] |
J08070023+4130268 | L8p | 1600 | 4.5 | 0.50 | 0.85 | 10 | 119 | [22] |
J081958050335285 | T4 | 1200 | 5.2 | 0.25 | 0.55 | 2 | 36 | [16] |
J08203013+1037372 | L9.52 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 23 | [2] |
J08300825+4828482 | L8 | 1500 | 4.5 | 0.50 | 0.85 | 10 | 62 | [9] |
1600 | 5.2 | 0.75 | 0.55 | 1 | 200 | [9] | ||
J08304878+0128311 | T4.5 | 1100 | 4.5 | 0.25 | 0.55 | 2 | 36 | [11] |
J083717180000179 | T1 | 1600 | 4.5 | 0.25 | 0.55 | 1 | 17 | [33] |
J08523490+4720359 | L9.51 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 37 | [11] |
J08575849+5708514 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 700 | [11] |
J08583467+3256275 | T1 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 32 | [11] |
1500 | 4.5 | 0.75 | 0.85 | 10 | 19 | [2] | ||
J09002368+2539343 | L7 | 1600 | 4.5 | 0.00 | 0.55 | 2 | 52 | [25] |
J09025899+6708331 | L7p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 631 | [31] |
J09293364+3429527 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 37 | [11] |
J09310955+0327331 | L7.51.5 | 1600 | 4.5 | 0.25 | 0.55 | 10 | 32 | [33] |
J095533360208403 | L7p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 60 | [31] |
J095608101447065 | L9.5 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 40 | [30] |
J10071185+1930563 | L81.5 | 1600 | 4.5 | 0.00 | 0.55 | 2 | 30 | [2] |
1600 | 4.5 | 0.75 | 0.85 | 10 | 19 | [11] | ||
J10255227+3212349 | L7.52.5 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 16 | [2] |
1600 | 4.5 | 0.50 | 0.85 | 10 | 8 | [25] | ||
J10271549+5445175 | L7 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 93 | [32] |
J10302678+0213057 | L9.51 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 16 | [11] |
J10430758+2225236 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 44 | [5] |
J10433508+1213149 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 18 | [12] |
1600 | 4.5 | 0.50 | 0.85 | 10 | 53 | [11] | ||
J10440942+0429376 | L7 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 55 | [11] |
J10475385+2124234 | T6.5 | 900 | 5.0 | 0.25 | 0.55 | 1 | 62 | [9] |
J10482926+0919373 | T2.5 | 1400 | 5.0 | 0.50 | 0.10 | 1 | 27 | [11] |
J104915575319061A | L8 | 1500 | 4.5 | 1.00 | 0.85 | 10 | 778 | [18] |
J104915575319061B | T0.5 | 1600 | 4.5 | 0.25 | 0.55 | 1 | 252 | [18] |
J10595185+3042059 | T4 | 1200 | 5.2 | 0.25 | 0.55 | 1 | 13 | [10] |
J112208263512363 | T2 | 1300 | 4.5 | 0.25 | 0.55 | 2 | 86 | [3] |
J114724212040204 | L7 | 1500 | 4.5 | 1.00 | 0.85 | 10 | 203 | [29] |
J11555389+0559577 | L7.5 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 78 | [11] |
J12010457+5730042 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 19 | [31] |
J12060248+2813293 | T3 | 1200 | 5.2 | 0.25 | 0.55 | 10 | 26 | [2] |
J12144089+6316434 | T3.51 | 1200 | 5.0 | 0.75 | 0.55 | 2 | 37 | [2] |
J12195156+3128497 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 61 | [11] |
J12314753+0847331 | T5.5 | 1100 | 4.5 | 0.25 | 0.55 | 1 | 91 | [1] |
J125453930122474 | T2 | 1400 | 5.2 | 0.25 | 0.55 | 2 | 80 | [1] |
J130729840558146 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 30 | [28] |
J13184567+3626138 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 34 | [32] |
J13240767+1906259 | T3.5 | 1200 | 5.2 | 0.00 | 0.55 | 2 | 118 | [13] |
J132629810038314 | L8: | 1600 | 4.5 | 0.75 | 0.85 | 10 | 196 | [11] |
J134646340031501 | T6.5 | 1000 | 4.5 | 0.00 | 0.55 | 1 | 30 | [4] |
J13480699+6603284 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 18 | [16] |
J13585269+3747137 | T4.51 | 1000 | 5.0 | 0.50 | 0.55 | 1 | 36 | [2] |
J14002320+4338222 | L71 | 1600 | 5.0 | 0.50 | 0.85 | 10 | 37 | [11] |
J14321117+3244338 | L7 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 52 | [31] |
J14564268+6450097 | L7p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 154 | [31] |
J15031961+2525196 | T5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 114 | [1] |
J150502372853452 | T1.5 | 1500 | 5.0 | 0.00 | 0.55 | 2 | 55 | [13] |
J15131461+4019356 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 8 | [22] |
J15150607+4436483 | L7.51.5 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 39 | [11] |
J15164306+3053443 | T0.51 | 1500 | 4.5 | 0.75 | 0.85 | 10 | 32 | [11] |
J15203974+3546210 | T01 | 1600 | 5.2 | 0.00 | 0.55 | 2 | 15 | [33] |
1500 | 4.5 | 0.25 | 0.85 | 10 | 106 | [11] | ||
J15210327+0131426 | T2: | 1300 | 5.0 | 0.25 | 0.55 | 2 | 19 | [11] |
J15232263+3014562 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 26 | [11] |
J15261405+2043414 | L7 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 102 | [1] |
J15335852+4757069 | L8: | 1600 | 4.5 | 0.75 | 0.85 | 10 | 29 | [31] |
J15400942+3742316 | L91.5 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 20 | [2] |
1600 | 4.5 | 0.50 | 0.85 | 10 | 26 | [11] | ||
J154236300045452 | L7 | 1600 | 4.5 | 0.25 | 0.55 | 10 | 26 | [15] |
J15450901+3555271 | L7.5 | 1600 | 5.0 | 0.50 | 0.85 | 10 | 10 | [33] |
J15461460+4932113 | T2.51 | 1300 | 4.5 | 0.25 | 0.55 | 2 | 28 | [11] |
J154627183325111 | T5.5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 26 | [9] |
J16241436+0029158 | T6 | 1000 | 5.2 | 0.00 | 0.55 | 1 | 64 | [4] |
J16272565+3255246 | T6 | 900 | 4.5 | 0.50 | 0.55 | 1 | 33 | [16] |
J16291855+0335355 | T2 | 1400 | 5.2 | 0.25 | 0.55 | 2 | 188 | [17] |
J16303054+4344032 | L71.5 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 37 | [11] |
J16322911+1904407 | L8 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 42 | [8] |
J16323934+4150048 | T1: | 1400 | 4.5 | 0.25 | 0.55 | 2 | 8 | [11] |
J16471557+5632083 | L9p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 42 | [16] |
J17342104+5023508 | T4 | 1200 | 5.2 | 0.00 | 0.55 | 1 | 46 | [21] |
J17385927+6142421 | L9p | 1500 | 4.5 | 1.00 | 0.85 | 10 | 128 | [21] |
J17502385+4222373 | T2 | 1500 | 4.5 | 0.25 | 0.55 | 2 | 23 | [3] |
J17503293+1759042 | T3.5 | 1200 | 5.2 | 0.25 | 0.55 | 2 | 37 | [1] |
J17545447+1649196 | T5.5 | 1100 | 4.5 | 0.25 | 0.55 | 1 | 58 | [9] |
J17551028+1803202 | T2 | 1400 | 4.5 | 1.00 | 0.10 | 1 | 40 | [24] |
J18002660+0134531 | L7.5 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 580 | [14] |
J180952530448125 | T0.5 | 1600 | 4.5 | 0.25 | 0.55 | 2 | 47 | [21] |
1600 | 4.5 | 0.25 | 0.55 | 2 | 24 | [20] | ||
1600 | 4.5 | 0.25 | 0.55 | 2 | 24 | [27] | ||
J18305856+4542574 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 11 | [16] |
J19010601+4718136 | T5 | 1100 | 5.0 | 0.25 | 0.55 | 1 | 69 | [1] |
J19062474+4508071 | T6 | 1000 | 5.0 | 0.25 | 0.55 | 1 | 24 | [16] |
J19284135+2356049 | T6 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 518 | [21] |
1100 | 5.2 | 0.25 | 0.55 | 1 | 121 | [33] | ||
J20304279+0749347 | T1.5 | 1400 | 5.0 | 0.00 | 0.55 | 2 | 721 | [21] |
1400 | 4.5 | 0.50 | 0.55 | 2 | 43 | [20] | ||
J204317691551031 | L9 | 1500 | 5.0 | 0.50 | 0.85 | 10 | 13 | [11] |
J20435642+6220490 | T1.5 | 1500 | 5.2 | 0.50 | 0.10 | 1 | 18 | [21] |
J204749590718176 | T0: | 1600 | 5.2 | 0.00 | 0.55 | 2 | 26 | [11] |
J205201742043119 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 28 | [28] |
J211714442940052 | T0: | 1600 | 5.2 | 0.00 | 0.55 | 2 | 5 | [27] |
J21241387+0059599 | T5 | 1100 | 5.0 | 0.25 | 0.55 | 1 | 26 | [3] |
J21243865+1849263 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 36 | [26] |
J213248981452544 | T4 | 1200 | 5.2 | 0.00 | 0.55 | 2 | 14 | [33] |
J21392676+0220226 | T1.5 | 1500 | 4.5 | 0.25 | 0.55 | 1 | 882 | [3] |
J21442847+1446077 | T2.5 | 1300 | 4.5 | 0.50 | 0.55 | 2 | 113 | [7] |
J215138394853542 | T4 | 1200 | 5.2 | 0.00 | 0.55 | 2 | 35 | [3] |
J21513979+3402444 | L7p | 1600 | 5.0 | 0.50 | 0.85 | 10 | 24 | [12] |
J215424941023022 | T4.5 | 1200 | 5.2 | 0.00 | 0.55 | 1 | 20 | [6] |
J220921832711329 | T2.5 | 1300 | 4.5 | 0.25 | 0.55 | 2 | 64 | [33] |
J22221993+3026014 | L9 | 1600 | 5.2 | 0.50 | 0.85 | 10 | 25 | [25] |
J22361759+5105519 | T5.5 | 1100 | 4.5 | 0.25 | 0.55 | 2 | 176 | [20] |
J22393755+1617161 | T3 | 1300 | 5.0 | 0.25 | 0.55 | 2 | 106 | [16] |
J22541892+3123498 | T4 | 1200 | 4.5 | 0.00 | 0.55 | 2 | 123 | [1] |
J23032925+3150210 | T2p | 1300 | 5.0 | 0.50 | 0.55 | 2 | 44 | [30] |
1300 | 5.0 | 0.50 | 0.55 | 2 | 41 | [28] | ||
J231854971301106 | T5 | 1200 | 4.5 | 0.00 | 0.55 | 1 | 53 | [33] |
J23230708+0541130 | L8p | 1600 | 4.5 | 0.50 | 0.85 | 10 | 16 | [31] |
J23254530+4251488 | L8 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 13 | [11] |
J232728742730566 | L9 | 1600 | 4.5 | 0.50 | 0.85 | 10 | 26 | [16] |
J233051160844533 | L8.0 | 1600 | 4.5 | 0.75 | 0.85 | 10 | 25 | [33] |
J23333346+0251284 | L8p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 31 | [31] |
J23352734+4511442 | L9p | 1600 | 4.5 | 0.75 | 0.85 | 10 | 157 | [27] |
J23391025+1352284 | T5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 38 | [3] |
J235654771553111 | T5.5 | 1100 | 5.2 | 0.25 | 0.55 | 2 | 29 | [3] |
References. — [1] Burgasser et al. (2004); [2] Chiu et al. (2006); [3] Burgasser et al. (2006b); [4] Burgasser et al. (2006a); [5] Siegler et al. (2007); [6] Looper et al. (2007); [7] Luhman et al. (2007); [8] Burgasser (2007b); [9] Burgasser et al. (2008a); [10] Sheppard & Cushing (2009); [11] Burgasser et al. (2010a); [12] Kirkpatrick et al. (2010); [13] Deacon et al. (2011); [14] Gizis et al. (2011); [15] Geissler et al. (2011); [16] Kirkpatrick et al. (2011); [17] Deacon et al. (2012b); [18] Burgasser et al. (2013); [19] Castro et al. (2013); [20] Best et al. (2013); [21] Mace et al. (2013b); [22] Thompson et al. (2013); [23] Bowler et al. (2014); [24] Luhman & Sheppard (2014); [25] Bardalez Gagliuffi et al. (2014); [26] Kellogg et al. (2015); [27] Best et al. (2015); [28] Schneider et al. (2016a); [29] Schneider et al. (2016b); [30] Robert et al. (2016); [31] Schneider et al. (2017); [32] Kellogg et al. (2017); [33] Burgasser et al. (in prep.)
Appendix B Bolometric Corrections from LOWZ Models
We provide in this section bolometric corrections computed from a subset of the LOWZ model grid (Meisner et al., 2021), specifically: = 5.0 (in cm/s2); = 2 (in cm2/s); C/O = 0.55; 2.5 [M/H] 1.0, and 500 K 1600 K. These corrections are for the MKO and WISE filters and are computed on the Vega system. See Section 7.1 for details on these calculations.
Bolometric Correction | ||||||||
[M/H] | ||||||||
(K) | (dex) | (mag) | (mag) | (mag) | (mag) | (mag) | (mag) | (mag) |
500 | 2.50 | 20.86 | 0.94 | 1.71 | 0.13 | 5.84 | 4.58 | 6.02 |
500 | 2.00 | 20.86 | 0.61 | 1.29 | 0.11 | 5.18 | 4.35 | 6.13 |
500 | 1.50 | 20.86 | 0.17 | 0.56 | 0.36 | 4.19 | 4.00 | 6.26 |
500 | 1.00 | 20.86 | 0.24 | 0.33 | 0.61 | 2.82 | 3.52 | 6.36 |
500 | 0.50 | 20.86 | 0.63 | 1.25 | 0.86 | 1.34 | 2.99 | 6.33 |
500 | 0.25 | 20.86 | 0.83 | 1.62 | 1.08 | 0.72 | 2.70 | 6.24 |
500 | 0.00 | 20.86 | 0.93 | 1.91 | 1.22 | 0.01 | 2.46 | 6.14 |
500 | 0.25 | 20.86 | 0.92 | 2.08 | 1.30 | 0.61 | 2.26 | 6.00 |
500 | 0.50 | 20.86 | 0.94 | 2.19 | 1.35 | 1.17 | 2.10 | 5.84 |
500 | 0.75 | 20.86 | 0.88 | 2.16 | 1.37 | 1.63 | 2.01 | 5.66 |
500 | 1.00 | 20.86 | 0.90 | 2.09 | 1.30 | 1.90 | 1.96 | 5.45 |
550 | 2.50 | 20.45 | 0.44 | 1.04 | 0.26 | 4.83 | 4.56 | 5.95 |
550 | 2.00 | 20.45 | 0.08 | 0.61 | 0.42 | 4.25 | 4.33 | 6.06 |
550 | 1.50 | 20.45 | 0.25 | 0.05 | 0.67 | 3.30 | 4.00 | 6.19 |
550 | 1.00 | 20.45 | 0.43 | 0.90 | 0.99 | 1.98 | 3.55 | 6.27 |
550 | 0.50 | 20.45 | 0.95 | 1.67 | 1.24 | 0.73 | 3.01 | 6.20 |
550 | 0.25 | 20.45 | 0.99 | 1.96 | 1.37 | 0.13 | 2.75 | 6.13 |
550 | 0.00 | 20.45 | 0.99 | 2.16 | 1.47 | 0.52 | 2.54 | 6.05 |
550 | 0.25 | 20.45 | 0.96 | 2.30 | 1.54 | 1.08 | 2.35 | 5.94 |
550 | 0.50 | 20.45 | 1.06 | 2.35 | 1.53 | 1.52 | 2.20 | 5.82 |
550 | 0.75 | 20.45 | 1.18 | 2.31 | 1.48 | 1.86 | 2.09 | 5.67 |
550 | 1.00 | 20.45 | 1.09 | 2.24 | 1.41 | 2.07 | 2.06 | 5.50 |
600 | 2.50 | 20.07 | 0.08 | 0.32 | 0.55 | 3.79 | 4.57 | 5.93 |
600 | 2.00 | 20.07 | 0.18 | 0.07 | 0.79 | 3.28 | 4.35 | 6.02 |
600 | 1.50 | 20.07 | 0.62 | 0.67 | 1.04 | 2.49 | 3.99 | 6.11 |
600 | 1.00 | 20.07 | 0.98 | 1.35 | 1.25 | 1.30 | 3.54 | 6.17 |
600 | 0.50 | 20.07 | 1.10 | 1.98 | 1.51 | 0.22 | 3.02 | 6.08 |
600 | 0.25 | 20.07 | 1.13 | 2.19 | 1.62 | 0.35 | 2.78 | 6.01 |
600 | 0.00 | 20.07 | 1.22 | 2.36 | 1.71 | 0.92 | 2.58 | 5.91 |
600 | 0.25 | 20.07 | 1.19 | 2.45 | 1.74 | 1.43 | 2.41 | 5.77 |
600 | 0.50 | 20.07 | 1.20 | 2.50 | 1.75 | 1.87 | 2.32 | 5.61 |
600 | 0.75 | 20.07 | 1.17 | 2.46 | 1.70 | 2.17 | 2.29 | 5.41 |
600 | 1.00 | 20.07 | 1.13 | 2.44 | 1.67 | 2.35 | 2.29 | 5.20 |
650 | 2.50 | 19.72 | 0.23 | 0.10 | 0.81 | 3.12 | 4.53 | 5.84 |
650 | 2.00 | 19.72 | 0.54 | 0.48 | 0.98 | 2.58 | 4.29 | 5.92 |
650 | 1.50 | 19.72 | 0.88 | 1.02 | 1.18 | 1.79 | 3.95 | 6.01 |
650 | 1.00 | 19.72 | 1.08 | 1.66 | 1.48 | 0.78 | 3.51 | 6.05 |
650 | 0.50 | 19.72 | 1.30 | 2.20 | 1.69 | 0.20 | 3.02 | 5.97 |
650 | 0.25 | 19.72 | 1.26 | 2.37 | 1.79 | 0.72 | 2.80 | 5.90 |
650 | 0.00 | 19.72 | 1.07 | 2.38 | 1.85 | 1.25 | 2.61 | 5.82 |
650 | 0.25 | 19.72 | 1.24 | 2.55 | 1.86 | 1.70 | 2.45 | 5.72 |
650 | 0.50 | 19.72 | 1.27 | 2.59 | 1.85 | 2.07 | 2.36 | 5.61 |
650 | 0.75 | 19.72 | 1.22 | 2.55 | 1.78 | 2.28 | 2.30 | 5.46 |
650 | 1.00 | 19.72 | 1.24 | 2.57 | 1.76 | 2.43 | 2.31 | 5.27 |
700 | 2.50 | 19.40 | 0.39 | 0.67 | 1.06 | 2.27 | 4.52 | 5.79 |
700 | 2.00 | 19.40 | 0.60 | 1.00 | 1.27 | 1.79 | 4.30 | 5.87 |
700 | 1.50 | 19.40 | 0.95 | 1.42 | 1.47 | 1.13 | 3.94 | 5.93 |
700 | 1.00 | 19.40 | 1.35 | 1.95 | 1.65 | 0.36 | 3.47 | 5.93 |
700 | 0.50 | 19.40 | 1.39 | 2.37 | 1.86 | 0.51 | 2.99 | 5.84 |
700 | 0.25 | 19.40 | 1.38 | 2.52 | 1.95 | 1.02 | 2.79 | 5.78 |
700 | 0.00 | 19.40 | 1.38 | 2.59 | 1.97 | 1.50 | 2.62 | 5.68 |
700 | 0.25 | 19.40 | 1.35 | 2.62 | 1.96 | 1.93 | 2.50 | 5.55 |
700 | 0.50 | 19.40 | 1.34 | 2.64 | 1.95 | 2.26 | 2.44 | 5.39 |
700 | 0.75 | 19.40 | 1.34 | 2.67 | 1.94 | 2.48 | 2.45 | 5.21 |
700 | 1.00 | 19.40 | 1.41 | 2.77 | 1.89 | 2.59 | 2.49 | 5.02 |
750 | 2.50 | 19.10 | 0.61 | 0.91 | 1.21 | 1.79 | 4.46 | 5.69 |
750 | 2.00 | 19.10 | 0.84 | 1.26 | 1.42 | 1.31 | 4.23 | 5.76 |
750 | 1.50 | 19.10 | 1.19 | 1.68 | 1.60 | 0.70 | 3.88 | 5.81 |
750 | 1.00 | 19.10 | 1.43 | 2.12 | 1.77 | 0.02 | 3.41 | 5.81 |
750 | 0.50 | 19.10 | 1.46 | 2.49 | 2.00 | 0.82 | 2.98 | 5.73 |
750 | 0.25 | 19.10 | 1.44 | 2.61 | 2.05 | 1.29 | 2.80 | 5.63 |
750 | 0.00 | 19.10 | 1.43 | 2.67 | 2.08 | 1.74 | 2.65 | 5.52 |
750 | 0.25 | 19.10 | 1.46 | 2.75 | 2.11 | 2.17 | 2.57 | 5.35 |
750 | 0.50 | 19.10 | 1.40 | 2.71 | 2.07 | 2.43 | 2.52 | 5.16 |
750 | 0.75 | 19.10 | 1.52 | 2.84 | 2.08 | 2.59 | 2.53 | 4.96 |
750 | 1.00 | 19.10 | 1.39 | 2.77 | 2.03 | 2.71 | 2.64 | 4.86 |
800 | 2.50 | 18.82 | 0.69 | 1.20 | 1.39 | 1.27 | 4.42 | 5.59 |
800 | 2.00 | 18.82 | 0.97 | 1.49 | 1.54 | 0.84 | 4.17 | 5.65 |
800 | 1.50 | 18.82 | 1.21 | 1.85 | 1.74 | 0.29 | 3.83 | 5.71 |
800 | 1.00 | 18.82 | 1.34 | 2.24 | 1.96 | 0.36 | 3.40 | 5.71 |
800 | 0.50 | 18.82 | 1.41 | 2.51 | 2.10 | 1.09 | 2.97 | 5.63 |
800 | 0.25 | 18.82 | 1.53 | 2.66 | 2.15 | 1.51 | 2.79 | 5.56 |
800 | 0.00 | 18.82 | 1.49 | 2.73 | 2.12 | 1.88 | 2.63 | 5.48 |
800 | 0.25 | 18.82 | 1.53 | 2.81 | 2.13 | 2.25 | 2.55 | 5.38 |
800 | 0.50 | 18.82 | 1.56 | 2.87 | 2.11 | 2.48 | 2.52 | 5.22 |
800 | 0.75 | 18.82 | 1.60 | 2.95 | 2.13 | 2.64 | 2.56 | 5.05 |
800 | 1.00 | 18.82 | 1.42 | 2.79 | 2.05 | 2.71 | 2.62 | 4.89 |
850 | 2.50 | 18.56 | 0.83 | 1.43 | 1.50 | 0.83 | 4.36 | 5.49 |
850 | 2.00 | 18.56 | 0.87 | 1.69 | 1.74 | 0.34 | 4.16 | 5.58 |
850 | 1.50 | 18.56 | 1.09 | 1.99 | 1.94 | 0.14 | 3.82 | 5.63 |
850 | 1.00 | 18.56 | 1.40 | 2.33 | 2.07 | 0.63 | 3.36 | 5.59 |
850 | 0.50 | 18.56 | 1.56 | 2.65 | 2.16 | 1.26 | 2.92 | 5.50 |
850 | 0.25 | 18.56 | 1.53 | 2.69 | 2.20 | 1.68 | 2.78 | 5.44 |
850 | 0.00 | 18.56 | 1.48 | 2.67 | 2.22 | 2.09 | 2.68 | 5.34 |
850 | 0.25 | 18.56 | 1.59 | 2.75 | 2.23 | 2.42 | 2.63 | 5.18 |
850 | 0.50 | 18.56 | 1.85 | 3.00 | 2.33 | 2.61 | 2.62 | 5.02 |
850 | 0.75 | 18.56 | 1.85 | 3.16 | 2.30 | 2.72 | 2.65 | 4.86 |
850 | 1.00 | 18.56 | 1.63 | 2.97 | 2.23 | 2.80 | 2.72 | 4.74 |
900 | 2.50 | 18.31 | 0.96 | 1.55 | 1.58 | 0.54 | 4.29 | 5.39 |
900 | 2.00 | 18.31 | 1.06 | 1.83 | 1.80 | 0.09 | 4.08 | 5.46 |
900 | 1.50 | 18.31 | 1.25 | 2.12 | 2.00 | 0.36 | 3.74 | 5.50 |
900 | 1.00 | 18.31 | 1.45 | 2.41 | 2.14 | 0.86 | 3.31 | 5.49 |
900 | 0.50 | 18.31 | 1.48 | 2.57 | 2.33 | 1.56 | 2.96 | 5.41 |
900 | 0.25 | 18.31 | 1.54 | 2.70 | 2.31 | 1.90 | 2.81 | 5.32 |
900 | 0.00 | 18.31 | 1.65 | 2.86 | 2.33 | 2.25 | 2.72 | 5.18 |
900 | 0.25 | 18.31 | 1.70 | 2.92 | 2.34 | 2.51 | 2.69 | 5.00 |
900 | 0.50 | 18.31 | 1.65 | 2.84 | 2.34 | 2.68 | 2.72 | 4.86 |
900 | 0.75 | 18.31 | 1.78 | 2.95 | 2.40 | 2.79 | 2.76 | 4.72 |
900 | 1.00 | 18.31 | 1.44 | 2.80 | 2.33 | 2.88 | 2.82 | 4.60 |
950 | 2.50 | 18.07 | 1.08 | 1.68 | 1.65 | 0.27 | 4.22 | 5.28 |
950 | 2.00 | 18.07 | 1.09 | 1.93 | 1.88 | 0.16 | 4.02 | 5.35 |
950 | 1.50 | 18.07 | 1.19 | 2.17 | 2.11 | 0.64 | 3.70 | 5.42 |
950 | 1.00 | 18.07 | 1.23 | 2.32 | 2.37 | 1.25 | 3.35 | 5.45 |
950 | 0.50 | 18.07 | 1.54 | 2.63 | 2.38 | 1.70 | 2.93 | 5.33 |
950 | 0.25 | 18.07 | 1.64 | 2.68 | 2.43 | 2.04 | 2.81 | 5.27 |
950 | 0.00 | 18.07 | 1.60 | 2.66 | 2.44 | 2.39 | 2.76 | 5.20 |
950 | 0.25 | 18.07 | 1.61 | 2.86 | 2.35 | 2.54 | 2.67 | 5.05 |
950 | 0.50 | 18.07 | 1.58 | 2.83 | 2.39 | 2.70 | 2.70 | 4.90 |
950 | 0.75 | 18.07 | 1.63 | 2.81 | 2.39 | 2.79 | 2.74 | 4.76 |
950 | 1.00 | 18.07 | 1.82 | 2.94 | 2.52 | 2.90 | 2.82 | 4.64 |
1000 | 2.50 | 17.85 | 1.01 | 1.79 | 1.77 | 0.08 | 4.19 | 5.20 |
1000 | 2.00 | 17.85 | 0.98 | 1.97 | 2.04 | 0.54 | 4.01 | 5.30 |
1000 | 1.50 | 17.85 | 1.04 | 2.12 | 2.28 | 0.99 | 3.71 | 5.37 |
1000 | 1.00 | 17.85 | 1.31 | 2.37 | 2.41 | 1.39 | 3.30 | 5.34 |
1000 | 0.50 | 17.85 | 1.49 | 2.62 | 2.46 | 1.87 | 2.93 | 5.24 |
1000 | 0.25 | 17.85 | 1.57 | 2.71 | 2.47 | 2.17 | 2.82 | 5.15 |
1000 | 0.00 | 17.85 | 1.60 | 2.73 | 2.52 | 2.48 | 2.79 | 5.04 |
1000 | 0.25 | 17.85 | 1.65 | 2.78 | 2.49 | 2.63 | 2.76 | 4.90 |
1000 | 0.50 | 17.85 | 1.70 | 2.92 | 2.52 | 2.74 | 2.77 | 4.75 |
1000 | 0.75 | 17.85 | 1.86 | 3.05 | 2.67 | 2.88 | 2.82 | 4.63 |
1000 | 1.00 | 17.85 | 1.65 | 2.86 | 2.62 | 2.95 | 2.89 | 4.51 |
1100 | 2.50 | 17.44 | 1.16 | 1.94 | 1.84 | 0.40 | 4.04 | 4.99 |
1100 | 2.00 | 17.44 | 1.18 | 2.11 | 2.10 | 0.83 | 3.87 | 5.09 |
1100 | 1.50 | 17.44 | 1.18 | 2.28 | 2.34 | 1.24 | 3.58 | 5.16 |
1100 | 1.00 | 17.44 | 1.07 | 2.22 | 2.61 | 1.88 | 3.33 | 5.23 |
1100 | 0.50 | 17.44 | 1.35 | 2.47 | 2.62 | 2.22 | 2.97 | 5.08 |
1100 | 0.25 | 17.44 | 1.34 | 2.45 | 2.66 | 2.49 | 2.91 | 5.02 |
1100 | 0.00 | 17.44 | 1.34 | 2.42 | 2.69 | 2.70 | 2.90 | 4.93 |
1100 | 0.25 | 17.44 | 1.61 | 2.77 | 2.69 | 2.72 | 2.81 | 4.78 |
1100 | 0.50 | 17.44 | 1.56 | 2.72 | 2.76 | 2.86 | 2.86 | 4.66 |
1100 | 0.75 | 17.44 | 1.58 | 2.72 | 2.87 | 3.02 | 2.98 | 4.41 |
1100 | 1.00 | 17.44 | 1.73 | 2.84 | 2.83 | 3.04 | 3.01 | 4.27 |
1200 | 2.50 | 17.06 | 1.13 | 2.05 | 1.93 | 0.75 | 3.94 | 4.82 |
1200 | 2.00 | 17.06 | 1.02 | 2.15 | 2.23 | 1.23 | 3.81 | 4.96 |
1200 | 1.50 | 17.06 | 1.08 | 2.20 | 2.48 | 1.65 | 3.55 | 5.04 |
1200 | 1.00 | 17.06 | 1.20 | 2.33 | 2.64 | 2.00 | 3.21 | 5.01 |
1200 | 0.50 | 17.06 | 1.14 | 2.27 | 2.76 | 2.54 | 3.06 | 4.89 |
1200 | 0.25 | 17.06 | 1.16 | 2.29 | 2.78 | 2.71 | 3.02 | 4.80 |
1200 | 0.00 | 17.06 | 1.29 | 2.42 | 2.77 | 2.76 | 2.95 | 4.68 |
1200 | 0.25 | 17.06 | 1.26 | 2.40 | 2.83 | 2.91 | 2.99 | 4.58 |
1200 | 0.50 | 17.06 | 1.58 | 2.69 | 2.84 | 2.89 | 2.94 | 4.41 |
1200 | 0.75 | 17.06 | 1.51 | 2.61 | 2.88 | 3.02 | 3.01 | 4.31 |
1200 | 1.00 | 17.06 | 1.47 | 2.68 | 2.85 | 3.07 | 3.05 | 4.18 |
1300 | 2.50 | 16.71 | 1.18 | 2.11 | 1.96 | 0.94 | 3.81 | 4.65 |
1300 | 2.00 | 16.71 | 1.07 | 2.22 | 2.27 | 1.41 | 3.70 | 4.78 |
1300 | 1.50 | 16.71 | 0.95 | 2.21 | 2.55 | 1.91 | 3.50 | 4.89 |
1300 | 1.00 | 16.71 | 1.05 | 2.20 | 2.72 | 2.31 | 3.23 | 4.88 |
1300 | 0.50 | 16.71 | 1.23 | 2.32 | 2.79 | 2.60 | 3.01 | 4.72 |
1300 | 0.25 | 16.71 | 1.13 | 2.24 | 2.82 | 2.79 | 3.02 | 4.66 |
1300 | 0.00 | 16.71 | 1.05 | 2.19 | 2.85 | 2.94 | 3.06 | 4.58 |
1300 | 0.25 | 16.71 | 1.40 | 2.49 | 2.96 | 2.99 | 3.02 | 4.45 |
1300 | 0.50 | 16.71 | 1.37 | 2.45 | 2.99 | 3.16 | 3.13 | 4.26 |
1300 | 0.75 | 16.71 | 1.48 | 2.53 | 2.88 | 3.06 | 3.07 | 4.09 |
1300 | 1.00 | 16.71 | 1.46 | 2.47 | 2.87 | 3.19 | 3.18 | 3.96 |
1400 | 2.50 | 16.39 | 1.20 | 2.16 | 2.00 | 1.13 | 3.71 | 4.49 |
1400 | 2.00 | 16.39 | 1.02 | 2.25 | 2.31 | 1.60 | 3.62 | 4.64 |
1400 | 1.50 | 16.39 | 0.95 | 2.23 | 2.58 | 2.04 | 3.41 | 4.73 |
1400 | 1.00 | 16.39 | 0.91 | 2.19 | 2.76 | 2.47 | 3.20 | 4.68 |
1400 | 0.50 | 16.39 | 0.98 | 2.13 | 2.84 | 2.81 | 3.13 | 4.56 |
1400 | 0.25 | 16.39 | 0.91 | 2.08 | 2.85 | 2.94 | 3.14 | 4.49 |
1400 | 0.00 | 16.39 | 1.11 | 2.20 | 2.89 | 3.00 | 3.12 | 4.39 |
1400 | 0.25 | 16.39 | 1.03 | 2.13 | 2.87 | 3.15 | 3.21 | 4.29 |
1400 | 0.50 | 16.39 | 1.02 | 2.12 | 2.88 | 3.28 | 3.32 | 4.17 |
1400 | 0.75 | 16.39 | 1.26 | 2.30 | 2.91 | 3.28 | 3.25 | 4.02 |
1400 | 1.00 | 16.39 | 1.24 | 2.24 | 2.88 | 3.39 | 3.44 | 3.80 |
1500 | 2.50 | 16.09 | 1.22 | 2.19 | 2.01 | 1.24 | 3.59 | 4.34 |
1500 | 2.00 | 16.09 | 1.08 | 2.28 | 2.32 | 1.69 | 3.50 | 4.48 |
1500 | 1.50 | 16.09 | 0.89 | 2.24 | 2.60 | 2.19 | 3.36 | 4.58 |
1500 | 1.00 | 16.09 | 0.85 | 2.16 | 2.78 | 2.57 | 3.18 | 4.52 |
1500 | 0.50 | 16.09 | 0.80 | 2.06 | 2.85 | 2.90 | 3.18 | 4.43 |
1500 | 0.25 | 16.09 | 0.79 | 2.04 | 2.86 | 3.02 | 3.21 | 4.36 |
1500 | 0.00 | 16.09 | 0.88 | 2.02 | 2.85 | 3.15 | 3.28 | 4.29 |
1500 | 0.25 | 16.09 | 0.90 | 2.04 | 2.87 | 3.26 | 3.37 | 4.15 |
1500 | 0.50 | 16.09 | 1.02 | 2.07 | 2.83 | 3.31 | 3.39 | 4.01 |
1500 | 0.75 | 16.09 | 1.03 | 2.06 | 2.82 | 3.41 | 3.53 | 3.89 |
1500 | 1.00 | 16.09 | 1.19 | 2.21 | 2.81 | 3.38 | 3.47 | 3.72 |
1600 | 2.50 | 15.81 | 1.25 | 2.20 | 2.01 | 1.34 | 3.49 | 4.22 |
1600 | 2.00 | 15.81 | 1.05 | 2.29 | 2.34 | 1.81 | 3.43 | 4.35 |
1600 | 1.50 | 15.81 | 0.93 | 2.26 | 2.61 | 2.25 | 3.27 | 4.41 |
1600 | 1.00 | 15.81 | 0.79 | 2.15 | 2.79 | 2.66 | 3.20 | 4.37 |
1600 | 0.50 | 15.81 | 0.73 | 2.04 | 2.85 | 2.97 | 3.27 | 4.27 |
1600 | 0.25 | 15.81 | 0.73 | 1.96 | 2.84 | 3.13 | 3.35 | 4.23 |
1600 | 0.00 | 15.81 | 0.77 | 1.99 | 2.85 | 3.21 | 3.37 | 4.14 |
1600 | 0.25 | 15.81 | 0.79 | 1.98 | 2.84 | 3.33 | 3.49 | 4.05 |
1600 | 0.50 | 15.81 | 0.96 | 2.02 | 2.81 | 3.40 | 3.57 | 3.95 |
1600 | 0.75 | 15.81 | 0.96 | 2.00 | 2.76 | 3.43 | 3.65 | 3.82 |
1600 | 1.00 | 15.81 | 1.01 | 2.04 | 2.77 | 3.46 | 3.73 | 3.65 |
Note. — Magnitudes are on the Vega system, and evaluation of assumes R = 0.077 R⊙. We further constrain the LOWZ models to have = 5 (cm/s2), = 2 (cm2/s), and C/O = 0.55.